### v9

# The Race between Technology and Woman:

Abstract
The era of technological change entails complex patterns of changes in wages and employment. We develop a unified framework to evaluate the effects of capital-embodied technological change on, as well as the contributions of factor inputs to, the relative wages and income shares of different types of labor. We obtain the aggregate elasticities of substitution among different types of capital and labor by estimating and aggregating sectoral production function parameters with cross-country cross-industry panel data from OECD countries. We show that advances in information, communication, and computation technologies contribute significantly to narrowing the gender wage gap, widening the skill wage gap, and declining labor shares.
Keywords: Gender wage gap; skill wage gap; labor share; capital–skill complementarity; capital-embodied technological change.
JEL classification: E23, E25, J16, J23, J31, O33.
1 Introduction
Over the decades, there have been substantial changes in the structure of wages and employment. The wage gap between male and female workers has narrowed in many countries, while that between skilled and unskilled workers has widened in some countries (Krueger, Perri, Pistaferri and Violante, 2010). At the same time, the share of labor in national income has declined in many countries (Karabarbounis and Neiman, 2014). During the period, there have been tremendous advances in information, communication, and computation technologies in industrialized nations. Such technological advances have been recognized as one of the causes for the widening of the skill wage gap and the decline in the labor share (Hornstein, Krusell and Violante, 2005; Acemoglu and Autor, 2011; Grossman and Oberfield, 2022; Karabarbounis, 2024). Nevertheless, changes in relative wages and income shares vary in direction, magnitude, or both across different types of labor classified by gender and skill, as described later.
The direction and magnitude of changes in the relative wages and income shares of different types of labor due to the expansion of technologically advanced new equipment depend on the presence and degree of capital–skill or capital–gender complementarity. Male and female workers possess different sets of skills (Welch, 2000; Borghans, ter Weel and Weinberg, 2014; Cortes, Jaimovich and Siu, 2023; Rendall, 2024), as skilled and unskilled workers do (Griliches, 1969). Communication skills, in which women can possibly have a comparative advantage, are more likely to be complementary to information and communication technologies. Physical skills, in which men can possibly have a comparative advantage, are more likely to be substitutable with computation and automation technologies. The widespread use of new technologies might raise the demand for brains relative to brawn, which could result in not only a widening of the skill wage gap but also a narrowing of the gender wage gap and a decline in the income share of male unskilled labor.
This study develops a unified framework to evaluate the effects of capital-embodied technological change on, as well as the contributions of factor inputs to, the relative wages and income shares of different types of labor. Our framework builds upon a multi-sector, multi-factor extension of the Greenwood, Hercowitz and Krusell (1997) model of investment-specific technological change. [^1] The terms, capital-embodied technological change and investment-specific technological change, are used interchangeably in this paper. We incorporate technological change embodied in new capital into an economy consisting of multiple sectors, where multiple consumption and investment goods are produced from multiple types of capital and labor. The presence and degrees of capital–skill and capital–gender complementarities in the multi-sector economy are determined by the aggregate elasticities of substitution among different types of capital and labor. We show that the aggregate elasticities of substitution are sufficient statistics for measuring the contributions of factor inputs to the relative wages and income shares of different types of labor. In this context, we generalize a model that explains changes in the wage premium to skill in terms of the race between technology and education (Tinbergen, 1974; Goldin and Katz, 2010) and a model that explains changes in the labor share in terms of capital deepening (Hicks, 1932; Elsby, Hobijn and Şahin, 2013). We derive analytical expressions that characterize not only the contribution of each input, holding the other inputs constant, but also the effects of capital-embodied technological change, allowing all inputs to vary. We further show that the aggregate elasticities of substitution are sufficient statistics for quantifying the effects of capital-embodied technological change on the relative wages and income shares of different types of labor. In this context, we synthesize models that explain a rise in the skill premium or a decline in the labor share in terms of a fall in the relative price of capital equipment (Krusell, Ohanian, Ríos-Rull and Violante, 2000; Karabarbounis and Neiman, 2014).
We apply this framework to advanced economies consisting of multiple sectors, where the production of goods and services requires multiple types of capital and labor. Given the fact that capital-embodied technological progress was accelerated by information processing equipment and software in the 1980s and 1990s (Cummins and Violante, 2002), we distinguish equipment related to information, communication, and computation technologies from other types of capital. We refer to the former as ICCT capital and the latter as non-ICCT capital. We classify labor by gender and skill into four types: male skilled, female skilled, male unskilled, and female unskilled labor. In doing so, we analyze how technological change embodied in ICCT capital shapes the observed differences in the direction and magnitude of changes in the gender wage gap by skill, the skill wage gap by gender, and labor shares by gender and skill.
Although it is notoriously difficult to identify the elasticities of substitution in the aggregate production function using time-series data (Diamond, McFadden and Rodriguez, 1978; Acemoglu, 2002), it is possible to obtain the aggregate elasticities of substitution by estimating and aggregating the disaggregated production function parameters with cross-sectional variation in disaggregated data (Sato, 1975; Oberfield and Raval, 2021). We estimate the sectoral production function parameters and obtain the aggregate elasticities of substitution between ICCT capital and the four types of labor using cross-country cross-industry data from OECD countries. The advantages of cross-country cross-industry data over firm-level data are that it enables us to measure equipment related to information, communication, and computation technologies and to classify labor by gender and skill. Our estimates of the aggregate elasticities of substitution indicate that ICCT capital is more complementary not only to skilled labor than unskilled labor but also to female labor than male labor. This finding is robust to the choice of instrumental variables and the presence of labor market institutions, factor-biased technological change, and price markups.
We decompose changes in the relative wages and labor shares of the four types of labor into components attributable to the factor complementarity effect associated with progress in ICCT and the relative quantity effect associated with advances in educational attainment and female employment, and then further into components attributable to the two types of capital and the four types of labor. Our results suggest that the expansion of ICCT equipment contributes to the widening of the skill wage gap, the narrowing of the gender wage gap, and a fall in the income share of male unskilled labor that accounts for most of the decline in the labor share. The observed difference in the rate of decrease (increase) in the gender (skill) wage gap by skill (gender) can be explained in terms of the race between progress in ICCT and advances in female employment (educational attainment). From the 1980s to the 2000s in OECD countries, the rate of decline in the gender wage gap was greater among unskilled workers than among skilled workers, while the rate of increase in the skill wage gap was greater among male workers than among female workers. Our results suggest that these differences do not arise from an insufficient increase in the relative demand for female skilled labor, but rather from a substantial increase in the relative supply of female skilled labor. Furthermore, we quantify the general equilibrium effects of technological change embodied in ICCT equipment on the gender wage gap by skill, the skill wage gap by gender, and labor shares by gender and skill. On average across OECD countries, if all else were held constant, a fall in the relative price of ICCT equipment between the years 1980 and 2005 would have reduced the skilled (unskilled) gender wage gap by 49 (10) percentage points and raised the male (female) skill wage gap by 26 (77) percentage points, while it would have raised the income share of male (female) skilled labor by 1.4 (2.0) percentage points and reduced the income share of male (female) unskilled labor by 7.0 (0.7) percentage points.
The remainder of this paper is structured as follows. The next section reviews the related literature. Section 3 establishes the theoretical framework to evaluate the effects of technological change on, as well as the contributions of factor inputs to, the relative wages and income shares of different types of labor. Section 4 describes the data used in the analysis and shows changes in the relative prices, relative quantities, and income shares of factors. Section 5 discusses the specification, identification, estimation, and aggregation of parameters in the sectoral production function. Section 6 presents the results regarding the sources of changes in the relative wages and income shares of different types of labor. The final section provides conclusions.
2 Related Literature
This study adds to several strands of literature. First, it is related to the literature on technological change and wage inequality. This strand of literature highlights the importance of factor-biased technological change in accounting for changes in wage inequality in the United States and other OECD countries, typically by estimating the aggregate production function. The widening of the skill wage gap is attributed to skill-biased technological change (Bound and Johnson, 1992; Katz and Murphy, 1992; Goldin and Katz, 2010), technological change embodied in capital equipment (Krusell, Ohanian, Ríos-Rull and Violante, 2000; Maliar, Maliar and Tsener, 2022; Taniguchi and Yamada, 2022; Ohanian, Orak and Shen, 2023), or sector-specific disembodied technological change (Buera, Kaboski, Rogerson and Vizcaino, 2022). The narrowing of the gender wage gap is also attributed to gender-biased technological change (Johnson and Keane, 2013). Relative to these studies, we obtain the aggregate elasticities of substitution by estimating and aggregating sectoral production function parameters, and examine the sources of changes in the four wage gaps: skilled gender wage gap, unskilled gender wage gap, male skill wage gap, and female skill wage gap.
Second, this study is related to the literature on the labor market impact of new technologies. This strand of literature indicates advances in information, communication, computation, and automation technologies as causes of changes in the structure of wages and employment in the United States and other OECD countries. The rate of increase in the wage bill share of skilled or female labor tends to be higher in ICCT or computer-intensive industries (Autor, Katz and Krueger, 1998; Michaels, Natraj and Van Reenen, 2014; Raveh, 2015). The increased employment and wages of skilled or female labor are associated with a rise in occupations requiring nonroutine analytical and interactive tasks that are complementary to ICCT and a decline in routine cognitive and manual tasks that are substitutable with ICCT (Autor, Levy and Murnane, 2003; Black and Spitz-Oener, 2010). The widening of the skill wage gap is attributed to a relative decline in industries requiring routine tasks that are substitutable with automation technology (Acemoglu and Restrepo, 2022). The widening of the skill wage gap is also attributed to the increased use of computers and other capital equipment, whereas the narrowing of the gender wage gap is not necessarily attributable to that increased use (Beaudry and Lewis, 2014; Burstein, Morales and Vogel, 2019; Caunedo, Jaume and Keller, 2023). [^2] Beaudry and Lewis (2014) consider an aggregate production function with a unit elasticity of substitution between computer equipment and skilled labor and an infinite elasticity of substitution between computer equipment and unskilled labor. Burstein, Morales and Vogel (2019) consider an aggregate production function, in which final output is produced from occupational outputs, which are in turn produced by production units, each using one type of capital (computing or other equipment) and one type of labor (classified by gender, education, and age). They specify an occupational production function with a unit elasticity of capital–labor substitution for all types of capital and labor. Caunedo, Jaume and Keller (2023) consider an aggregate production function, in which final output is produced from occupational outputs, which are in turn produced from capital and labor. They specify an occupational production function with an common elasticity of capital–labor substitution for all types of labor and an infinite elasticity of labor–labor substitution for all types of labor. Relative to these studies, we allow all four types of labor disaggregated by gender and skill to be imperfect substitutes for one another and substitutable to varying degrees with ICCT equipment within sectors, thereby permitting the separate identification of capital–skill and capital–gender complementarities. We impose neither specific values on the elasticities of substitution among the four types of labor nor on those between ICCT equipment and labor.
Third, this study is related to the literature on technological change and the labor share of income. This strand of literature points to the relevance of technological change to the decline in the labor share in the United States and elsewhere. The decline in the labor share is attributed to capital-embodied technological change (Karabarbounis and Neiman, 2014; Eden and Gaggl, 2018), factor-biased technological change (Oberfield and Raval, 2021), or the adoption of robotic technology (Acemoglu and Restrepo, 2020). Relative to these studies, we examine the sources of changes in labor shares by gender and skill. We decompose changes in the income share of the four types of labor into components attributable to the relative quantity effect, whose sign depends on whether the aggregate elasticities of substitution exceed one, and to the factor complementarity effect, whose sign is determined by the relative magnitude of the aggregate elasticities of substitution.
Finally, from a methodological point of view, this study is related to Baqaee and Farhi (2019) and Oberfield and Raval (2021), who develop a framework to characterize and estimate the aggregate elasticity of substitution in an economy consisting of multiple sectors. Relative to these studies, we consider a dynamic economy in which new capital accumulates more rapidly as technology progresses. We derive analytical expressions that decomposes changes in relative wages and income shares into components attributable to factor inputs and quantifies the general equilibrium effects of capital-embodied technological change on relative wages and income shares. We further extend the framework to allow for endogenous labor supply and imperfect labor mobility across sectors. As in the wage inequality literature, we focus on the value-added output.
This study contributes to these strands of literature in three ways. First, we provide the first estimates of the elasticities of substitution between ICCT equipment and four types of labor disaggregated by gender and skill. These estimates enable us to explain the observed differences in the direction and magnitude of changes in the gender wage gap by skill, the skill wage gap by gender, and labor shares by gender and skill in terms of technological change embodied in ICCT equipment. Second, we analytically derive and quantitatively evaluate the contributions of specific factor inputs to changes in the relative wages and income shares of different types of labor in OECD countries. Our results suggest that technological change embodied in ICCT equipment underlies gender-biased technological change, as well as skill-biased technological change noted in the literature. Finally, we analytically derive and quantitatively evaluate the general equilibrium effects of technological change embodied in ICCT equipment on the relative wages and income shares of different types of labor in OECD countries. Our results capture the effects of capital-embodied technological change more comprehensively because the elasticity of capital–labor substitution is allowed to vary within and across sectors.
3 Theory
We develop a unified framework to evaluate the effects of capital-embodied technological change on, as well as the contributions of factor inputs to, the relative wages and income shares of different types of labor. Our framework builds upon the seminal work of Greenwood, Hercowitz and Krusell (1997) on investment-specific technological change. We consider a multi-sector economy in which multiple consumption and investment goods are produced in each sector from multiple types of capital and multiple types of labor, whereas Greenwood, Hercowitz and Krusell (1997) consider a two-sector economy in which consumption and investment goods are separately produced in different sectors from two types of capital and one type of labor. We first show how the relative wages and income shares of different types of labor can change in response to each input, holding all other inputs constant. We then show how they can change in response to a fundamental determinant, technological change, allowing all inputs to vary. Furthermore, we extend the framework to allow for endogenous labor supply and imperfect labor mobility across sectors in Appendix A. Our theoretical results do not require specifying any aggregate or sectoral production function.
3.1 Environment
3.1.1 Household
The representative household has preferences over sequences of consumption $\{c_{ft}\}_{t=0}^{\infty}$ characterized by
$\sum_{t=0}^{\infty}\beta^{t}\sum_{f=1}^{F_{\ell}}\mu_{f}\mathcal{U}\left(c_{ft}\right),$ (1)
where $\beta\in(0,1)$ is a discount factor, $\mu_{f}$ is the share of individuals of type $f\in\{1,\ldots,F_{\ell}\}$, and $\mathcal{U}(\cdot)$ is an instantaneous utility function. Aggregate consumption by an individual of type $f$ in period $t$ is a constant elasticity of substitution (CES) composite of goods produced in each sector:
$c_{ft}=\left(\sum_{n=1}^{N}\theta_{n}c_{fnt}^{\eta}\right)^{\frac{1}{\eta}},$ (2)
where $c_{fnt}$ is consumption goods produced in sector $n$. The share parameters $\theta_{n}$ lie between zero and one, and the substitution parameter $\eta$ is less than one.
The household is endowed with an initial capital stock $k_{f0}$. The individual of type $f$ is endowed with time $l_{ft}$ in each period. Capital accumulates according to the law of motion:
$k_{f,t+1}=\left(1-\delta_{f}\right)k_{ft}+x_{ft},$ (3)
where $x_{ft}$ is investment in capital $f\in\{1,\ldots,F_{k}\}$, and $\delta_{f}\in(0,1)$ is a depreciation rate. Aggregate investment in capital $f$ is also a CES composite of goods produced in each sector:
$x_{ft}=q_{ft}\left(\sum_{n=1}^{N}\theta_{n}x_{fnt}^{\eta}\right)^{\frac{1}{\eta}},$ (4)
where $x_{fnt}$ is investment goods produced in sector $n$, and $q_{ft}$ is investment-specific technology. In each period, the household faces the budget constraint:
$\sum_{n=1}^{N}p_{nt}\sum_{f=1}^{F_{\ell}}\mu_{f}c_{fnt}+\sum_{n=1}^{N}p_{nt}\sum_{f=1}^{F_{k}}x_{fnt}=\sum_{f=1}^{F_{\ell}}\mu_{f}w_{ft}l_{ft}+\sum_{f=1}^{F_{k}}r_{ft}k_{ft},$ (5)
where $p_{nt}$ is the price of goods produced in sector $n$, $w_{ft}$ is the wages of labor $f$, and $r_{ft}$ is the rental price of capital $f$. The supply of labor is assumed to be exogenously given, but this assumption will be relaxed in Appendix A.
The household maximizes its utility by equalizing consumption across individuals: $c_{fnt}=c_{nt}$ and $c_{ft}=c_{t}$ for all $f$. The expenditure share of final goods $n$ in period $t$ is then defined as $\zeta_{nt}=p_{nt}y_{nt}/\sum_{n=1}^{N}p_{nt}y_{nt}=p_{nt}y_{nt}/p_{t}y_{t}$, where the aggregate price of consumption goods is $p_{t}=(\sum_{n=1}^{N}\theta_{n}^{1/(1-\eta)}p_{nt}^{-\eta/(1-\eta)})^{-(1-\eta)/\eta}$, and the aggregate output is $y_{t}=c_{t}+\sum_{f=1}^{F_{k}}(x_{ft}/q_{ft})$. The expenditure share can be derived as
$\zeta_{nt}=\text{$\theta_{n}^{\frac{1}{1-\eta}}\left(\frac{p_{nt}}{p_{t}}\right)^{-\frac{\eta}{1-\eta}}$}.$ (6)
As in Greenwood, Hercowitz and Krusell (1997), investment-specific technology can be measured by the ratio of the aggregate price of investment goods to the aggregate price of consumption goods.
$\frac{1}{q_{ft}}=\frac{p_{ft}}{p_{t}}.$ (7)
where the aggregate price of investment goods is $p_{ft}=(1/q_{ft})(\sum_{n=1}^{N}\theta_{n}^{1/(1-\eta)}p_{nt}^{-\eta/(1-\eta)})^{-(1-\eta)/\eta}$.
3.1.2 Firms
The firm in sector $n$ in period $t$ maximizes its profits:
$p_{nt}y_{nt}-\sum_{f=1}^{F_{\ell}}w_{ft}\ell_{fnt}-\sum_{f=1}^{F_{k}}r_{ft}k_{fnt}$ (8)
subject to a constant returns to scale technology:
$y_{nt}=A_{nt}\mathcal{F}_{n}\left(\ell_{1nt},\ldots,\ell_{F_{\ell}nt},k_{1nt},\ldots,k_{F_{k}nt}\right),$ (9)
where $y_{nt}$ is the real output in sector $n$ in period $t$, $\ell_{fnt}$ is the quantity of labor $f\in\{1,\ldots,F_{\ell}\}$, $k_{fnt}$ is the quantity of capital $f\in\{1,\ldots,F_{k}\}$, $A_{nt}$ is sector-specific factor-neutral technology, and $\mathcal{F}_{n}(\cdot)$ is a sectoral production function.
Below, to avoid notational clutter, we let $\ell_{fnt}$ denote not only labor for $f\in\{1,\ldots,F_{\ell}\}$ but also capital for $f\in\{F_{\ell}+1,\ldots,F\}$ and $w_{ft}$ denote not only wages for $f\in\{1,\ldots,F_{\ell}\}$ but also the rental price of capital for $f\in\{F_{\ell}+1,\ldots,F\}$.
The income share of factor $f$ in sector $n$ in period $t$ is defined as $\lambda_{\ell_{f}nt}=w_{ft}\ell_{fnt}/p_{nt}y_{nt}$. When the production function is homogeneous of degree one, the profits (8) can be written as $p_{nt}y_{nt}-y_{nt}\mathcal{\widetilde{C}}_{n}(w_{1t},\ldots,w_{Ft})/A_{nt}$, where $\mathcal{\mathcal{\widetilde{C}}}_{n}(\cdot)/A_{nt}$ is a unit cost function in sector $n$ in period $t$, and the demand for factor $f$ in sector $n$ in period $t$ can be written as $\ell_{fnt}=y_{nt}\widetilde{\mathcal{G}}_{fn}(w_{1t},\ldots,w_{Ft})/A_{nt}$, where $\widetilde{\mathcal{G}}_{fn}(\cdot)/A_{nt}$ is the demand for factor $f$ per unit output in sector $n$ in period $t$. By Shephard’s lemma, $\lambda_{\ell_{f}nt}=\partial\ln\mathcal{\mathcal{\widetilde{C}}}_{n}(w_{1t},\ldots,w_{Ft})/\partial\ln w_{ft}$ and $\partial\ln\lambda_{\ell_{f}nt}/\partial\ln w_{gt}=\partial\ln\widetilde{\mathcal{G}}_{fn}\left(w_{1t},\ldots,w_{Ft}\right)/\partial\ln w_{gt}-\lambda_{\ell_{g}nt}+I_{\left(f,g\right)},$where $I_{(f,g)}$ is the $fg$-th element of the $F\times F$ identity matrix. While $\lambda_{\ell_{f}nt}$ can be observed directly from data, $\partial\ln\lambda_{\ell_{f}nt}/\partial\ln w_{g}$ can be estimated after $\mathcal{F}_{n}(\cdot)$ is specified and $\widetilde{\mathcal{G}}_{fn}(\cdot)$ is derived.
3.1.3 Equilibrium
Given a set of technologies $A_{nt}$, $q_{ft}$} and an initial capital stock $k_{f0}$, a competitive equilibrium is a set of prices $p_{nt}$, $w_{ft}$, $r_{ft}$} and a set of quantities $c_{fnt}$, $x_{fnt}$, $\ell_{fnt}$, $k_{fnt}$} such that: (i) given prices, the representative household chooses $c_{fnt}$, $x_{fnt}$, and $k_{f,t+1}$ to maximize the utility (1) subject to the budget constraint (5), the capital law of motion (3), and the consumption and investment aggregators (2) and (4), (ii) given prices, each firm chooses $\ell_{fnt}$ and $k_{fnt}$ to maximize its profits (8) subject to the technology (9), (iii) goods markets clear:
$c_{nt}+\sum_{f=1}^{F_{k}}x_{fnt}=y_{nt},$ (10)
and (iv) factor markets clear:
$\sum_{n=1}^{N}\ell_{fnt}=\mu_{f}l_{ft}\qquad\text{and}\qquad\sum_{n=1}^{N}k_{fnt}=k_{ft}.$ (11)
We consider the competitive equilibrium that satisfies conditions (ii) and (iii) throughout this paper. When we extend this model to allow for endogenous labor supply or imperfect labor mobility, we modify condition (i) or (iv) as described in Appendix A.
3.2 Aggregate elasticity of substitution
The aggregate elasticity of substitution can be expressed in terms of the aggregate production or cost function ($\mathcal{F}$ or $\mathcal{C}$). [^3] Baqaee and Farhi (2019) describe the aggregate production function as a maximal attainable output subject to consumer’s preferences, producer’s technology, and resource constraints and the aggregate cost function as a minimal attainable cost of production subject to consumer’s preferences, producer’s technology, and resource constraints in a static economy, where all factors are exogenously given and perfectly mobile across sectors. The Morishima (1967) elasticities of substitution between factors $f$ and $g$ are defined as
$\frac{1}{\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{F}}}=\frac{\partial\ln\left(\left.\frac{\partial\mathcal{F}}{\partial\ell_{f}}\right/\frac{\partial\mathcal{F}}{\partial\ell_{g}}\right)}{\partial\ln\left(\left.\ell_{g}\right/\ell_{f}\right)},$
$\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}=\frac{\partial\ln\left(\left.\frac{\partial\mathcal{C}}{\partial w_{f}}\right/\frac{\partial\mathcal{C}}{\partial w_{g}}\right)}{\partial\ln\left(\left.w_{g}\right/w_{f}\right)},$
where $\ell_{f}=\sum_{n=1}^{N}\ell_{fn}$ for $f\in\{1,\ldots,F\}$. [^4] The Morishima elasticity of substitution is a measure of curvature of the isoquant in line with the original concept by Hicks (1932) and a sufficient statistic for measuring the contributions of factor prices to relative factor shares, as noted by Blackorby and Russell (1989). The partial derivative of the log of the production (cost) function with respect to the log of the factor quantity (price) equals the factor share of income in the aggregate economy, defined as $\Lambda_{\ell_{f}}=w_{f}\ell_{f}/py$. The aggregate elasticities of substitution can be derived by looking at the extent to which the factor shares of income vary according to the quantity or price of factors. Throughout this and next two subsections, we suppress the time subscript $t$ for notational simplicity. The following lemma recasts the results of Baqaee and Farhi (2019) for value-added output. [^5] Baqaee and Farhi (2019) derive expressions for the elasticities of substitution in the aggregate production and cost functions when a micro-level production function takes a nested CES form, and firms are connected through input–output linkages (Propositions 3 and 8). In the appendix, they refine these expressions without imposing any functional form on the micro-level production function. Oberfield and Raval (2021) derive expressions for the elasticity of substitution in the aggregate cost function both when the micro-level production function takes a nested CES form (Proposition 2) and when no functional form is imposed on the micro-level production function (Proposition 2’).
Lemma 1.
The aggregate elasticities of substitution between factors $f$ and $g$ are given by
$\frac{1}{\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{F}}}=$ $\Psi_{\left(f,g\right)}^{\ell}-\Psi_{\left(g,g\right)}^{\ell}+1,$ (12)
$\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}=$ $\Psi_{\left(f,g\right)}^{w}-\Psi_{\left(g,g\right)}^{w}+1,$ (13)
where $\Psi_{(f,g)}^{\ell}$ is the $fg$-th element of the $F\times F$ matrix $\Psi^{\ell}=-(I-\Psi^{w})^{-1}\Psi^{w}$, and $\Psi^{w}$ is the $F\times F$ matrix whose $fg$-th element is given by $\Psi_{(f,g)}^{w}=\sum_{n=1}^{N}(\zeta_{n}\lambda_{\ell_{f}n}/\Lambda_{\ell_{f}})[\partial\ln\lambda_{\ell_{f}n}/\partial\ln w_{g}+[1-1/(1-\eta)](\lambda_{\ell_{g}n}-\Lambda_{\ell_{g}})]$.
Appendix B contains the proofs of this lemma and all propositions. The lemma and propositions apply not only to sector-level production functions but also to firm- and establishment-level production functions.
The aggregate elasticities of substitution do not generally coincide with the sectoral elasticities of substitution for two reasons. First, the elasticities of substitution in one sector may differ from those in another sector. Second, factor intensities in one sector may differ from those in another sector. The aggregate elasticities of substitution depend not only on substitution within sectors but also on reallocation across sectors.
Corollary 1.
The aggregate elasticity of substitution (13) can be expressed as a weighted average of the elasticities of substitution in production and consumption:
$\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}=\left[\sum_{n=1}^{N}\left(\frac{\ell_{fn}}{\ell_{f}}\right)\epsilon_{\ell_{f}\ell_{g}n}^{\mathcal{C}}+\sum_{n=1}^{N}\left(\frac{\ell_{gn}}{\ell_{g}}-\frac{\ell_{fn}}{\ell_{f}}\right)\sum_{f\neq g}^{F}\lambda_{\ell_{f}n}\epsilon_{\ell_{f}\ell_{g}n}^{\mathcal{C}}\right]+\left[\sum_{n=1}^{N}\left(\frac{\ell_{gn}}{\ell_{g}}-\frac{\ell_{fn}}{\ell_{f}}\right)\lambda_{\ell_{g}n}\frac{1}{1-\eta}\right],$ (14)
where the elasticity of substitution in the sectoral cost function is defined as $\epsilon_{\ell_{f}\ell_{g}n}^{\mathcal{C}}=\partial\ln[(\partial\mathcal{C}_{n}/\partial w_{f})$ $(\partial\mathcal{C}_{n}/\partial w_{g})]/\partial\ln(w_{g}/w_{f})$.
The first square bracket captures the extent to which the producer’s optimal mix of factors changes in response to relative factor prices. If factor intensities are equal across sectors (i.e., $\ell_{fn}/\ell_{gn}=\ell_{fn^{\prime}}/\ell_{gn^{\prime}}=\ell_{f}/\ell_{g}$ for $n\neq n^{\prime}$), the aggregate elasticity (14) is the weighted average of the sectoral elasticities; that is, $\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}=\sum_{n=1}^{N}(\ell_{fn}/\ell_{f})\epsilon_{\ell_{f}\ell_{g}n}^{\mathcal{C}}$, where the weight is the quantity share of factor $f$ in each sector. Furthermore, if the elasticities of substitution are equal across sectors (i.e., $\epsilon_{\ell_{f}\ell_{g}n}^{\mathcal{C}}=\epsilon_{\ell_{f}\ell_{g}n^{\prime}}^{\mathcal{C}}$ for $n\neq n^{\prime}$), the aggregate elasticity coincides with the sectoral elasticity. The second square bracket captures the extent to which the consumer’s optimal mix of goods (i.e., the size of each sector) changes in response to relative factor prices. An increase in the factor price raises the price of goods due to an increase in the marginal cost of production, which in turn reduces consumption. An increase in the price of factor $g$ reduces the equilibrium quantity of labor $g$ relative to factor $f$ (or equivalently raises the equilibrium quantity of factor $f$ relative to factor $g$) in sectors more intensive in factor $g$ than in factor $f$ (i.e., $\ell_{fn}/\ell_{f}<\ell_{gn}/\ell_{g}$). [^6] Our representation in equation (14) differs from Baqaee and Farhi (2019), who consider a production function with capital, labor, and intermediate inputs and underscore input–output linkages, and from Oberfield and Raval (2021), who consider a production function with capital, labor, and material inputs and focus on the elasticity of substitution between capital and labor.
The aggregate elasticities (12) and (13) can be expressed in a simple form if production requires only one type of capital and one type of labor, as described in Corollary 5 in Appendix C.
3.3 Contributions of factor inputs
3.3.1 Labor shares
The aggregate elasticities of substitution, $\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{F}}$ ($\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}$), are sufficient statistics for measuring the contributions of factor quantities (prices) to the factor shares of income if the factor shares of income are known. We denote by $\Delta\ln z$ the log change of variable $z$ from one period to another.
Proposition 1.
Changes in factor shares are given by
$\Delta\ln\Lambda_{\ell_{f}}=-\sum_{g\neq f}^{F}\left(\frac{1-\epsilon_{\ell_{g}\ell_{f}}^{\mathcal{F}}}{\epsilon_{\ell_{g}\ell_{f}}^{\mathcal{F}}}\right)\Lambda_{\ell_{g}}\Delta\ln\ell_{f}+\sum_{g\neq f}^{F}\left(\frac{1-\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{F}}}{\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{F}}}\right)\Lambda_{\ell_{g}}\Delta\ln\ell_{g}\\ +\sum_{g\neq f}^{F}\sum_{h\neq f,g}^{F}\left(\frac{\epsilon_{\ell_{h}\ell_{g}}^{\mathcal{F}}-\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{F}}}{\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{F}}\epsilon_{\ell_{h}\ell_{g}}^{\mathcal{F}}}\right)\Lambda_{\ell_{h}}\Delta\ln\ell_{g}+\psi_{\left(f\right)}^{\ell}$ (15)
or
$\Delta\ln\Lambda_{\ell_{f}}=\sum_{g\neq f}^{F}\left(1-\epsilon_{\ell_{g}\ell_{f}}^{\mathcal{C}}\right)\Lambda_{\ell_{g}}\Delta\ln\left(\frac{w_{f}}{p}\right)-\sum_{g\neq f}^{F}\left(1-\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}\right)\Lambda_{\ell_{g}}\Delta\ln\left(\frac{w_{g}}{p}\right)\\ -\sum_{g\neq f}^{F}\sum_{h\neq f,g}^{F}\left(\epsilon_{\ell_{h}\ell_{g}}^{\mathcal{C}}-\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}\right)\Lambda_{\ell_{h}}\Delta\ln\left(\frac{w_{g}}{p}\right)+\psi_{\left(f\right)}^{w},$ (16)
where $\psi_{(f)}^{\ell}$ is the $f$-th element of the $F\times 1$ vector $\psi^{\ell}=(I-\Psi^{w})^{-1}\psi^{w}$, and $\psi^{w}$ is the $F\times 1$ vector whose $f$-th element is given by $\psi_{(f)}^{w}=-[1-1/(1-\eta)]\sum_{n=1}^{N}(\zeta_{n}\lambda_{\ell_{f}n}/\Lambda_{\ell_{f}})(\Delta\ln A_{n}-\sum_{m=1}^{N}\zeta_{m}\Delta\ln A_{m})$.
We refer to the first two terms in equation (15) as the relative quantity effect, the first two terms in equation (16) as the relative price effect, and the third term in equations (15) and (16) as the factor complementarity effect. The direction of the relative price and quantity effects on the income share of factor $f$ depends on whether the elasticity of substitution between factors $f$ and $g$ exceeds one, whereas the direction of the factor complementarity effect on the income share of factor $f$ depends on whether factor $g$ is more complementary to factor $f$ than to factor $h$.
If production requires only one type of capital and one type of labor (or equivalently, if all types of labor are perfectly substitutable, and all types of capital are perfectly substitutable), the factor complementarity effect may disappear. In this case, the direction and magnitude of changes in the labor share are determined by the elasticity of substitution between capital and labor and the degree of capital deepening, as described in Corollary 6 in Appendix C.
However, when multiple types of capital and labor are used to produce goods and services, the labor share can change not only due to the relative price or quantity effect but also due to the factor complementarity effect. Equations (15) and (16) show how the income share of factor $f$ can change in response to its own price or quantity change and other price or quantity change in such a general case. On the one hand, a rise in the price or a fall in the quantity of factor $f$ causes a rise in the relative price of factor $f$ and a fall in the relative quantity of factor $f$. A rise in the price (a fall in the quantity) of factor $f$ can result in a decline in the income share of factor $f$ if the latter exceeds the former, in which case $\epsilon_{\ell_{g}\ell_{f}}^{\mathcal{C}}>1$ ($\epsilon_{\ell_{g}\ell_{f}}^{\mathcal{F}}>1$). On the other hand, a fall in the price or a rise in the quantity of factor $g$ causes a rise in the relative price of factor $f$ and a fall in the relative quantity of factor $f$. A fall in the price (a rise in the quantity) of factor $g$ can result in a decline in the income share of factor $f$ if the latter exceeds the former, in which case $\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}>1$ ($\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{F}}>1$). A fall in the price (a rise in the quantity) of factor $g$ can further reduce the income share of factor $f$ if factor $g$ is more substitutable with factor $f$ than with factor $h$, in which case $\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}>\epsilon_{\ell_{h}\ell_{g}}^{\mathcal{C}}$ ($\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{F}}>\epsilon_{\ell_{h}\ell_{g}}^{\mathcal{F}}$). The last term arises from sector-specific disembodied technological change.
When production requires multiple types of labor, the rate of change in the labor share can be expressed as the weighted average of the rates of changes in the income shares of different types of labor:
$\Delta\ln\Lambda_{\ell}=\sum_{f=1}^{F_{\ell}}\left(\frac{w_{f}\ell_{f}}{\sum_{f=1}^{F_{\ell}}w_{f}\ell_{f}}\right)\Delta\ln\Lambda_{\ell_{f}},$ (17)
where $\Delta\ln\Lambda_{\ell_{f}}$ is given by equation (15), and the weight is the wage bill share of each type of labor. Given the fact that the majority of labor is male unskilled, the largest weight is assigned to the rate of change in the income share of male unskilled labor.
3.3.2 Relative wages
The aggregate elasticities of substitution, $\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{F}}$ ($\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}$), are sufficient statistics for measuring the contributions of factor quantities (prices) to the relative wages of (relative demand for) different types of labor.
Proposition 2.
Changes in relative factor prices are given by
$\Delta\ln\left(\frac{w_{f}}{w_{g}}\right)=-\sum_{h\neq f,g}\left(\frac{\epsilon_{\ell_{f}\ell_{h}}^{\mathcal{F}}-\epsilon_{\ell_{g}\ell_{h}}^{\mathcal{F}}}{\epsilon_{\ell_{f}\ell_{h}}^{\mathcal{F}}\epsilon_{\ell_{g}\ell_{h}}^{\mathcal{F}}}\right)\Delta\ln\ell_{h}-\frac{1}{\epsilon_{\ell_{g}\ell_{f}}^{\mathcal{F}}}\Delta\ln\ell_{f}+\frac{1}{\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{F}}}\Delta\ln\ell_{g}+\left(\psi_{\left(f\right)}^{\ell}-\psi_{\left(g\right)}^{\ell}\right).$ (18)
Changes in relative factor demand are given by
$\Delta\ln\left(\frac{\ell_{f}}{\ell_{g}}\right)=\sum_{h\neq f,g}\left(\epsilon_{\ell_{f}\ell_{h}}^{\mathcal{C}}-\epsilon_{\ell_{g}\ell_{h}}^{\mathcal{C}}\right)\Delta\ln\left(\frac{w_{h}}{p}\right)-\epsilon_{\ell_{g}\ell_{f}}^{\mathcal{C}}\Delta\ln\left(\frac{w_{f}}{p}\right)+\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}\Delta\ln\left(\frac{w_{g}}{p}\right)+\left(\psi_{\left(f\right)}^{w}-\psi_{\left(g\right)}^{w}\right).$ (19)
We refer to the first term as the factor complementarity effect and the second and third terms in equation (18) as the relative quantity effect again. The direction of the factor complementarity effect on the wages of factor $f$ relative to factor $g$ depends on whether factor $h$ is more complementary to factor $f$ than to factor $g$, whereas the direction of the relative quantity effect depends on whether each factor increases. Equation (18) is a multi-sector, multi-factor general equilibrium version of the model of the race between technology and skills, which nests the standard models of skill differentials, as described in Corollaries 7 and 8 in Appendix C.
Equation (18) shows how gender and skill premia can vary according to the shift in demand due to progress in ICCT and the shift in supply due to advances in educational attainment and female employment. The first term captures the shift in demand. The wages of factor $f$ relative to factor $g$ can increase with a rise in factor $h$ if factor $h$ is more complementary to factor $f$ than to factor $g$ (i.e., $\epsilon_{\ell_{g}\ell_{h}}^{\mathcal{F}}>\epsilon_{\ell_{f}\ell_{h}}^{\mathcal{F}}$). In particular, if ICCT capital is more complementary to skilled (female) labor than to unskilled (male) labor, the skill (gender) wage gap increases (decreases) with a rise in ICCT capital. The magnitude of this factor complementarity effect is directly proportional to the difference between $\epsilon_{\ell_{g}k_{i}}^{\mathcal{F}}$ and $\epsilon_{\ell_{f}k_{i}}^{\mathcal{F}}$, as well as the rate of increase in $k_{i}$. The second and third terms capture the shift in supply. The wages of factor $f$ relative to factor $g$ can decrease with a rise (fall) in the quantity of factor $f$ ($g$). The magnitude of this relative quantity effect is inversely proportional to $\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{F}}$ and $\epsilon_{\ell_{g}\ell_{f}}^{\mathcal{F}}$ and directly proportional to the rate of increase in $\ell_{f}$ relative to $\ell_{g}$. Given the fact that the rate of increase in skilled (female) labor is greater than that in unskilled (male) labor, the relative quantity effect should be negative (positive) for the skilled–unskilled (male–female) wage gap. Consequently, the factor complementarity effect and the relative quantity effect would work in opposite directions. Whether the skill (gender) wage gap eventually widens (narrows) depends on the relative magnitude of these two effects.
When production requires four or more types of labor, it may be of interest to examine the sources of change in the average wage for one group relative to that for another group. The average wage of labor $h$ in group $H$ is defined as $\overline{w}_{H}=\sum_{h\in H}(\ell_{h}/\sum_{h\in H}\ell_{h})w_{h}$. The rate of change in the average wage for group $H$ relative to that for group $U$ can then be expressed as
$\Delta\ln\left(\frac{\overline{w}_{H}}{\overline{w}_{U}}\right)=\sum_{h\in H}\left(\frac{w_{h}\ell_{h}}{\sum_{h\in H}w_{h}\ell_{h}}\Delta\ln\Lambda_{\ell_{h}}-\frac{\ell_{h}}{\sum_{h\in H}\ell_{h}}\Delta\ln\ell_{h}\right)\\ -\sum_{u\in U}\left(\frac{w_{u}\ell_{u}}{\sum_{u\in U}w_{u}\ell_{u}}\Delta\ln\Lambda_{\ell_{u}}-\frac{\ell_{u}}{\sum_{u\in U}\ell_{u}}\Delta\ln\ell_{u}\right),$ (20)
where $\Delta\ln\Lambda_{\ell_{f}}$ is given by equation (15). The rate of change in the ratio of average wages between two groups equals the difference between the weighted average of the rates of change in the income shares of different types of labor, with each type weighted by its wage bill share, and the weighted average of the rates of change in the quantities of different types of labor, with each type weighted by its quantity share.
3.4 Effects of technological change
The analysis thus far is useful for isolating the sources of changes in the relative wages and income shares of different types of labor. We can quantify the extent to which each factor contributes to changes in relative wages and labor shares using equations (15) and (18) for given values of the aggregate elasticities of substitution. However, if the equilibrium prices and quantities of factors can change in response to technological change, those equations are insufficient to fully capture the effects of technological change. Here, we derive the analytical expressions for changes in the equilibrium relative wages and income shares of different types of labor due to capital-embodied technological change and sector-specific disembodied technological change. Throughout this and next subsections, we focus on steady-state equilibria in which all variables are constant over time and examine changes before and after information revolution. We denote by $d\ln z$ the log change of variable $z$ from one steady state to another.
Proposition 3.
Changes in relative wages due to technological change are given by
$d\ln\left(\frac{w_{f}}{w_{g}}\right)=\sum_{h=1}^{F_{k}}\left(\Upsilon_{\left(f,h\right)}^{q}-\Upsilon_{\left(g,h\right)}^{q}\right)d\ln q_{h}+\sum_{n=1}^{N}\left(\Upsilon_{\left(f,n\right)}^{A}-\Upsilon_{\left(g,n\right)}^{A}\right)d\ln A_{n}\quad\text{for }f,g=1,\ldots,F_{\ell},$ (21)
where $\Upsilon_{\left(f,g\right)}^{q}$ is the $fg$-th element of the $F_{\ell}\times F_{k}$ matrix $\Upsilon^{q}$, and $\Upsilon_{\left(f,n\right)}^{A}$ is the $fn$-th element of the $F_{\ell}\times N$ matrix $\Upsilon^{A}$. The matrices $\Upsilon^{q}$ and $\Upsilon^{A}$ are described in Appendix B. Changes in labor shares due to technological change are given by
$d\ln\Lambda_{\ell_{f}}=\sum_{h=1}^{F_{k}}\left\{\sum_{g\neq f}^{F}\left(1-\epsilon_{\ell_{g}\ell_{f}}^{\mathcal{C}}\right)\Lambda_{\ell_{g}}\Upsilon_{\left(f,h\right)}^{q}+\sum_{g\neq f}^{F_{\ell}}\left[-\left(1-\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}\right)\Lambda_{\ell_{g}}+\sum_{s\neq f,g}^{F}\left(\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}-\epsilon_{\ell_{s}\ell_{g}}^{\mathcal{C}}\right)\Lambda_{\ell_{s}}\right]\Upsilon_{\left(g,h\right)}^{q}\right.\\ \left.-\left[-\left(1-\epsilon_{\ell_{f}k_{h}}^{\mathcal{C}}\right)\Lambda_{k_{h}}+\sum_{s\neq f,h}^{F}\left(\epsilon_{\ell_{f}k_{h}}^{\mathcal{C}}-\epsilon_{\ell_{s}k_{h}}^{\mathcal{C}}\right)\Lambda_{\ell_{s}}\right]\right\}d\ln q_{h}+\sum_{n=1}^{N}\left\{\Phi_{\left(f,n\right)}^{w}+\sum_{g\neq f}^{F}\left(1-\epsilon_{\ell_{g}\ell_{f}}^{\mathcal{C}}\right)\Lambda_{\ell_{g}}\Upsilon_{\left(f,n\right)}^{A}\right.\\ \left.+\sum_{g\neq f}^{F_{\ell}}\left[-\left(1-\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}\right)\Lambda_{\ell_{g}}+\sum_{s\neq f,g}^{F}\left(\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}-\epsilon_{\ell_{s}\ell_{g}}^{\mathcal{C}}\right)\Lambda_{\ell_{s}}\right]\Upsilon_{\left(g,n\right)}^{A}\right\}d\ln A_{n}\quad\text{for }f=1,\ldots,F_{\ell},$ (22)
where $\Phi^{w}$ is the $F\times N$ matrix whose $fn$-th element is given by $\Phi_{\left(f,n\right)}^{w}=-[1-1/(1-\eta)]\zeta_{n}(\lambda_{\ell_{f}n}-\Lambda_{\ell_{f}})/\Lambda_{\ell_{f}}$.
Equations (21) and (22) express how the relative wages and income shares of different types of labor can change in response to capital-embodied technological change and sector-specific disembodied technological change in a multi-sector, multi-factor general equilibrium model of investment-specific technological change. The effect of capital-embodied technological change on relative wages, which is represented by the first term in equation (21), is related to the factor complementarity effect in equation (18). The effect of capital-embodied technological change on labor shares, which is represented by the first term in equation (22), is related to the relative price effect and the factor complementarity effect in equation (16). The first term in equation (22) can be divided into two parts, one in the first line and the other in the second line. The two parts capture the effects through changes in wages and rental prices, respectively. Most elements of the matrix $\Upsilon^{q}$ consist of the aggregate elasticities of substitution $\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}$, and the others consist of the factor shares of income $\Lambda_{\ell_{f}}$. Equations (21) and (22) show that the aggregate elasticities of substitution are sufficient statistics for quantifying the effects of technological change on the relative wages and income shares of different types of labor if the income shares of each factor are known.
We now consider two examples to demonstrate how the direction and magnitude of the effects of technological change depend on the aggregate elasticities of substitution among different types of capital and labor.
Corollary 2.
Suppose that production requires only one type of capital, $k$, and one type of labor, $\ell$. Equation (22) becomes the Karabarbounis and Neiman (2014) model.
$d\ln\Lambda_{\ell}=\left(1-\epsilon\right)\left(\frac{1-\Lambda_{\ell}}{\Lambda_{\ell}}\right)d\ln q+\sum_{n=1}^{N}\zeta_{n}\left[\left(1-\epsilon\right)\left(\frac{1-\Lambda_{\ell}}{\Lambda_{\ell}}\right)+\left(1-\frac{1}{1-\eta}\right)\left(\frac{\Lambda_{\ell}-\lambda_{\ell n}}{\Lambda_{\ell}}\right)\right]d\ln A_{n}.$
This equation implies that the direction of the effect of capital-embodied technological progress depends entirely on whether the aggregate elasticity of substitution between capital and labor exceeds one. The labor share can decline with capital-embodied technological progress if the aggregate elasticity exceeds one, and the rate of decline becomes greater as the aggregate elasticity increases. The magnitude of the effect also becomes greater as the labor share decreases. The effect of sector-specific disembodied technological change can be divided into two parts. The first part in the square bracket takes the same form as the first term and captures the effect through changes in wages. The second part arises from the direct effect denoted by $\psi^{w}$ in equation (16) and is unlikely to be large unless factor shares differ widely across sectors. Thus, the relative magnitude of the first and second terms depends on that of the two types of technological change.
Corollary 3.
Suppose that production requires capital equipment, $k_{e}$, capital structures, $k_{s}$, skilled labor, $\ell_{h}$, and unskilled labor, $\ell_{u}$, the aggregate elasticities of substitution satisfy $\epsilon_{\ell_{h}k_{s}}^{\mathcal{C}}=\epsilon_{\ell_{u}k_{s}}^{\mathcal{C}}=1$ and $\epsilon_{\ell_{u}\ell_{h}}^{\mathcal{C}}-\epsilon_{\ell_{h}\ell_{u}}^{\mathcal{C}}=\epsilon_{\ell_{h}k_{e}}^{\mathcal{C}}-\epsilon_{\ell_{u}k_{e}}^{\mathcal{C}}$, and there is no technological change embodied in capital structure, as in Krusell et al. (2000). Changes in the skill premium due to technological change are given by
$d\ln\left(\frac{w_{h}}{w_{u}}\right)=\left(\frac{\epsilon_{\ell_{u}k_{e}}^{\mathcal{C}}-\epsilon_{\ell_{h}k_{e}}^{\mathcal{C}}}{\overline{\epsilon}_{\ell_{h}\ell_{u}}^{\mathcal{C}}}\right)\left(\frac{\Lambda_{k_{e}}+\Lambda_{\ell_{h}}+\Lambda_{\ell_{u}}}{\Lambda_{\ell_{h}}+\Lambda_{\ell_{u}}}\right)d\ln q_{e}\\ +\sum_{n=1}^{N}\left[-\left(\frac{\epsilon_{\ell_{u}k_{e}}^{\mathcal{C}}-\epsilon_{\ell_{h}k_{e}}^{\mathcal{C}}}{\overline{\epsilon}_{\ell_{h}\ell_{u}}^{\mathcal{C}}}\right)\left(\frac{\zeta_{n}}{\Lambda_{\ell_{h}}+\Lambda_{\ell_{u}}}\right)+\frac{1}{\overline{\epsilon}_{\ell_{h}\ell_{u}}^{\mathcal{C}}}\left(1-\frac{1}{1-\eta}\right)\left(\frac{\zeta_{n}\lambda_{\ell_{u}n}}{\Lambda_{\ell_{u}}}-\frac{\zeta_{n}\lambda_{\ell_{h}n}}{\Lambda_{\ell_{h}}}\right)\right]d\ln A_{n},$
where $\overline{\epsilon}_{\ell_{h}\ell_{u}}^{\mathcal{C}}=[\Lambda_{\ell_{h}}/(\Lambda_{\ell_{h}}+\Lambda_{\ell_{u}})]\epsilon_{\ell_{h}\ell_{u}}^{\mathcal{C}}+[\Lambda_{\ell_{u}}/(\Lambda_{\ell_{h}}+\Lambda_{\ell_{u}})]\epsilon_{\ell_{u}\ell_{h}}^{\mathcal{C}}$, which is known as the McFadden (1963) elasticity of substitution.
This equation is a multi-sector, multi-factor general equilibrium version of the Krusell, Ohanian, Ríos-Rull and Violante (2000) model. It implies that the direction and magnitude of the effect of capital-embodied technological progress on the skill wage gap depend on the presence and degree of capital–skill complementarity. The skill wage gap can increase with capital-embodied technological progress if capital equipment is more complementary to skilled labor than to unskilled labor (i.e., $\epsilon_{\ell_{u}k_{e}}^{\mathcal{F}}>\epsilon_{\ell_{h}k_{e}}^{\mathcal{F}}$). The rate of the increase becomes greater as the difference between the two elasticities or the income share of capital equipment increases. The effect of sector-specific disembodied technological change on the skill wage gap can be divided into two parts. The first part in the square bracket takes a similar form to the effect of capital-embodied technological change and captures the effect through a change in capital equipment. However, this has the opposite sign to that of the first term because sector-specific disembodied technological progress reduces the need for the use of capital equipment. The second part is related to the direct effect denoted by $\psi^{\ell}$ in equation (18). The sign of the first part is negative in the presence of capital–skill complementarity, whereas that of the second part is likely to be positive. In this setting, sector-specific disembodied technological progress is unlikely to account for the widening of the skill wage gap.
4 Data
We have shown in Propositions 1 to 3 that the aggregate elasticities of substitution and the factor shares of income determine the direction and magnitude of changes in relative wages and income shares of different types of labor attributable to (i) changes in specific inputs, under which other inputs are held constant, and (ii) changes in technologies, under which all inputs are allowed to vary. This section begins by describing the data sources, sample, and variables used to calculate the factor shares of income and to estimate the aggregate elasticities of substitution. It then documents trends in the relative prices, relative quantities, and income shares of factors. The section ends by looking at the substitutability between ICCT capital and the four types of labor.
4.1 Data sources, sample, and variables
The data used in the analysis are mainly from the EU KLEMS database, which collects detailed and internationally comparable information on the prices and quantities of capital and labor in major OECD countries. Our analysis is based on the March 2008 version because it contains the longest time series covering the 1980s, during which there were dramatic changes in technology and inequality. For the years 1980 to 2005, we include all countries and years in the sample for which required data are available. Our sample comprises 11 OECD countries: Australia, Austria, the Czech Republic, Denmark, Finland, Germany, Italy, Japan, the Netherlands, the United Kingdom, and the United States. The economy of each country is divided into five sectors: primary, manufacturing, utilities, construction, and services. The sample comprises four non-primary sectors and includes 258 country-year observations for each sector. [^7] On average in our sample, the male share of employment is higher in the construction and utilities sectors than in any subsector of manufacturing or services.
Labor is divided into skilled and unskilled labor, each of which is further divided into male and female labor. Skilled labor consists of workers who have completed college, and unskilled labor consists of workers who do not enter or complete college. We calculate wages by gender and skill by dividing the total labor compensation by the total hours worked. When we calculate wages and hours worked, we adjust for changes in the age and education composition of the labor force over time. Appendix D.1 provides the details of the adjustment procedure.
Capital is divided into ICCT and non-ICCT capital. ICCT capital consists of computing equipment, communications equipment, and software. Non-ICCT capital consists of transport equipment, other machinery and equipment, non-residential structures and infrastructures, residential structures, and other assets. We calculate the rental price of capital (Jorgenson, 1963), based on either the internal or external rate of return. The former approach has the advantage of maintaining consistency between national income and production accounts, whereas the latter approach has the advantage of not requiring the assumption of competitive markets. We employ the latter approach to examine the robustness of our results. Appendix D.2 provides the details of the calculation procedure.
All variables measured in monetary values are converted into U.S. dollars by the purchasing power parity index and deflated by the gross value-added deflator with 1995 as the base year. The purchasing power parity index is obtained from the 1997 Productivity Level Database.
In addition, the Barro and Lee (2013) database is used to calculate the population size by gender and skill. The collective bargaining coverage, the strictness of employment protection legislation, and the level of minimum wages are obtained from OECD.Stat. [^8] The collective bargaining coverage is measured by the percentage of employees with the right to bargain. The strictness of employment protection legislation is measured in terms of the regulation of individual and collective dismissals of workers on regular contracts and the regulation for hiring workers on temporary contracts.
4.2 Factor prices, quantities, and shares
During the period between the years 1980 and 2005, there was a difference in the trends of the gender wage gap between skilled and unskilled workers in OECD countries (Figure 1a). The male–female wage gap declined among unskilled workers, but not among skilled workers. At the same time, there was a difference in the trends of the skill wage gap between male and female workers in OECD countries (Figure 1b). The skilled–unskilled wage gap increased among male workers after a slight drop in the early 1980s, while it did not increase among female workers except for the period between the years 1987 and 1995. Taken together, the rate of decrease in the male–female wage gap was greater among unskilled workers than among skilled workers, while the rate of increase in the skilled–unskilled wage gap was greater among male workers than among female workers. This finding is perhaps surprising but not inconsistent with the fact that the rate of decline in the gender wage gap was greater at the middle or bottom than at the top of the wage distribution in the United States from the 1980s to the 2000s (Blau and Kahn, 2017). The patterns of changes in gender and skill premia are more visible in the United States (Figures 2a and 2b). [^9] Since the mid-2000s, the widening of the skill wage gap has largely stabilized, while the narrowing of the gender wage gap has decelerated (Heathcote, Perri, Violante and Zhang, 2023). Across most countries in our sample, the rate of decline in the gender wage gap tends to be greater among unskilled workers than among skilled workers, while the rate of increase in the skill wage gap tends to be greater among male workers than among female workers. [^10] This pattern is observed in Australia, the Czech Republic, Denmark, Germany, Japan, and the United States. For the rest of the countries, there is no notable difference in the magnitude of changes in the gender wage gap between skilled and unskilled workers and in the skill wage gap between male and female workers.
Figure 1: Gender and skill premia
(a) Male vs. female
(b) Skilled vs. unskilled
Notes: The wages of male skilled, female skilled, male unskilled, and female unskilled labor are denoted by $w_{mh}$, $w_{fh}$, $w_{mu}$, and $w_{fu}$, respectively. All the series are logarithmically transformed and normalized to zero in the year 1980. The 1980 values of the skilled gender wage gap, $w_{mh}/w_{fh}$, unskilled gender wage gap, $w_{mu}/w_{fu}$, male skill wage gap, $w_{mh}/w_{mu}$, and female skill wage gap, $w_{fh}/w_{fu}$, are 1.35, 1.43, 1.59, and 1.67, respectively.
Figure 2: Gender and skill premia in the United States
(a) Male vs. female
(b) Skilled vs. unskilled
Notes: The wages of male skilled, female skilled, male unskilled, and female unskilled labor are denoted by $w_{mh}$, $w_{fh}$, $w_{mu}$, and $w_{fu}$, respectively. All the series are logarithmically transformed and normalized to zero in the year 1980. The 1980 values of the skilled gender wage gap, $w_{mh}/w_{fh}$, unskilled gender wage gap, $w_{mu}/w_{fu}$, male skill wage gap, $w_{mh}/w_{mu}$, and female skill wage gap, $w_{fh}/w_{fu}$, are 1.49, 1.69, 1.39, and 1.57, respectively.
Turning to the relative quantities of labor, there was a decline in the quantity of male labor relative to female labor, and an increase in the quantity of skilled labor relative to unskilled labor. The rate of decline in the relative quantity of male labor was much greater for skilled labor than for unskilled labor (Figure 3a). At the same time, the rate of increase in the relative quantity of skilled labor was much greater for female labor than for male labor (Figure 3b). These observations suggest that the relative supply of labor alone could not explain the differences in the trends of the gender wage gap across skill groups and of the skill wage gap across gender groups.
Figure 3: Relative labor quantities
(a) Male vs. female
(b) Skilled vs. unskilled
Notes: The quantities of male skilled, female skilled, male unskilled, and female unskilled labor are denoted by $\ell_{mh}$, $\ell_{fh}$, $\ell_{mu}$, and $\ell_{fu}$, respectively. All the series are logarithmically transformed and normalized to zero in the year 1980. The 1980 values of the skilled male–female ratio, $\ell_{mh}/\ell_{fh}$, unskilled male–female ratio, $\ell_{mu}/\ell_{fu}$, male skilled–unskilled ratio, $\ell_{mh}/\ell_{mu}$, and female skilled–unskilled ratio, $\ell_{fh}/\ell_{fu}$, are 4.66, 1.77, 0.15, and 0.09, respectively.
Figure 4: ICCT and non-ICCT capital
(a) Investment prices
(b) Quantities
Notes: The investment prices of ICCT and non-ICCT capital are denoted by $p_{i}$ and $p_{o}$, respectively, and the output price is denoted by $p$. The quantities of ICCT and non-ICCT capital are denoted by $k_{i}$ and $k_{o}$, respectively. All the series are logarithmically transformed and normalized to zero in the year 1980.
The trends in the rental prices of capital differ significantly between ICCT and non-ICCT capital (Figure 4a). The rental price of ICCT capital fell dramatically, but that of non-ICCT capital remained almost unchanged. Meanwhile, the quantities of both ICCT and non-ICCT capital increased. However, the rate of increase in ICCT capital was far greater than that in non-ICCT capital (Figure 4b). These observations indicate substantial progress in ICCT during the period. [^11] Since the mid-2000s, the deepening of ICT capital has slowed markedly, while the decline in ICT investment prices has flattened considerably (Byrne and Corrado, 2017; Sichel, 2019; Goldin, Koutroumpis, Lafond and Winkler, 2024). Appendix D.3 shows that even if non-ICCT equipment is distinguished from non-ICCT structures, there is no significant difference in the trends of the prices and quantities between them.
The direction and magnitude of changes in labor shares vary by gender and skill (Figure 5). The income shares of male and female unskilled labor decreased, while the income shares of male and female skilled labor increased. The magnitude of the decline in the income share of male unskilled labor is substantially greater than that of the changes in the income shares of the other three types of labor. Taken together, the decline in the labor share is attributable to the fall in the income share of male unskilled labor. This finding is consistent with evidence of a decline in the income share of routine labor (Eden and Gaggl, 2018).
Figure 5: Income shares of the four types of labor
Notes: The income shares of male skilled labor, female skilled labor, male unskilled labor, and female unskilled labor are denoted by $w_{mh}\ell_{mh}/py$, $w_{fh}\ell_{fh}/py$, $w_{mu}\ell_{mu}/py$, and $w_{fu}\ell_{fu}/py$, respectively.
4.3 Factor substitution
Before we specify the sectoral production function, we conduct a preliminary analysis to acquire a general idea of substitutability between ICCT capital and the four types of labor. For labor of type $f$ in sector $n$, we estimate the following equation:
$\ln\left(\frac{k_{i,njt}}{\ell_{f,njt}}\right)=\varepsilon_{f,n}\ln\left(\frac{w_{f,njt}}{r_{i,njt}}\right)+\alpha_{nj}+\nu_{njt},$ (23)
where $\alpha_{nj}$ is a country-specific unobserved effect, $\nu_{njt}$ is the error term, and $j$ and $t$ are indices for countries and years. The parameter of interest is $\varepsilon_{f,n}$, which captures the percentage change in the capital–labor ratio associated with a one percent change in the relative price of labor to capital. The greater $\varepsilon_{f,n}$ is, the more substitutable labor of type $f$ is with ICCT capital in sector $n$.
We address two econometric issues: (i) unobserved heterogeneity and (ii) simultaneity. First, we control for country-specific unobserved effects by first differencing. In doing so, we take 5-year differences to deal with the measurement error problem. Then, we make use of two supply shifters as instrumental variables to identify equation (23) regarding the relative demand for factors. The first instrument is the lagged population size of each demographic group. The second instrument is the lagged price of ICCT investment relative to consumption, which measures investment-specific technology accelerating capital accumulation.
Table 1: Substitutability between ICCT capital and the four types of labor
OLS First-difference GMM1 First-difference GMM2
$\varepsilon_{fh}$ $\varepsilon_{mh}$ $\varepsilon_{fu}$ $\varepsilon_{mu}$ $\varepsilon_{fh}$ $\varepsilon_{mh}$ $\varepsilon_{fu}$ $\varepsilon_{mu}$ $\varepsilon_{fh}$ $\varepsilon_{mh}$ $\varepsilon_{fu}$ $\varsigma_{mu}$
1980–2005 1980–2005 1980–2005
Manufacturing 0.737 1.092 1.135 1.225 0.592 1.182 1.537 1.613 0.562 0.986 1.305 1.339
(0.136) (0.144) (0.179) (0.154) {0.378, 0.829} {1.041, 1.364} {1.298, 1.823} {1.392, 1.893} {0.373, 0.751} {0.832, 1.160} {1.050, 1.607} {1.113, 1.604}
Utilities 0.369 0.869 0.941 1.207 0.472 1.088 1.283 1.513 0.482 0.975 1.156 1.319
(0.234) (0.159) (0.058) (0.143) {0.321, 0.689} {0.841, 1.513} {1.044, 1.683} {1.235, 2.009} {0.347, 0.647} {0.696, 1.339} {0.886, 1.504} {1.029, 1.692}
Construction 0.728 1.016 1.015 1.004 1.107 1.828 1.819 2.027 0.923 1.373 1.456 1.583
(0.338) (0.445) (0.201) (0.254) {0.413, 2.047} {1.109, 2.865} {1.160, 2.753} {1.368, 3.024} {0.454, 1.558} {0.874, 2.053} {0.970, 2.123} {1.092, 2.277}
Services 0.652 0.929 0.993 1.137 0.713 1.109 1.350 1.554 0.687 0.963 1.189 1.306
(0.148) (0.106) (0.079) (0.077) {0.592, 0.858} {0.978, 1.302} {1.193, 1.573} {1.344, 1.886} {0.577, 0.798} {0.843, 1.100} {1.052, 1.345} {1.136, 1.508}
1980–1995 1980–1995 1980–1995
Manufacturing 0.969 1.092 0.962 1.062 0.599 1.564 1.967 2.232 0.592 1.398 1.836 2.039
(0.349) (0.370) (0.264) (0.294) {0.199, 1.009} {1.200, 2.109} {1.531, 2.545} {1.650, 3.247} {0.161, 0.996} {1.083, 1.853} {1.447, 2.366} {1.542, 2.903}
Utilities 0.277 0.581 0.751 0.987 0.490 1.397 1.586 2.047 0.545 1.540 1.743 2.222
(0.376) (0.302) (0.126) (0.217) {0.090, 1.000} {0.899, 3.007} {1.110, 2.709} {1.401, 4.071} {0.180, 0.929} {1.152, 2.289} {1.329, 2.498} {1.643, 3.481}
Construction 0.283 0.409 0.465 0.305 1.137 2.712 2.474 3.046 1.041 2.340 2.156 2.672
(0.640) (0.716) (0.304) (0.319) {0.023, 2.622} {1.292, 5.372} {1.251, 4.577} {1.737, 5.588} {–0.063, 2.410} {0.985, 4.474} {1.089, 3.820} {1.571, 4.577}
Services 0.492 0.588 0.784 0.783 0.662 1.450 1.668 2.170 0.655 1.441 1.668 2.152
(0.348) (0.253) (0.114) (0.146) {0.339, 0.950} {1.172, 2.079} {1.388, 2.254} {1.695, 3.397} {0.349, 0.928} {1.176, 2.010} {1.403, 2.158} {1.716, 3.061}
1990–2005 1990–2005 1990–2005
Manufacturing 0.765 1.048 0.880 0.965 0.576 0.986 1.298 1.296 0.548 0.891 1.180 1.183
(0.152) (0.164) (0.185) (0.160) {0.415, 0.763} {0.872, 1.144} {1.104, 1.579} {1.134, 1.531} {0.401, 0.700} {0.753, 1.050} {0.955, 1.461} {0.990, 1.415}
Utilities 0.281 0.735 0.656 0.968 0.440 0.950 1.118 1.260 0.457 0.877 1.037 1.157
(0.278) (0.113) (0.160) (0.161) {0.293, 0.630} {0.700, 1.287} {0.895, 1.435} {1.034, 1.587} {0.319, 0.629} {0.603, 1.242} {0.775, 1.386} {0.891, 1.503}
Construction 0.637 0.832 0.775 0.663 1.066 1.414 1.458 1.514 0.888 1.176 1.289 1.342
(0.328) (0.447) (0.143) (0.169) {0.440, 2.093} {0.798, 2.335} {0.885, 2.325} {0.973, 2.345} {0.488, 1.508} {0.749, 1.789} {0.848, 1.932} {0.920, 1.971}
Services 0.691 0.905 0.788 0.920 0.727 0.977 1.208 1.298 0.689 0.890 1.104 1.167
(0.189) (0.100) (0.094) (0.079) {0.580, 0.885} {0.849, 1.153} {1.057, 1.418} {1.137, 1.543} {0.564, 0.814} {0.772, 1.019} {0.978, 1.249} {1.035, 1.322}
Notes: Estimates of $\epsilon_{f,n}$ in equation (23) for female skilled labor ($f=fh$), male skilled labor ($f=mh$), female unskilled labor ($f=fu$), and male unskilled labor ($f=mu$) are reported. Standard errors are in parentheses of the first four columns. The 95% Anderson–Rubin confidence sets are in curly brackets of the last eight columns. Both of them are robust to clustering at the country level. Lagged population size and lagged ICCT investment price are used as instrumental variables in the fifth to eighth columns and the ninth to the last columns, respectively.
Table 1 presents the ordinary least squares (OLS) and first-difference generalized method of moments (GMM) estimates of $\varepsilon_{f,n}$ for female skilled labor ($f=fh$), male skilled labor ($f=mh$), female unskilled labor ($f=fu$), and male unskilled labor ($f=mu$) in manufacturing, utilities, construction, and services sectors. Both estimates indicate that the degrees of substitution between ICCT capital and the four types of labor are smaller in the order of female skilled, male skilled, female unskilled, and male unskilled labor for all sectors (i.e., $\varepsilon_{fh,n}<\varepsilon_{mh,n}<\varepsilon_{fu,n}<\varepsilon_{mu,n}$ for all $n$). The first-difference GMM estimates differ by the type of labor more than the OLS estimates. In both cases, the differences in the estimates by the type of labor are more significant for manufacturing, utilities, and services sectors than for the construction sector. Similar results are obtained for the 1980–1995 and 1990–2005 periods. Although the degrees of substitution between ICCT capital and unskilled labor become smaller from the former to the latter period, ICCT capital is more substitutable with male and female unskilled labor than with male and female skilled labor. The results described thus far are reinforced by the Anderson and Rubin (1949) confidence sets robust to weak instruments. The first-difference GMM estimates remain largely unchanged regardless of whether the lagged population size or the lagged investment price is used as an instrument. Overall, ICCT capital is more complementary not only to skilled labor than unskilled labor but also to female labor than male labor.
5 Estimation
This section describes the specification, identification, estimation, and aggregation of the sectoral production function. We consider the sectoral production function in which all four types of labor are imperfect substitutes for one another and substitutable to varying degrees with ICCT capital.
5.1 Specification
The economy consists of multiple sectors. In each sector, production requires two types of capital (ICCT capital, $k_{i}$, and non-ICCT capital, $k_{o}$) and four types of labor (male skilled labor, $\ell_{mh}$, female skilled labor, $\ell_{fh}$, male unskilled labor, $\ell_{mu}$, and female unskilled labor, $\ell_{fu}$). Building upon the analysis by Fallon and Layard (1975) and Krusell et al. (2000) and in Sections 4.2 and 4.3, the sectoral production function is specified as
$y_{n}=A_{n}k_{o,n}^{\theta_{o,n}}\left[\left(1-\theta_{mu,n}\right)B_{n}^{\sigma_{mu,n}}+\theta_{mu,n}\ell_{mu,n}^{\sigma_{mu,n}}\right]^{\frac{1-\theta_{o,n}}{\sigma_{mu,n}}},$ (24)
where
$B_{n}=$ $\left[\left(1-\theta_{fu,n}\right)C_{n}^{\sigma_{fu,n}}+\theta_{fu,n}\ell_{fu,n}^{\sigma_{fu,n}}\right]^{\frac{1}{\sigma_{fu,n}}},$ (25)
$C_{n}=$ $\left[\left(1-\theta_{mh,n}\right)D_{n}^{\sigma_{mh,n}}+\theta_{mh,n}\ell_{mh,n}^{\sigma_{mh,n}}\right]^{\frac{1}{\sigma_{mh,n}}},$ (26)
$D_{n}=$ $\left[\left(1-\theta_{fh,n}\right)k_{i,n}^{\sigma_{fh,n}}+\theta_{fh,n}\ell_{fh,n}^{\sigma_{fh,n}}\right]^{\frac{1}{\sigma_{fh,n}}}.$ (27)
This production function involves four substitution parameters ($\sigma_{mh}$, $\sigma_{fh}$, $\sigma_{mu}$, and $\sigma_{fu}$) that are less than one and five share parameters ($\theta_{mh}$, $\theta_{fh}$, $\theta_{mu}$, $\theta_{fu}$, and $\theta_{o}$) that lie between zero and one for each sector. The production function exhibits capital–skill complementarity for male labor if $\sigma_{mu}>\sigma_{mh}$ and for female labor if $\sigma_{fu}>\sigma_{fh}$, and capital–gender complementarity for skilled labor if $\sigma_{mh}>\sigma_{fh}$ and for unskilled labor if $\sigma_{mu}>\sigma_{fu}$. We allow for different degrees of capital–skill complementarity for male and female labor and different degrees of capital–gender complementarity for skilled and unskilled labor. In addition, we allow for different degrees of capital–skill–gender complementarity across sectors.
The multi-level nested CES production function (24) is not only analytically tractable enough to derive the general equilibrium effects of technological change on the relative wages and income shares of different types of labor but also flexible enough to allow for different degrees of capital–skill and capital–gender complementarities. There are three features in the specification of the production function. First, ICCT capital is located inside the CES function to allow for the possibility that technological change embodied in ICCT capital can influence the gender wage gap and the skill wage gap. There were dramatic changes in the price and quantity of ICCT equipment (see Figure 4b in the last section), whereas there were no such changes in the prices and quantities of non-ICCT equipment or structures (see Figure A1 in Appendix D.3). Second, ICCT capital is placed in the lowest nest in the four-level CES function to allow for the possibility that the four types of labor may be substitutable to varying degrees with ICCT capital. Otherwise, the rate of change in the gender (skill) wage gap due to progress in ICCT would be the same for skilled and unskilled (male and female) workers, which is not consistent with the data. Finally, the production function is required to be concave (i.e., the substitution parameters must be less than one) to satisfy the economic principle that the relative price of skills should decrease with a rise in the relative supply of skills. For this reason, skilled labor is placed in a nest lower than that of unskilled labor, as in Fallon and Layard (1975) and Krusell et al. (2000), and female labor is placed in a nest lower than that of male labor for each skill type. Our specification is consistent with key features of the data described in Sections 4.2 and 4.3. Appendices E.1 and E.2 provide further details on the assumptions, implications, and relevance of alternative specifications.
5.2 Identification
Cost minimization implies that the marginal rate of technical substitution equals the ratio of input prices irrespective of the degree of competition in goods markets. For the purpose of our analysis, we make use of five conditions, concerning (i) the relative price of female skilled labor to ICCT capital, (ii) the skilled gender wage gap, (iii) the female skill wage gap, (iv) the unskilled gender wage gap, and (v) the male skill wage gap, to estimate the production function parameters. The conditions lead to the following first-difference equations:
$\Delta\ln\left(\frac{w_{fh,n}}{r_{i,n}}\right)=$ $-\left(1-\sigma_{fh,n}\right)\Delta\ln\left(\frac{\ell_{fh,n}}{k_{i,n}}\right)+\Delta v_{fh,n},$ (28)
$\Delta\ln\left(\frac{w_{mh,n}}{w_{fh,,n}}\right)=$ $-\left(1-\sigma_{mh,n}\right)\Delta\ln\ell_{mh,n}+\left(1-\sigma_{fh,n}\right)\Delta\ln\ell_{fh,n}$
$-\left(\sigma_{mh,n}-\sigma_{fh,n}\right)\Delta\ln\widetilde{D}_{n}+\Delta v_{mh,n},$ (29)
$\Delta\ln\left(\frac{w_{fh,n}}{w_{fu,n}}\right)=$ $-\left(1-\sigma_{fh,n}\right)\Delta\ln\ell_{fh,n}+\left(1-\sigma_{fu,n}\right)\Delta\ln\ell_{fu,n}$
$+\left(\sigma_{mh,n}-\sigma_{fh,n}\right)\Delta\ln\widetilde{D}_{n}+\left(\sigma_{fu,n}-\sigma_{mh,n}\right)\Delta\ln\widetilde{C}_{n}+\Delta v_{fu,n},$ (30)
$\Delta\ln\left(\frac{w_{mu,n}}{w_{fu,n}}\right)=$ $-\left(1-\sigma_{mu,n}\right)\Delta\ln\ell_{mu,n}+\left(1-\sigma_{fu,n}\right)\Delta\ln\ell_{fu,n}$
$-\left(\sigma_{mu,n}-\sigma_{fu,n}\right)\Delta\ln\widetilde{B}_{n}+\Delta v_{mu1,n},$ (31)
$\Delta\ln\left(\frac{w_{mh,n}}{w_{mu,n}}\right)=$ $\left(1-\sigma_{mu,n}\right)\Delta\ln\ell_{mu,n}-\left(1-\sigma_{mh,n}\right)\Delta\ln\ell_{mh,n}$
$+\left(\sigma_{fu,n}-\sigma_{mh,n}\right)\Delta\ln\widetilde{C}_{n}+\left(\sigma_{mu,n}-\sigma_{fu,n}\right)\Delta\ln\widetilde{B}_{n}+\Delta v_{mu2,n}.$ (32)
where the error terms, $\Delta v$, are added to account for measurement errors and shocks to non-neutral disembodied technologies, as detailed in Appendix E.3, and the log changes of the CES aggregates are approximated by the weighted average of the log changes of factor inputs. It can be shown by Taylor expansion that
$\Delta\ln D_{n}\simeq$ $\left(1-\varphi_{fh,n}\right)\Delta\ln k_{i,n}+\varphi_{fh,n}\Delta\ln\ell_{fh,n}\equiv\Delta\ln\widetilde{D}_{n},$
$\Delta\ln C_{n}\simeq$ $\left(1-\varphi_{mh,n}\right)\Delta\ln\widetilde{D}_{n}+\varphi_{mh,n}\Delta\ln\ell_{mh,n}\equiv\Delta\ln\widetilde{C}_{n},$
$\Delta\ln B_{n}\simeq$ $\left(1-\varphi_{fu,n}\right)\Delta\ln\widetilde{C}_{n}+\varphi_{fu,n}\Delta\ln\ell_{fu,n}\equiv\Delta\ln\widetilde{B}_{n},$
where the weights can be calculated directly from the data as $\varphi_{fh,n}=w_{fh,n}\ell_{fh,n}/(r_{i,n}k_{i,n}+w_{fh,n}\ell_{fh,n})$, $\varphi_{mh,n}=w_{mh,n}\ell_{mh,n}/(r_{i,n}k_{i,n}+w_{fh,n}\ell_{fh,n}+w_{mh,n}\ell_{mh,n})$, and $\varphi_{fu,n}=w_{fu,n}\ell_{fu,n}/(r_{i,n}k_{i,n}+w_{fh,n}\ell_{fh,n}+w_{mh,n}\ell_{mh,n}+w_{fu,n}\ell_{fu,n})$. [^12] As a result of these approximations, the share parameters are eliminated from the system of equations (28)–(32). The substitution parameters can be estimated solely using equations (28)–(32). However, it is unnecessary to estimate the share parameters in our analysis because they are not required to compute the aggregate elasticities of substitution. Alternatively, the share and substitution parameters can be estimated sequentially using the levels and log differences of the marginal-rate-of-technical substitution equations. We confirm that the approximations do not affect the estimates of the substitution parameters by comparing the results obtained from both approaches. Because all time-invariant factors are eliminated by first differencing, the estimates of the system of equations (28)–(32) are immune to the omitted variable bias due to unobserved country-specific characteristics. Differences are taken over five years to avoid the attenuation bias due to measurement error.
Because prices and quantities are jointly determined in markets, identifying equations (28)–(32) for factor demand requires using supply shifters as instrumental variables. The instruments used are the lagged population size of each demographic group and the lagged price of ICCT investment relative to consumption. This means that we exploit the variation in past birth rates and university enrollment capacity or the variation in investment-specific technology across countries over time by taking advantage of cross-country panel data. With either of the instruments, it is straightforward to see that the substitution parameters $\sigma_{fh}$, $\sigma_{mh}$, $\sigma_{fu}$, and $\sigma_{mu}$ can be identified sequentially from equations (28)–(32) in the order of those in equations (27), (26), (25), and (24) from the lowest to the highest nest. The error terms are presumably correlated across equations (28)–(32). The substitution parameters in the system of equations (28)–(32) are estimated jointly by GMM to improve efficiency. Standard errors are clustered at the country level to account for heteroscedasticity and serial correlation. In addition, we control for other factors such as labor market institutions, disembodied factor-biased technological change, and markups.
5.3 Aggregation
When the sectoral production function is of the nested CES form (24), the aggregate elasticity of substitution (13) can be expressed as the weighted average of the substitution parameters in production and consumption:
$\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}=$ $\sum_{n=1}^{N}\pi_{\ell_{f}\ell_{g},n}^{fh}\left(\frac{1}{1-\sigma_{fh,n}}\right)+\sum_{n=1}^{N}\pi_{\ell_{f}\ell_{g},n}^{mh}\left(\frac{1}{1-\sigma_{mh,n}}\right)+\sum_{n=1}^{N}\pi_{\ell_{f}\ell_{g},n}^{fu}\left(\frac{1}{1-\sigma_{fu,n}}\right)+\sum_{n=1}^{N}\pi_{\ell_{f}\ell_{g},n}^{mu}\left(\frac{1}{1-\sigma_{mu,n}}\right)$
$+\sum_{n=1}^{N}\pi_{\ell_{f}\ell_{g},n}^{o}+\sum_{n=1}^{N}\pi_{\ell_{f}\ell_{g},n}^{c}\left(\frac{1}{1-\eta}\right),$ (33)
where the sum of the weights on the substitution parameters adds up to one. The weights on the substitution parameters depend on the income shares of each factor, as detailed in Appendix E.4. The elasticity of substitution in consumption is set at $1/(1-\eta)=0.9$, as in Baqaee and Farhi (2019).
After computing the aggregate elasticities of substitution using the estimated substitution parameters, we can evaluate the effects of technological change on, as well as the contributions of factor inputs to, the relative wages and income shares of the four types of labor. The direction and magnitude of the effects of technological change embodied in ICCT capital depend on the presence and degree of capital–skill or capital–gender complementarity (i.e., the sign and magnitude of differences between the aggregate elasticities, $\epsilon_{\ell_{f}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{g}k_{i}}^{\mathcal{C}}$) and the rate of progress in ICCT. The degrees of capital–skill and capital–gender complementarities are approximately proportional to the differences in the substitution parameters and the income share of ICCT capital relative to the income shares of each type of labor. Appendix E.5 provides the exact expressions for the effects of technological change on the relative wages and income shares of the four types of labor.
6 Results
This section begins by presenting the estimates of sectoral production function parameters and assessing the robustness of the estimates. It then reports the estimates of the aggregate elasticities of substitution and the quantitative contributions of factor inputs to the relative wages and income shares of the four types of labor. The section ends with findings on the effects of technological change.
6.1 Sectoral results
6.1.1 Production function estimates
Table 2 presents the GMM estimates of the substitution parameters in the production functions for manufacturing, utilities, construction and services sectors. For comparison it also presents corresponding estimates under the assumption that the aggregate production function takes the form of equation (24). Significant differences are observed in the estimates of the four substitution parameters for three out of four sectors. For manufacturing, utilities, and services sectors, the null hypotheses of $\sigma_{fh}=\sigma_{mh}$, $\sigma_{mh}=\sigma_{fu}$, and $\sigma_{fu}=\sigma_{mu}$ are rejected at the 2 percent significance level (see Panel D of Table A1 in Appendix E.1). The estimates suggest that ICCT capital is more complementary not only to skilled labor than unskilled labor but also to female labor than male labor (i.e., $\sigma_{fh}<\sigma_{mh}<\sigma_{fu}<\sigma_{mu}$). Moreover, the null hypothesis that the substitution parameters are equal across the three sectors is rejected at the 2 percent significance level. [^13] The null hypotheses that the substitution parameters are equal between the manufacturing and utilities sectors, between the utilities and services sectors, and between the manufacturing and services sectors are rejected at the 10 percent, 3 percent, and 13 percent significance levels, respectively. For the construction sector, the substitution parameters are estimated under the null hypothesis of equal parameters (i.e., $\sigma_{fh}=\sigma_{mh}=\sigma_{fu}=\sigma_{mu}$) because the null hypothesis cannot be rejected at the 39 percent significance level. [^14] On average in our sample, the shares of college graduates and women each account for less than 10 percent in the construction sector. This pattern is consistent with the absence of capital–skill and capital–gender complementarities. These results are not an artifact of an arbitrary nesting structure of the CES production function. The relative magnitude of the estimated degrees of factor substitution is consistent with that reported in Table 1. The estimates remain similar regardless of whether the lagged population size or the lagged investment price is used as an instrument. The estimates obtained with the lagged population size are used in the subsequent analyses because they are more precise. When the substitution parameters are estimated under the assumption that the aggregate production function for the non-primary sector exists, the parameter estimates differ from those obtained for the manufacturing, utilities, and construction sectors.
Table 2: Production function estimates for each sector
GMM1 GMM2
$\sigma_{fh}$ $\sigma_{mh}$ $\sigma_{fu}$ $\sigma_{mu}$ $\sigma_{fh}$ $\sigma_{mh}$ $\sigma_{fu}$ $\sigma_{mu}$
Manufacturing –0. 689 0. 288 0. 600 0. 776 –0. 780 0. 136 0. 540 0. 693
(0. 319) (0. 046) (0. 043) (0. 054) (0. 303) (0. 074) (0. 069) (0. 052)
Utilities –0. 962 0. 184 0. 404 0. 631 –1. 076 0. 064 0. 329 0. 551
(0. 362) (0. 104) (0. 063) (0. 058) (0. 324) (0. 158) (0. 084) (0. 051)
Construction 0. 628 0. 628 0. 628 0. 628 0. 506 0. 506 0. 506 0. 506
(0. 077) (0. 077) (0. 077) (0. 077) (0. 092) (0. 092) (0. 092) (0. 092)
Services –0. 413 0. 340 0. 598 0. 861 –0. 456 0. 203 0. 561 0. 789
(0. 134) (0. 067) (0. 034) (0. 048) (0. 119) (0. 075) (0. 036) (0. 058)
Non-primary –0. 477 0. 333 0. 604 0. 847 –0. 517 0. 193 0. 560 0. 784
(0. 171) (0. 061) (0. 036) (0. 047) (0. 145) (0. 071) (0. 041) (0. 044)
Notes: Standard errors in parentheses are clustered at the country level. Lagged population size and lagged ICCT investment price are used as instrumental variables in the first to fourth columns and the fifth to the last columns, respectively. The production function parameters for the construction sector are estimated under the restriction of $\sigma_{fh}=\sigma_{mh}=\sigma_{fu}=\sigma_{mu}$ because the restriction cannot be rejected.
6.1.2 Robustness checks
Table 3 presents the robustness of the estimates to controlling for the influence of labor market institutions, goods market imperfections, and disembodied factor-biased technological change. The first three rows report the estimates of the substitution parameters after controlling for the influence of labor market institutions. We allow for the possibility that the actual wage may deviate from the competitive wage depending on the collective bargaining coverage, the strictness of employment protection legislation, and the presence and level of minimum wages. Specifically, we add the collective bargaining coverage and the employment protection legislation in log-difference form and the presence of minimum wages and its interaction with their level in first-difference form to equations (28)–(32). The relative magnitude of the estimated substitution parameters remains unchanged with the additional controls. Although the first three rows of Table 3 report the estimates when the three sets of variables are added individually to avoid the multicollinearity problem, the relative magnitude of the estimated substitution parameters remains unchanged even if they are added jointly.
The fourth row of Table 3 reports the estimates of the substitution parameters after taking into account imperfect competition in goods markets. We can relax the assumption of competitive markets by recalculating the rental price of capital based on the external rate of return because all estimating equations hold irrespective of the degree of markup. The estimates of the substitution parameters remain essentially unchanged.
The last row of Table 3 reports the estimates of the substitution parameters after controlling for disembodied factor-biased technological change. We take this into account by adding trend polynomials with country-specific coefficients in equations (28)–(32). Trend polynomials capture changes in the direction and magnitude of disembodied technological change, while country-specific coefficients capture cross-country differences in the speed and timing of disembodied technological change. We choose the order of the polynomials to fit the trends in relative wages for each sector and country. Basically, we include trend polynomials when they are statistically significant at the 10 percent level. The estimates of the substitution parameters remain essentially unchanged. The results are robust to country-specific nonlinear time trends regardless of their causes even though the trends could be attributable in part to other unobserved factors such as discrimination and social norms.
Table 3: Production function estimates with additional controls
Manufacturing sector Construction sector
$\sigma_{fh}$ $\sigma_{mh}$ $\sigma_{fu}$ $\sigma_{mu}$ $\sigma_{fh}$ $\sigma_{mh}$ $\sigma_{fu}$ $\sigma_{mu}$
Collective bargaining –0. 601 0. 307 0. 626 0. 720 0. 676 0. 676 0. 676 0. 676
   coverage (0. 408) (0. 074) (0. 047) (0. 065) (0. 074) (0. 074) (0. 074) (0. 074)
Employment protection –0. 855 0. 166 0. 553 0. 707 0. 496 0. 496 0. 496 0. 496
   legislation (0. 328) (0. 056) (0. 055) (0. 065) (0. 082) (0. 082) (0. 082) (0. 082)
Minimum wages –0. 668 0. 387 0. 635 0. 784 0. 592 0. 592 0. 592 0. 592
(0. 439) (0. 057) (0. 053) (0. 094) (0. 071) (0. 071) (0. 071) (0. 071)
Goods market –0. 928 0. 191 0. 559 0. 733 0. 518 0. 518 0. 518 0. 518
   imperfections (0. 368) (0. 061) (0. 049) (0. 058) (0. 087) (0. 087) (0. 087) (0. 087)
Disembodied factor-biased –0. 722 0. 336 0. 542 0. 854 0. 674 0. 674 0. 674 0. 674
   technological change (0. 141) (0. 047) (0. 028) (0. 030) (0. 060) (0. 060) (0. 060) (0. 060)
Utilities sector Services sector
$\sigma_{fh}$ $\sigma_{mh}$ $\sigma_{fu}$ $\sigma_{mu}$ $\sigma_{fh}$ $\sigma_{mh}$ $\sigma_{fu}$ $\sigma_{mu}$
Collective bargaining –0. 891 0. 204 0. 375 0. 566 –0. 458 0. 282 0. 584 0. 809
   coverage (0. 421) (0. 139) (0. 083) (0. 054) (0. 186) (0. 074) (0. 040) (0. 064)
Employment protection –1. 344 0. 152 0. 348 0. 594 –0. 563 0. 246 0. 567 0. 828
   legislation (0. 400) (0. 113) (0. 054) (0. 042) (0. 159) (0. 065) (0. 041) (0. 060)
Minimum wages –1. 117 0. 211 0. 388 0. 636 –0. 418 0. 361 0. 604 0. 872
(0. 365) (0. 109) (0. 066) (0. 063) (0. 139) (0. 081) (0. 040) (0. 059)
Goods market –1. 200 0. 076 0. 330 0. 565 –0. 626 0. 212 0. 518 0. 791
   imperfections (0. 363) (0. 088) (0. 064) (0. 054) (0. 169) (0. 073) (0. 034) (0. 047)
Disembodied factor-biased –0. 661 0. 439 0. 705 0. 908 –0. 371 0. 473 0. 627 0. 899
   technological change (0. 246) (0. 132) (0. 081) (0. 050) (0. 095) (0. 086) (0. 049) (0. 043)
Notes: Standard errors in parentheses are clustered at the country level. Lagged population size is used as an instrumental variable. The production function parameters for the construction sector are estimated under the restriction of $\sigma_{fh}=\sigma_{mh}=\sigma_{fu}=\sigma_{mu}$ because it cannot be rejected.
6.2 Aggregate results
6.2.1 Elasticities of substitution
Table 4 presents the estimates of the elasticities of substitution in the aggregate production and cost functions. The estimated elasticities of substitution between ICCT capital and the four types of labor are greater in the order of male unskilled, female unskilled, male skilled, and female skilled labor in both cases. The elasticity estimates imply that the expansion of ICCT equipment would narrow the gender wage gap and widen the skill wage gap. In this sense, the estimates are consistent with the literature that points to technological advances as causes of changes in the structure of wages and employment (Autor, Levy and Murnane, 2003; Black and Spitz-Oener, 2010). The magnitude of the factor complementarity effects associated with a rise in ICCT capital is directly proportional to the differences in the elasticities of substitution between ICCT capital and the four types of labor as well as the rate of increase in ICCT capital. The elasticity estimates imply that the effects of ICCT capital on the gender (skill) wage gap would be greater for skilled (female) labor than for unskilled (male) labor. All else held constant, a one percent increase in ICCT capital would reduce the skilled (unskilled) gender wage gap by 0.30 (0.02) percent and raise the male (female) skill wage gap by 0.08 (0.35) percent. At the same time, the magnitude of the relative quantity effects is inversely proportional to the elasticities of labor–labor substitution and directly proportional to the rates of increases in the quantities of labor. The elasticity estimates imply that the effects of labor quantities would be greater in the order of female skilled, male skilled, female unskilled, and male unskilled labor.
Table 4: Aggregate elasticities of substitution
$\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{F}}$ $\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}$
$k_{i}$ $\ell_{fh}$ $\ell_{mh}$ $\ell_{fu}$ $\ell_{mu}$ $k_{o}$ $k_{i}$ $\ell_{fh}$ $\ell_{mh}$ $\ell_{fu}$ $\ell_{mu}$ $k_{o}$
$k_{i}$ 0. 78 1. 52 2. 49 5. 66 1. 00 $k_{i}$ 0. 79 1. 53 2. 50 5. 86 1. 00
(0. 04) (0. 08) (0. 14) (0. 82) (0. 00) (0. 04) (0. 09) (0. 14) (0. 99) (0. 00)
$\ell_{fh}$ 0. 76 1. 54 2. 48 6. 22 1. 00 $\ell_{fh}$ 0. 75 1. 52 2. 47 5. 93 1. 00
(0. 04) (0. 09) (0. 14) (1. 44) (0. 00) (0. 04) (0. 08) (0. 13) (1. 04) (0. 00)
$\ell_{mh}$ 0. 99 1. 06 2. 50 5. 80 1. 00 $\ell_{mh}$ 1. 12 1. 20 2. 51 5. 86 1. 00
(0. 04) (0. 05) (0. 14) (0. 93) (0. 00) (0. 05) (0. 07) (0. 14) (0. 99) (0. 00)
$\ell_{fu}$ 1. 05 1. 12 1. 96 5. 90 1. 00 $\ell_{fu}$ 1. 31 1. 41 2. 07 5. 89 1. 00
(0. 05) (0. 05) (0. 10) (1. 01) (0. 00) (0. 06) (0. 07) (0. 10) (1. 01) (0. 00)
$\ell_{mu}$ 1. 07 1. 16 2. 26 3. 44 1. 00 $\ell_{mu}$ 1. 67 1. 83 3. 07 4. 11 1. 00
(0. 05) (0. 06) (0. 13) (0. 27) (0. 00) (0. 14) (0. 18) (0. 34) (0. 53) (0. 00)
$k_{o}$ 1. 03 1. 10 1. 76 2. 03 1. 73 $k_{o}$ 1. 40 1. 53 2. 33 2. 93 3. 49
(0. 04) (0. 05) (0. 07) (0. 07) (0. 04) (0. 09) (0. 11) (0. 19) (0. 29) (0. 51)
Notes: The elasticities of substitution are computed using equations (12) and (13) of Lemma 1 and evaluated at the sample means across all countries and years. Standard errors in parentheses are computed using bootstrap with 500 replications.
The estimated elasticities of labor–labor substitution in the aggregate production (cost) function range from 1.1 to 6.2 (1.2 to 5.9). The elasticity estimates are within the range of standard estimates in the literature, although they are not strictly comparable to those in other studies, such as Acemoglu, Autor and Lyle (2004) and Johnson and Keane (2013), because of differences in the degree of classification of labor. The estimated aggregate elasticities of substitution between female skilled labor and the other three types of labor are greater in the order of male unskilled, female unskilled, and male skilled labor. Both female skilled and unskilled labor are more complementary to male skilled labor than to male unskilled labor.
The Morishima (1967) elasticities of substitution are asymmetric by definition. The differences in the estimated elasticities of relative demand for capital with respect to wages are greater than those in the estimated elasticities of relative demand for labor with respect to the rental prices of capital. However, in both cases, ICCT capital is more complementary not only to skilled labor than unskilled labor but also to female labor than male labor.
The estimated elasticity of substitution between ICCT capital and male unskilled labor in the aggregate cost (production) function exceeds one at $\epsilon_{\ell_{mu}k_{i}}^{\mathcal{C}}=1.67$ ($\epsilon_{\ell_{mu}k_{i}}^{\mathcal{F}}=1.07$). ICCT capital is consistently more substitutable with male unskilled labor than with the other three types of labor. The estimates suggest that the income share of male unskilled labor is likely to decline with a fall in the price or a rise in the quantity of ICCT capital.
6.2.2 Contributions of factor inputs
Gender and skill premia
Table 5 presents the quantitative contributions of specific factor inputs to changes in the gender premium by skill and the skill premium by gender. The first column reports the observed changes in gender and skill premia, while the second column reports the sum of the changes attributable to factor inputs. These results are obtained without adjusting any parameters to match relative wages in the aggregate economy. The rate of decline in the gender wage gap is greater among unskilled workers than among skilled workers, while the rate of increase in the skill wage gap is greater among male workers than among female workers. From this point of view, the observed pattern of changes in gender and skill premia is consistent with the total contribution of factor inputs.
Table 5: Contributions of factor inputs to gender and skill premia, 1980–2005
Data Total FCE RQE $\Delta\ln k_{i}$ $\Delta\ln\ell_{fh}$ $\Delta\ln\ell_{mh}$ $\Delta\ln\ell_{fu}$ $\Delta\ln\ell_{mu}$ $\Delta\ln k_{o}$
$\Delta\ln\left(\frac{w_{mh}}{w_{fh}}\right)$ 0. 012 0. 070 –0. 885 0. 954 –0. 884 1. 423 –0. 469 –0. 001 0. 000 0. 000
(0. 050) (0. 070) (0. 047) (0. 070) (0. 067) (0. 027) (0. 001) (0. 000) (0. 000)
$\Delta\ln\left(\frac{w_{mu}}{w_{fu}}\right)$ –0. 057 –0. 103 –0. 164 0. 061 –0. 070 –0. 045 –0. 049 0. 059 0. 001 0. 000
(0. 015) (0. 013) (0. 005) (0. 005) (0. 004) (0. 004) (0. 004) (0. 000) (0. 000)
$\Delta\ln\left(\frac{w_{mh}}{w_{mu}}\right)$ 0. 118 0. 052 0. 373 –0. 321 0. 222 0. 129 –0. 320 0. 022 –0. 001 0. 000
(0. 019) (0. 023) (0. 019) (0. 014) (0. 010) (0. 019) (0. 002) (0. 000) (0. 000)
$\Delta\ln\left(\frac{w_{fh}}{w_{fu}}\right)$ 0. 049 –0. 121 1. 136 –1. 257 1. 036 –1. 340 0. 100 0. 082 0. 000 0. 000
(0. 046) (0. 071) (0. 059) (0. 069) (0. 062) (0. 013) (0. 005) (0. 000) (0. 000)
$\Delta\ln\left(\frac{\overline{w}_{M}}{\overline{w}_{F}}\right)$ –0. 083 –0. 111 –0. 307 0. 196 –0. 228 0. 147 –0. 094 0. 063 0. 001 0. 000
(0. 020) (0. 022) (0. 012) (0. 017) (0. 014) (0. 006) (0. 004) (0. 000) (0. 000)
$\Delta\ln\left(\frac{\overline{w}_{H}}{\overline{w}_{U}}\right)$ 0. 081 –0. 017 0. 574 –0. 590 0. 454 –0. 356 –0. 172 0. 058 0. 000 0. 000
(0. 022) (0. 030) (0. 028) (0. 024) (0. 016) (0. 011) (0. 002) (0. 000) (0. 000)
Notes: The first column reports the observed changes in the logarithms of relative wages. The second column reports the sum of the factor complementarity effect (FCE) in the third column and the relative quantity effect (RQE) in the fourth column, or equivalently, the sum of the individual contributions of factor inputs in the fifth to last columns. The figures in the third to last columns are computed using equation (18) of Proposition 2. Standard errors in parentheses are computed using bootstrap with 500 replications.
The third and fourth columns report the changes attributable to the factor complementarity effect and the relative quantity effect. The fifth to last columns report the changes attributable to each factor input. This period shows a substantial increase in ICCT capital and female skilled labor, a moderate increase in non-ICCT capital and male skilled labor, a small increase in female unskilled labor, and a small decrease in male unskilled labor. The increase in ICCT capital contributes to narrowing the gender wage gap and widening the skill wage gap. Meanwhile, the increase in male and female skilled labor contributes to widening the gender wage gap and narrowing the skill wage gap. Consequently, whether the gender (skill) wage gap narrows (widens) depends on the outcome of the race between progress in ICCT and advances in female employment (educational attainment). These results hold not only for the four wage gaps (skilled gender wage gap, unskilled gender wage gap, male skill wage gap, and female skill wage gap) reported in the first four rows, but also for the two aggregate wage gaps (aggregate gender wage gap and aggregate skill wage gap) reported in the last two rows.
The factor complementarity effect associated with a rise in ICCT capital is greater for the skilled gender wage gap and the female skill wage gap, reflecting the result that ICCT capital is significantly more complementary to female skilled labor than to the other three types of labor. If all else were held constant, the expansion of ICCT equipment during the period would have reduced the skilled gender wage gap by 88 percent and raised the female skill wage gap by 104 percent. At the same time, the relative quantity effect is greater for the skilled gender wage gap and the female skill wage gap, reflecting the fact that the rate of increase in female skilled labor is significantly greater than that in the other three types of labor. If all else were held constant, a fall in the relative quantity of male skilled labor to female skilled labor during the period would have raised the skilled gender wage gap by 95 ($(1.423-0.469)\times 100$) percent, and a rise in the relative quantity of female skilled labor to female unskilled labor would have reduced the female skill wage gap by 126 ($(\text{\textendash}1.340+0.082)\times 100$) percent. Consequently, the skilled gender wage gap did not narrow as the two large effects canceled out. Similarly, the female skill wage gap did not widen as the two large effects canceled out. Meanwhile, the unskilled gender wage gap narrowed as the factor complementarity effect exceeded the relative quantity effect. Similarly, the male skill wage gap widened as the factor complementarity effect exceeded the relative quantity effect. A rise in female skilled labor also increased the male skill wage gap as female skilled labor was more complementary to male skilled labor than to male unskilled labor. Given the results that the factor complementarity effects on the skilled gender wage gap and the female skill wage gap are so large, if there were no change in the relative supply of female skilled labor, the skilled gender wage gap would have narrowed more than the unskilled gender wage gap, and the female skill wage gap would have widened more than the male skill wage gap. Taken together, a smaller rate of decline in the gender wage gap for skilled workers than for unskilled workers and a smaller rate of increase in the skill wage gap for female workers than for male workers are attributable not to a lack of demand for female skilled labor but to an abundance in the supply of female skilled labor.
Labor shares
Changes in the labor share can be decomposed into those in the income shares of the four types of labor $\ell_{f}$ for $f\in L=\left\{mh,fh,mu,fu\right\}$, as shown in equation (17).
$\Delta\ln\left(\frac{\sum_{f\in L}w_{f}\ell_{f}}{py}\right)=\sum_{f\in L}\left(\frac{w_{f}\ell_{f}}{\sum_{f\in L}w_{f}\ell_{f}}\right)\Delta\ln\left(\frac{w_{f}\ell_{f}}{py}\right),$ (34)
where $py=\sum_{f\in L}w_{f}\ell_{f}+r_{i}k_{i}+r_{o}k_{o}$. We evaluate the weights on the rates of changes in the income shares of each type of labor at the sample means across all countries and years. Changes in the labor share between the years 1980 and 2005 can be decomposed as
$\Delta\ln\left(\frac{\sum_{f\in L}w_{f}\ell_{f}}{py}\right)$ $=$ $\underset{\textrm{male skilled}}{\underbrace{0.165\times 0.457}}+\underset{\textrm{female skilled}}{\underbrace{0.066\times 1.230}}+\underset{\textrm{male unskilled}}{\underbrace{0.513\times\left(-0.389\right)}}+\underset{\textrm{female unskilled}}{\underbrace{0.256\times\left(-0.122\right)}}$
$=$ $0.075+0.081-0.200-0.031.$
The labor share did not decline uniformly across gender and skill groups. The income shares of male and female skilled labor increased, while the income shares of male and female unskilled labor declined. The reason for the decline in the labor share is that the majority of labor was unskilled to whom the share of income decreased. The decline in the labor share is attributable almost entirely to a fall in the income share of male unskilled labor.
Table 6: Contributions of factor inputs to labor shares, 1980–2005
Data Total FCE RQE $\Delta\ln k_{i}$ $\Delta\ln\ell_{fh}$ $\Delta\ln\ell_{mh}$ $\Delta\ln\ell_{fu}$ $\Delta\ln\ell_{mu}$ $\Delta\ln k_{o}$
$\Delta\ln\left(\frac{w_{mh}\ell_{mh}}{py}\right)$ 0. 457 0. 438 0. 147 0. 291 0. 092 0. 050 0. 312 –0. 019 0. 003 0. 000
(0. 015) (0. 020) (0. 011) (0. 015) (0. 010) (0. 016) (0. 001) (0. 000) (0. 000)
$\Delta\ln\left(\frac{w_{fh}\ell_{fh}}{py}\right)$ 1. 230 1. 153 1. 030 0. 123 0. 976 0. 134 0. 059 –0. 018 0. 003 0. 000
(0. 047) (0. 067) (0. 051) (0. 071) (0. 060) (0. 012) (0. 002) (0. 000) (0. 000)
$\Delta\ln\left(\frac{w_{mu}\ell_{mu}}{py}\right)$ –0. 389 –0. 343 –0. 259 –0. 084 –0. 130 –0. 079 –0. 091 –0. 041 –0. 003 0. 000
(0. 011) (0. 008) (0. 009) (0. 004) (0. 003) (0. 003) (0. 001) (0. 000) (0. 000)
$\Delta\ln\left(\frac{w_{fu}\ell_{fu}}{py}\right)$ –0. 122 –0. 029 –0. 086 0. 058 –0. 060 –0. 034 –0. 042 0. 104 0. 003 0. 000
(0. 009) (0. 009) (0. 007) (0. 005) (0. 003) (0. 003) (0. 003) (0. 000) (0. 000)
$\Delta\ln\left(\frac{\sum_{f}w_{f}\ell_{f}}{py}\right)$ –0. 075 –0. 035 –0. 063 0. 028 –0. 003 –0. 032 –0. 002 0. 001 0. 000 0. 000
(0. 005) (0. 003) (0. 002) (0. 007) (0. 002) (0. 001) (0. 000) (0. 000) (0. 000)
Notes: The first column reports the observed changes in the logarithms of labor shares. The second column reports the sum of the factor complementarity effect (FCE) in the third column and the relative quantity effect (RQE) in the fourth column, or equivalently, the sum of the individual contributions of factor inputs in the fifth to last columns. The figures in the third to last columns are computed using equation (15) of Proposition 1. Standard errors in parentheses are computed using bootstrap with 500 replications.
Table 6 presents the quantitative contributions of specific factor inputs to changes in the income shares of the four types of labor. The first column reports the observed changes in the income shares of the four types of labor, while the second column reports the sum of the changes attributable to factor inputs. These results are obtained without adjusting any parameters to match labor shares in the aggregate economy. The observed changes in the income shares of the four types of labor align with the total contribution of factor inputs. The total contribution to changes in the income shares of male and female skilled labor are positive and statistically significant, while that in the income shares of male and female unskilled labor are negative and statistically significant.
The third and fourth columns report the changes attributable to the factor complementarity effect and the relative quantity effect. The two effects work in the same direction for the income shares of male and female skilled labor and male unskilled labor. The former is much greater than the latter for the income shares of female skilled labor and male unskilled labor, although the converse is true for the income share of male skilled labor. The two effects work in opposite directions for the income share of female unskilled labor. However, the direction of the factor complementarity effect aligns with that of the actual change for the income shares of all types of labor. Overall, the factor complementarity effect is essential for understanding changes in the income shares of different types of labor. Both the factor complementarity effect and the relative quantity effect are obscured when aggregated across different types of labor, because the two effects tend to be positive for skilled labor but negative for unskilled labor.
The fifth to last columns report the changes attributable to each factor input. The expansion of ICCT equipment contributes significantly to the increase in the income shares of male and female skilled labor and the decrease in the income shares of male and female unskilled labor. The magnitude of the effect depends on the extent to which the rate of increase in the wages of male or female skilled (unskilled) labor is greater (smaller) than that in the nominal output. A rise in male and female skilled labor also contributes modestly to changes in labor shares.
6.2.3 Effects of technological change
The first three columns of Table 7 present the effects of three types of technological change on the relative wages and income shares of the four types of labor. The effects of technological change in Table 7 are measured in levels to facilitate interpretation of the magnitude of the effects on each relative wage and income share, whereas the contributions of factor inputs in Tables 5 and 6 are measured in logarithms to facilitate comparison of the contributions to the four wage gaps or four income shares. The first to third columns report the changes attributable to technological change embodied in ICCT capital, technological change embodied in non-ICCT capital, and sector-specific disembodied technological change, respectively. The effects of technological change embodied in ICCT capital are significant, whereas those of technological change embodied in non-ICCT capital and of sector-specific disembodied technological change are marginal. These results hold not only for the four wage gaps and the four labor shares but also for the aggregate gender wage gap and the aggregate skill wage gap. The direction of the effects of technological change embodied in ICCT capital aligns with that of the effects of the expansion of ICCT equipment in Tables 5 and 6, reflecting the presence of capital–skill–gender complementarity. The results indicate that a fall in the relative price of ICCT investment owing to technological progress contributes to a narrowing of the gender wage gap, a widening of the skill wage gap, an increase in the income share of skilled labor, and a decline in the income share of unskilled labor. The magnitude of the effect on relative wages is largest for female skilled labor in percentage-point terms, while that of the effect on labor shares is largest for male unskilled labor. The small effects of technological change embodied in non-ICCT capital reflect the lack of significant decrease in the relative price of non-ICCT investment and of complementarity among non-ICCT capital, skills, and gender. The small effects of sector-specific disembodied technological change reflect the lack of significant change in sector-specific disembodied technology.
Table 7: Effects of technological change, 1980–2005
$\Delta\ln q_{i}$ $\Delta\ln q_{o}$ $\Delta\ln A_{n}$ $\Delta\ln q_{i}$
    $1/\gamma$ 0 0 0 1/2 1 2
Relative wages Relative wages
$\Delta\left(\frac{w_{mh}}{w_{fh}}\right)$ –0. 886 –0. 015 0. 048 –0. 628 –0. 486 –0. 333
(0. 052) (0. 001) (0. 003) (0. 042) (0. 036) (0. 027)
$\Delta\left(\frac{w_{mu}}{w_{fu}}\right)$ –0. 074 –0. 001 0. 003 –0. 090 –0. 096 –0. 098
(0. 006) (0. 000) (0. 001) (0. 008) (0. 010) (0. 012)
$\Delta\left(\frac{w_{mh}}{w_{mu}}\right)$ 0. 262 0. 004 –0. 015 0. 272 0. 264 0. 239
(0. 014) (0. 000) (0. 001) (0. 013) (0. 013) (0. 015)
$\Delta\left(\frac{w_{fh}}{w_{fu}}\right)$ 1. 283 0. 021 –0. 070 0. 958 0. 767 0. 550
(0. 053) (0. 001) (0. 003) (0. 044) (0. 039) (0. 031)
$\Delta\left(\frac{\overline{w}_{M}}{\overline{w}_{F}}\right)$ –0. 249 –0. 004 0. 013 –0. 226 –0. 210 –0. 186
(0. 014) (0. 000) (0. 001) (0. 016) (0. 016) (0. 017)
$\Delta\left(\frac{\overline{w}_{H}}{\overline{w}_{U}}\right)$ 0. 564 0. 009 –0. 031 0. 465 0. 401 0. 319
(0. 013) (0. 000) (0. 001) (0. 015) (0. 016) (0. 017)
Labor shares Labor shares
$\Delta\left(\frac{w_{mh}\ell_{mh}}{py}\right)$ 0. 006 0. 000 0. 000 0. 010 0. 014 0. 019
(0. 001) (0. 000) (0. 000) (0. 001) (0. 001) (0. 001)
$\Delta\left(\frac{w_{fh}\ell_{fh}}{py}\right)$ 0. 017 0. 000 –0. 001 0. 019 0. 020 0. 022
(0. 001) (0. 000) (0. 000) (0. 001) (0. 001) (0. 001)
$\Delta\left(\frac{w_{mu}\ell_{mu}}{py}\right)$ –0. 039 –0. 001 0. 002 –0. 056 –0. 070 –0. 091
(0. 002) (0. 000) (0. 000) (0. 003) (0. 005) (0. 007)
$\Delta\left(\frac{w_{fu}\ell_{fu}}{py}\right)$ –0. 008 0. 000 0. 001 –0. 008 –0. 007 –0. 004
(0. 001) (0. 000) (0. 000) (0. 001) (0. 001) (0. 002)
$\Delta\left(\frac{\sum_{f}w_{f}\ell_{f}}{py}\right)$ –0. 001 0. 000 0. 000 –0. 006 –0. 009 –0. 013
(0. 004) (0. 000) (0. 000) (0. 004) (0. 004) (0. 004)
Notes: The changes in the levels of relative wages and labor shares attributable to technological change in the first to third columns are computed using equations (21) and (22) of Proposition 3, those in the fourth to last columns are computed using equations (A1) and (A2) of Proposition 4 for each value of the labor supply elasticity $1/\gamma$. Standard errors in parentheses are computed using bootstrap with 500 replications.
The last three columns of Table 7 presents the effects of technological change embodied in ICCT capital on the relative wages and income shares of the four types of labor for each value of the labor supply elasticity. For ease of reference, the first column reproduces the first column of Table 7, which corresponds to the case in which the labor supply elasticity is zero. Given the difficulty of selecting a single value of the aggregate hours elasticity (Chetty, Guren, Manoli and Weber, 2011; Keane and Rogerson, 2012), the second to fourth columns show how the effects vary according to the labor supply elasticity in the range of 0.5 to 2. The effects of technological change embodied in ICCT capital on the skilled gender wage gap and the female skill wage gap are greatest in absolute value when the supply of labor is exogenous and become smaller in absolute value as the supply of labor is more elastic. Progress in ICCT raises the value of the marginal product of female skilled labor relative to other types of labor, and thus, the relative wages of female skilled labor. If the supply of labor is endogenous, the quantities of male (female) skilled labor relative to female skilled (unskilled) labor can decrease (increase) in response to progress in ICCT. The decrease (increase) in the supply of male (female) skilled labor relative to female skilled (unskilled) labor moderates a fall (rise) in the skilled gender (female skill) wage gap. Such feedback effects become stronger as the supply of labor is more elastic. However, even if the labor supply elasticity is high at two, the effects of technological change embodied in ICCT capital on the skilled gender wage gap and the female skill wage gap remain greater than their observed changes. Moreover, the effects of technological change embodied in ICCT capital on the unskilled gender wage gap and the male skill wage gap do not vary substantially depending on the value of the labor supply elasticity. Even if the labor supply elasticity is high at two, the effects of technological change embodied in ICCT capital on the unskilled gender wage gap and the male skill wage gap are greater than their observed changes. All else held constant, when the supply of labor is unit elastic, the fall in the relative price of ICCT capital would have reduced the skilled (unskilled) gender wage gap by 49 (10) percentage points and raised the male (female) skill wage gap by 26 (77) percentage points.
The effects of technological change embodied in ICCT capital on the income shares of the four types of labor tend to be smallest in absolute value when the supply of labor is exogenous and become greater in absolute value as the supply of labor is more elastic. The rate of change in the income share of labor of type $f$ can be decomposed as $\Delta\ln\Lambda_{f}=\Delta\ln(w_{f}/p)+\Delta\ln\ell_{f}-\Delta\ln y$. As the labor supply elasticity increases, the rate of increase in equilibrium wages becomes smaller, whereas the rates of increases in equilibrium quantities of labor and output become greater. The effects of progress in ICCT on labor shares become greater (smaller) as the labor supply elasticity increases if the rate of increase in the equilibrium quantities of labor is large (small) relative to the rate of decrease in the equilibrium wages plus the rate of increase in the equilibrium quantities of output. The effects on the income shares of male and female skilled labor increase due to a greater increase in the equilibrium quantities of labor, whereas that on the income shares of male unskilled labor decrease due to a smaller increase in the equilibrium quantities of labor. The effect on the income share of female unskilled labor is insensitive to the value of the labor supply elasticity. All else held constant, the fall in the relative price of ICCT capital would have raised the income share of male (female) skilled labor by 1.4 (2.0) percentage points and reduced the income share of male (female) unskilled labor by 7.0 (0.7) percentage points in the case in which the supply of labor is unit elastic. In this case, the fall in the relative price of ICCT capital accounts for 34 (53) percent of the increase in the income share of male (female) skilled labor and 53 (34) percent of the decline in the income share of male (female) unskilled labor. The effect of technological change embodied in ICCT capital also becomes more visible as the labor supply elasticity increases, although it is obscured when aggregated across different types of labor.
Appendix F presents three sets of additional results on the effects of technological change. First, the magnitude of the effects of technological change embodied in ICCT capital remains largely unchanged regardless of whether labor is perfectly mobile across sectors (Table A2). Second, for all countries, technological change embodied in ICCT capital narrows the gender wage gap and widens the skill wage gap, while it increases the income share of skilled labor and decreases the income share of unskilled labor (Table A3). Finally, the magnitude of the changes in relative wages and income shares attributable to ICCT capital may differ depending on whether they are obtained by estimating and aggregating sectoral production function parameters or by assuming and estimating an aggregate production function (Table A4).
7 Conclusion
We developed a unified framework to evaluate the effects of capital-embodied technological change on, as well as the contributions of factor inputs to, the relative wages and income shares of different types of labor. We applied this framework to advanced economies consisting of multiple sectors, where two types of capital (ICCT and non-ICCT capital) and four types of labor (male skilled, female skilled, male unskilled, and female unskilled labor) are used to produce goods and services.
Our results show that the expansion of ICCT equipment results in an increase in the skill wage gap and a decline in the gender wage gap, while the increased supply of male and female skilled labor prevents an increase in the skill wage gap and a decline in the gender wage gap. The observed patterns of changes in the four wage gaps (skilled gender wage gap, unskilled gender wage gap, male skill wage gap, and female skill wage gap) can be explained in terms of the race between progress in ICCT and advances in educational attainment and female employment. The unskilled gender wage gap is more likely to decline than the skilled gender wage gap, while the male skill wage gap is more likely to increase than the female skill wage gap. The reason for this is not because the demand for female skilled labor is insufficient but because the supply of female skilled labor becomes abundant. Our results further show that the effects of technological change embodied in ICCT equipment are sufficiently large to account for the narrowing of the unskilled gender wage gap, the widening of the male skill wage gap, and about half of the decline in the income share of male unskilled labor that accounts for most of the decline in the labor share.
Our findings have implications for policies aimed at promoting female employment, technological progress, and educational attainment. The role of these policies is similar in that they could help raise national income but different in that they could widen or narrow the gender wage gap and the skill wage gap. The male–female wage gap would decrease as a result of policies that raise the relative demand for female labor (e.g., tax credits for and grant funding to research and development), whereas it would increase as a result of policies that raise the relative supply of female labor (e.g., maternity leave, childcare facilities, and the equal employment opportunity law). At the same time, the skilled–unskilled wage gap would decrease as a result of policies that raise the relative supply of skilled labor (e.g., tuition exemption and subsidies), whereas it would increase as a result of policies that raise the relative demand for skilled labor (e.g., tax credits for and grant funding to research and development). Our findings suggest not only that policies promoting educational attainment would be effective in reducing the skill wage gap but also that policies promoting technological progress would be effective in reducing the gender wage gap.
Appendix
Appendix A Extension
A.1 Endogenous labor supply
We have shown how the equilibrium relative wages and income shares of different types of labor can change in response to technological change under the assumption that the supply of labor is exogenous. In such a case, the equilibrium relative quantities of labor are invariant to technological change. However, if the supply of labor is endogenous, the equilibrium relative quantities of labor can change in response to a change in the equilibrium relative wages due to technological change. In such a general case, the equilibrium relative wages and income shares can undergo further changes as a consequence of changes in the relative quantities of labor. Here, we extend the framework to allow for endogenous labor supply.
We consider a representative household that has preferences over sequences of consumption $\{c_{fnt}\}_{t=0}^{\infty}$ and hours worked $\{l_{ft}\}_{t=0}^{\infty}$ characterized by
$\sum_{t=0}^{\infty}\beta^{t}\sum_{f=1}^{F_{\ell}}\mu_{f}\left[\mathcal{U}\left(c_{ft}\right)-\chi_{f}\frac{l_{ft}^{1+\gamma}}{1+\gamma}\right],$
where $\chi_{f}$ denotes the weight on the disutility of hours worked by an individual of type $f$.
Proposition 4.
Suppose that labor is supplied endogenously. Changes in relative wages due to technological change are given by
$d\ln\left(\frac{w_{f}}{w_{g}}\right)=\sum_{h=1}^{F_{k}}\left(\Xi_{\left(f,h\right)}^{q}-\Xi_{\left(g,h\right)}^{q}\right)d\ln q_{h}+\sum_{n=1}^{N}\left(\Xi_{\left(f,n\right)}^{A}-\Xi_{\left(g,n\right)}^{A}\right)d\ln A_{n}\quad\text{for }f,g=1,\ldots,F_{\ell},$ (A1)
where $\Xi_{\left(f,g\right)}^{q}$ is the $fg$-th element of the $F_{\ell}\times F_{k}$ matrix $\Xi^{q}$, and $\Xi_{\left(f,n\right)}^{A}$ is the $fn$-th element of the $F_{\ell}\times N$ matrix $\Xi^{A}$. The matrices $\Xi^{q}$ and $\Xi^{A}$ are described in Appendix B. Changes in labor shares due to technological change are given by
$d\ln\Lambda_{\ell_{f}}=\sum_{h=1}^{F_{k}}\left\{\sum_{g\neq f}^{F}\left(1-\epsilon_{\ell_{g}\ell_{f}}^{\mathcal{C}}\right)\Lambda_{\ell_{g}}\Xi_{\left(f,h\right)}^{q}+\sum_{g\neq f}^{F_{\ell}}\left[-\left(1-\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}\right)\Lambda_{\ell_{g}}+\sum_{s\neq f,g}^{F}\left(\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}-\epsilon_{\ell_{s}\ell_{g}}^{\mathcal{C}}\right)\Lambda_{\ell_{s}}\right]\Xi_{\left(g,h\right)}^{q}\right.\\ \left.-\left[-\left(1-\epsilon_{\ell_{f}k_{h}}^{\mathcal{C}}\right)\Lambda_{k_{h}}+\sum_{s\neq f,h}^{F}\left(\epsilon_{\ell_{f}k_{h}}^{\mathcal{C}}-\epsilon_{\ell_{s}k_{h}}^{\mathcal{C}}\right)\Lambda_{\ell_{s}}\right]\right\}d\ln q_{h}+\sum_{n=1}^{N}\left\{\Phi_{\left(f,n\right)}^{w}+\sum_{g\neq f}^{F}\left(1-\epsilon_{\ell_{g}\ell_{f}}^{\mathcal{C}}\right)\Lambda_{\ell_{g}}\Xi_{\left(f,n\right)}^{A}\right.\\ \left.+\sum_{g\neq f}^{F_{\ell}}\left[-\left(1-\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}\right)\Lambda_{\ell_{g}}+\sum_{s\neq f,g}^{F}\left(\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}-\epsilon_{\ell_{s}\ell_{g}}^{\mathcal{C}}\right)\Lambda_{\ell_{s}}\right]\Xi_{\left(g,n\right)}^{A}\right\}d\ln A_{n}\quad\text{for }f=1,\ldots,F_{\ell}.$ (A2)
Equations (A1) and (A2) take similar forms to equations (21) and (22), respectively, and nest them as special cases with zero elasticity of labor supply. Most elements of the matrix $\Xi^{q}$ consist of the aggregate elasticities of substitution $\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}$ only or both the aggregate elasticities of substitution and the aggregate elasticity of labor supply $1/\gamma$, and the others consist of the factor shares of income $\Lambda_{\ell_{f}}$. Equations (A1) and (A2) show that the aggregate elasticities of substitution and the aggregate elasticity of labor supply are sufficient statistics for quantifying the effects of technological change on the relative wages and income shares of different types of labor if the income shares of each factor are known.
Again, we consider two examples to demonstrate how the direction and magnitude of the effects of technological change depend on the aggregate elasticities of substitution among different types of capital and labor and the aggregate elasticity of labor supply. While Corollary 2 holds true regardless of whether the supply of labor is exogenous or endogenous, Corollary 3 can be generalized as follows.
Corollary 4.
Suppose that production requires capital equipment, $k_{e}$, capital structures, $k_{s}$, skilled labor, $\ell_{h}$, and unskilled labor, $\ell_{u}$, the aggregate elasticities of substitution satisfy $\epsilon_{\ell_{h}k_{s}}^{\mathcal{C}}=\epsilon_{\ell_{u}k_{s}}^{\mathcal{C}}=1$ and $\epsilon_{\ell_{u}\ell_{h}}^{\mathcal{C}}-\epsilon_{\ell_{h}\ell_{u}}^{\mathcal{C}}=\epsilon_{\ell_{h}k_{e}}^{\mathcal{C}}-\epsilon_{\ell_{u}k_{e}}^{\mathcal{C}}$, and there is no technological change embodied in capital structure, as in Krusell et al. (2000). Changes in the skill premium due to technological change are given by
$d\ln\left(\frac{w_{h}}{w_{u}}\right)=\left(\frac{\epsilon_{\ell_{u}k_{e}}^{\mathcal{C}}-\epsilon_{\ell_{h}k_{e}}^{\mathcal{C}}}{\overline{\epsilon}_{\ell_{h}\ell_{u}}^{\mathcal{C}}+1\left/\gamma\right.}\right)\left(\frac{\Lambda_{k_{e}}+\Lambda_{\ell_{h}}+\Lambda_{\ell_{u}}}{\Lambda_{\ell_{h}}+\Lambda_{\ell_{u}}}\right)d\ln q_{e}\\ +\sum_{n=1}^{N}\left[\left(\frac{1}{\overline{\epsilon}_{\ell_{h}\ell_{u}}^{\mathcal{C}}+1\left/\gamma\right.}\right)\left(1-\frac{1}{1-\eta}\right)\left(\frac{\zeta_{n}\lambda_{\ell_{u}n}}{\Lambda_{\ell_{u}}}-\frac{\zeta_{n}\lambda_{\ell_{h}n}}{\Lambda_{\ell_{h}}}\right)-\left(\frac{\epsilon_{\ell_{u}k_{e}}^{\mathcal{C}}-\epsilon_{\ell_{h}k_{e}}^{\mathcal{C}}}{\overline{\epsilon}_{\ell_{h}\ell_{u}}^{\mathcal{C}}+1\left/\gamma\right.}\right)\left(\frac{\zeta_{n}}{\Lambda_{\ell_{h}}+\Lambda_{\ell_{u}}}\right)\right]d\ln A_{n}.$
The direction of the effect of technological progress embodied in capital equipment on the skill wage gap remains the same regardless of whether the supply of labor is exogenous or endogenous. However, the magnitude of the effect becomes smaller as the labor supply elasticity increases.
A.2 Imperfect labor mobility
We have shown how the equilibrium relative wages and income shares of different types of labor can change in response to technological change under the assumption that labor is perfectly mobile across sectors. This assumption may be reasonable in the long run and useful for expressing the aggregate output as a function of the aggregate factor inputs. However, it may be costly for workers to switch jobs across sectors (Lee and Wolpin, 2006), and sectoral wages are not fully equalized in the data. Here, we relax the assumption of perfect labor mobility across sectors and consider the equilibrium in which the demand for factor $f$ equals the supply of factor $f$ for each sector: $\ell_{fn}=\mu_{f}l_{fn}$. In this case, the equilibrium wages vary across sectors. The gender (skill) wage gap is measured by the ratio of the average wages of male (skilled) labor to those of female (unskilled) labor. The average wages of labor $f$ is defined as $\overline{w}_{f}=\sum_{n=1}^{N}w_{fn}\ell_{fn}/\sum_{n=1}^{N}\ell_{fn}$, where $w_{fn}$ is the wages of labor $f$ in sector $n$. The income share of factor $f$ in the aggregate economy can be written as $\Lambda_{\ell_{f}}=\sum_{n=1}^{N}\Lambda_{\ell_{f}n}$, where the income share of sector $n$’s factor $f$ in the aggregate economy is defined as $\Lambda_{\ell_{f}n}=w_{fn}\ell_{fn}/py$.
Proposition 5.
Suppose that labor is not perfectly mobile across sectors. Changes in the relative wages due to technological change are given by
$d\ln\left(\frac{\overline{w}_{f}}{\overline{w}_{g}}\right)=\sum_{h=1}^{F_{k}}\sum_{n=1}^{N}\left(\frac{\Lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\Omega_{f\left(n,h\right)}^{q}-\frac{\Lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\Omega_{g\left(n,h\right)}^{q}\right)d\ln q_{h}\\ +\sum_{n=1}^{N}\sum_{m=1}^{N}\left(\frac{\Lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\Omega_{f\left(n,m\right)}^{A}-\frac{\Lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\Omega_{g\left(n,m\right)}^{A}\right)d\ln A_{m}\quad\text{for }f,g=1,\ldots,F_{\ell},$ (A3)
where $\Omega_{f(n,g)}^{q}$ is the $ng$-th element of the $N\times F_{k}$ matrix $\Omega_{f}^{q}$, and $\Omega_{f(n,m)}^{A}$ is the $nm$-th element of the $N\times N$ matrix $\Omega_{f}^{A}$. The matrices $\Omega_{f}^{q}$ and $\Omega_{f}^{A}$ are described in the proof below. Changes in labor shares due to technological change are given by
$d\ln\Lambda_{\ell_{f}}=\sum_{h=1}^{F_{k}}\sum_{n=1}^{N}\sum_{m=1}^{N}\frac{\Lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\left(\sum_{g=1}^{F_{\ell}}\Pi_{UL,fg\left(n,m\right)}^{w}\Omega_{g\left(m,h\right)}^{q}-\Pi_{UR,fh\left(n,m\right)}^{w}\right)d\ln q_{h}\\ +\sum_{n=1}^{N}\sum_{m=1}^{N}\frac{\Lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\left(\sum_{g=1}^{F_{\ell}}\sum_{l=1}^{N}\Pi_{UL,fg\left(n,l\right)}^{w}\Omega_{g\left(l,m\right)}^{A}+\Theta_{f\left(n,m\right)}^{w}\right)d\ln A_{m}\quad\text{for }f=1,\ldots,F_{\ell},$ (A4)
where $\Pi_{UR,fg(n,m)}^{w}$ is the $nm$-th element of the $N\times N$ matrix $\Pi_{UR,fg}^{w}$, $\Pi_{UL,fg(n,m)}^{w}$ is the $nm$-th element of the $N\times N$ matrix $\Pi_{UL,fg}^{w}$, and $\Theta_{f}^{w}$ is the $N\times N$ matrix whose $nm$-th element is given by $\Theta_{f(n,m)}^{w}=-[1-1/(1-\eta)](I_{(n,m)}-\zeta_{m})$. The matrices $\Pi_{UR,fg}^{w}$ and $\Pi_{UL,fg}^{w}$ are described in in the proof below.
Most elements of the matrix $\Omega_{f}^{q}$ consist of the sectoral elasticities of substitution defined as $\epsilon_{\ell_{fn}\ell_{gm}}^{\mathcal{C}}=\partial\ln(\ell_{fn}/\ell_{gm})/\partial\ln(w_{gm}/w_{fn})$, and the others are the factor shares of income $\Lambda_{\ell_{f}n}$. All elements of the matrices $\Pi_{UR,fg}^{w}$ and $\Pi_{UL,fg}^{w}$ consist of the sectoral elasticities of substitution $\epsilon_{\ell_{fn}\ell_{gm}}^{\mathcal{C}}$ and the factor shares of income $\Lambda_{\ell_{f}n}$. Equations (A3) and (A4) show that the sectoral elasticities of substitution are sufficient statistics for quantifying the effects of capital-embodied technological change on the relative wages and income shares of different types of labor if the income shares of each factor are known.
Appendix B Proofs
Proof of Lemma 1
Using the facts that $\partial\ln\mathcal{F}(\ell_{1},...,\ell_{F})/\partial\ln\ell_{f}=\Lambda_{\ell_{f}}$ and $\partial\ln\mathcal{C}(w_{1},...,w_{F},y)$ $\partial\ln w_{f}=\Lambda_{\ell_{f}}$, the aggregate elasticities of substitution can be rewritten as
$\frac{1}{\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{F}}}=$ $\frac{\partial\ln\Lambda_{\ell_{f}}}{\partial\ln\ell_{g}}-\frac{\partial\ln\Lambda_{\ell_{g}}}{\partial\ln\ell_{g}}+1,$ (A5)
$\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}=$ $\frac{\partial\ln\Lambda_{\ell_{f}}}{\partial\ln w_{g}}-\frac{\partial\ln\Lambda_{\ell_{g}}}{\partial\ln w_{g}}+1.$ (A6)
The first-order Taylor expansion of the factor share, $\Lambda_{\ell_{f}}=\sum_{n=1}^{N}\zeta_{n}\lambda_{\ell_{f}n}$, yields
$\Delta\ln\Lambda_{\ell_{f}}=\sum_{n=1}^{N}\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\Delta\ln\zeta_{n}+\sum_{n=1}^{N}\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\Delta\ln\lambda_{\ell_{f}n}.$ (A7)
The first-order Taylor expansion of the factor share in sector $n$, $\lambda_{\ell_{f}n}=w_{f}\ell_{fn}/p_{n}y_{n}$, yields
$\Delta\ln\lambda_{\ell_{f}n}=\sum_{g=1}^{F}\frac{\partial\ln\lambda_{\ell_{f}n}}{\partial\ln w_{g}}\Delta\ln w_{g}.$ (A8)
The first-order Taylor expansion of the expenditure share (6) yields
$\Delta\ln\zeta_{n}=-\frac{\eta}{1-\eta}\Delta\ln p_{n}+\frac{\eta}{1-\eta}\sum_{m=1}^{N}\zeta_{m}\Delta\ln p_{m}.$ (A9)
The first-order Taylor expansion of the profit maximizing condition, $p_{n}=\mathcal{\mathcal{\widetilde{C}}}_{n}(w_{1},\ldots,w_{F})/A_{n}$, yields
$\Delta\ln p_{n}=\sum_{f=1}^{F}\lambda_{\ell_{f}n}\Delta\ln w_{f}-\Delta\ln A_{n}.$ (A10)
Substituting this into equation (A9) yields
$\Delta\ln\zeta_{n}=\left(1-\frac{1}{1-\eta}\right)\left[\sum_{f=1}^{F}\left(\lambda_{\ell_{f}n}-\Lambda_{\ell_{f}}\right)\Delta\ln w_{f}-\Delta\ln A_{n}+\sum_{m=1}^{N}\zeta_{m}\Delta\ln A_{m}\right].$ (A11)
Substituting equations (A8) and (A11) into equation (A7) yields
$\Delta\ln\Lambda_{\ell_{f}}=\sum_{g=1}^{F}\Psi_{\left(f,g\right)}^{w}\Delta\ln w_{g}+\psi_{\left(f\right)}^{w}=\sum_{g=1}^{F}\Psi_{\left(f,g\right)}^{w}\Delta\ln\left(\frac{w_{g}}{p}\right)+\psi_{\left(f\right)}^{w}.$ (A12)
The second equality comes from the fact that equation (A12) is homogeneous of degree zero in factor prices because $\sum_{g=1}^{F}\Psi_{\left(f,g\right)}^{w}=0$. By the definition of factor shares, $\Delta\ln w_{f}=\Delta\ln\Lambda_{\ell_{f}}-\Delta\ln\ell_{f}+\Delta\ln(py)$. Substituting this into equation (A12) yields
$\Delta\ln\Lambda_{\ell_{f}}=\sum_{g=1}^{F}\Psi_{\left(f,g\right)}^{w}\Delta\ln\Lambda_{\ell_{g}}-\sum_{g=1}^{F}\Psi_{\left(f,g\right)}^{w}\Delta\ln\ell_{g}+\psi_{\left(f\right)}^{w}.$
Solving the system of equations for $\Delta\ln\Lambda_{\ell_{f}}$ yields
$\Delta\ln\Lambda_{\ell_{f}}=\sum_{g=1}^{F}\Psi_{\left(f,g\right)}^{\ell}\Delta\ln\ell_{g}+\psi_{\left(f\right)}^{\ell}.$ (A13)
Equation (A13) implies $\partial\ln\Lambda_{\ell_{f}}/\partial\ln\ell_{g}=\Psi_{\left(f,g\right)}^{\ell}$, while equation (A12) implies $\partial\ln\Lambda_{\ell_{f}}/\partial\ln w_{g}=\Psi_{\left(f,g\right)}^{w}$. Substituting these into equations (A5) and (A6) yields equations (12) and (13).
Proof of Proposition 1
Using the fact that $\sum_{f=1}^{F}\Lambda_{\ell_{f}}\Psi_{\left(f,g\right)}^{\ell}=0$, equation (A13) can be rearranged as
$\Delta\ln\Lambda_{\ell_{f}}=\sum_{g\neq f}^{F}\left[1-\left(\Psi_{\left(g,f\right)}^{\ell}-\Psi_{\left(f,f\right)}^{\ell}+1\right)\right]\Lambda_{\ell_{g}}\Delta\ln\ell_{f}+\sum_{g\neq f}^{F}\left[\left(\Psi_{\left(f,g\right)}^{\ell}-\Psi_{\left(g,g\right)}^{\ell}+1\right)-1\right]\Lambda_{\ell_{g}}\Delta\ln\ell_{g}\\ +\sum_{g\neq f}^{F}\sum_{h\neq f,g}^{F}\left[\left(\Psi_{\left(f,g\right)}^{\ell}-\Psi_{\left(g,g\right)}^{\ell}+1\right)-\left(\Psi_{\left(h,g\right)}^{\ell}-\Psi_{\left(g,g\right)}^{\ell}+1\right)\right]\Lambda_{\ell_{h}}\Delta\ln\ell_{g}+\psi_{\left(f\right)}^{\ell}.$
Substituting equation (12) into this yields equation (15).
Similarly, using the fact that $\sum_{f=1}^{F}\Lambda_{\ell_{f}}\Psi_{\left(f,g\right)}^{w}=0$, equation (A12) can be rearranged as
$\Delta\ln\Lambda_{\ell_{f}}=\sum_{g\neq f}^{F}\left[1-\left(\Psi_{\left(g,f\right)}^{w}-\Psi_{\left(f,f\right)}^{w}+1\right)\right]\Lambda_{\ell_{g}}\Delta\ln w_{f}+\sum_{g\neq f}^{F}\left[\left(\Psi_{\left(f,g\right)}^{w}-\Psi_{\left(g,g\right)}^{w}+1\right)-1\right]\Lambda_{\ell_{g}}\Delta\ln w_{g}\\ +\sum_{g\neq f}^{F}\sum_{h\neq f,g}^{F}\left[\left(\Psi_{\left(f,g\right)}^{w}-\Psi_{\left(g,g\right)}^{w}+1\right)-\left(\Psi_{\left(h,g\right)}^{w}-\Psi_{\left(g,g\right)}^{w}+1\right)\right]\Lambda_{\ell_{h}}\Delta\ln w_{g}+\psi_{\left(f\right)}^{w}.$
Substituting equation (13) into this yields equation (16).
Proof of Proposition 2
By definition, the log change of relative factor shares can be written as
$\Delta\ln\left(\frac{\Lambda_{\ell_{f}}}{\Lambda_{\ell_{g}}}\right)=\Delta\ln\left(\frac{w_{f}}{w_{g}}\right)+\Delta\ln\left(\frac{\ell_{f}}{\ell_{g}}\right).$ (A14)
Substituting equation (A13) into equation (A14) yields
$\Delta\ln\left(\frac{w_{f}}{w_{g}}\right)=\sum_{h=1}^{F}\left[\Psi_{\left(f,h\right)}^{\ell}-\Psi_{\left(h,h\right)}^{\ell}+1-\left(\Psi_{\left(g,h\right)}^{\ell}-\Psi_{\left(h,h\right)}^{\ell}+1\right)\right]\Delta\ln\ell_{h}-\Delta\ln\ell_{f}+\Delta\ln\ell_{g}+\psi_{\left(f\right)}^{\ell}-\psi_{\left(g\right)}^{\ell}.$
Using equation (12), this equation can be rewritten as equation (18).
Substituting (A12) into equation (A14) yields
$\Delta\ln\left(\frac{\ell_{f}}{\ell_{g}}\right)=\sum_{h=1}^{F}\left[\Psi_{\left(f,h\right)}^{w}-\Psi_{\left(h,h\right)}^{w}+1-\left(\Psi_{\left(g,h\right)}^{w}-\Psi_{\left(h,h\right)}^{w}+1\right)\right]\Delta\ln w_{h}-\Delta\ln w_{f}+\Delta\ln w_{g}+\psi_{\left(f\right)}^{w}-\psi_{\left(g\right)}^{w}.$
Using equation (13), this equation can be rewritten as equation (19).
Proof of Proposition 3
The Euler equation (A34) can be derived from the household’s problem when the interest rate is defined as $\iota_{t}=(1/\beta)[(\partial\mathcal{U}(c_{t-1})/\partial c_{t-1})/(\partial\mathcal{U}(c_{t})/\partial c_{t})]-1$. It follows that the log change in the rental price of capital between steady states are given by
$d\ln\left(\frac{r_{f}}{p}\right)=-d\ln q_{f}.$ (A15)
Subtracting $d\ln p$ from both sides of equation (A10) and substituting equation (A15) into this yields
$d\ln\left(\frac{p_{n}}{p}\right)=\sum_{f=1}^{F_{\ell}}\lambda_{\ell_{f}n}d\ln\left(\frac{w_{f}}{p}\right)-\sum_{f=1}^{F_{k}}\lambda_{k_{f}n}d\ln q_{f}-d\ln A_{n}.$ (A16)
The first-order Taylor expansion of the aggregate price, $p=(\sum_{n=1}^{N}\theta_{cn}^{1/(1-\eta)}p_{n}^{-\eta/(1-\eta)})^{-(1-\eta)/\eta}$, yields $\sum_{n=1}^{N}\zeta_{n}d\ln(p_{n}/p)=0$. Substituting equation (A16) into this yields
$\sum_{f=1}^{F_{\ell}}\Lambda_{\ell_{f}}d\ln\left(\frac{w_{f}}{p}\right)=\sum_{f=1}^{F_{k}}\Lambda_{k_{f}}d\ln q_{f}+\sum_{n=1}^{N}\zeta_{n}d\ln A_{n}.$ (A17)
Equation (A12) can be rewritten as
$d\ln\Lambda_{\ell_{f}}=\sum_{g=1}^{F}\Psi_{\left(f,g\right)}^{w}d\ln\left(\frac{w_{g}}{p}\right)+\sum_{n=1}^{N}\Phi_{\left(f,n\right)}^{w}d\ln A_{n},$
where $\Phi^{w}$ is the $F\times N$ matrix whose $fn$-th element is given by $\Phi_{\left(f,n\right)}^{w}=-[1-1/(1-\eta)](\zeta_{n}\lambda_{\ell_{f}n}/\Lambda_{\ell_{f}}-\zeta_{n})$. Let the matrix $\Psi^{w}$ be partitioned as
$\Psi^{w}=\left[\begin{array}[]{cc}\Psi_{UL}^{w}&\Psi_{UR}^{w}\\ \Psi_{LL}^{w}&\Psi_{LR}^{w}\end{array}\right],$
where $\Psi_{UL}^{w}$, $\Psi_{UR}^{w}$, $\Psi_{LL}^{w}$, and $\Psi_{LR}^{w}$ are $_{\ell}\times F_{\ell}, F×F$$_{\ell}\times F_{k}$, $F_{k}\times F_{\ell}$, and $F_{k}\times F_{k}$ matrices, respectively. The equation can be further rewritten as
$d\ln\Lambda_{\ell_{f}}=\sum_{g=1}^{F_{\ell}}\Psi_{UL\left(f,g\right)}^{w}d\ln\left(\frac{w_{g}}{p}\right)+\sum_{g=1}^{F_{k}}\Psi_{UR\left(f,g\right)}^{w}d\ln\left(\frac{r_{g}}{p}\right)+\sum_{n=1}^{N}\Phi_{\left(f,n\right)}^{w}d\ln A_{n}\quad\text{for }f=1,\ldots,F_{\ell}.$
Substituting equation (A15) into this yields
$d\ln\Lambda_{\ell_{f}}=\sum_{g=1}^{F_{\ell}}\Psi_{UL\left(f,g\right)}^{w}d\ln\left(\frac{w_{g}}{p}\right)-\sum_{g=1}^{F_{k}}\Psi_{UR\left(f,g\right)}^{w}d\ln q_{g}+\sum_{n=1}^{N}\Phi_{\left(f,n\right)}^{w}d\ln A_{n}\quad\text{for }f=1,\ldots,F_{\ell}.$ (A18)
Substituting $\Lambda_{\ell_{f}}=w_{f}\ell_{f}/py$ into this yields
$\sum_{g=1}^{F_{\ell}}\left(\Psi_{UL\left(f,g\right)}^{w}-I_{\left(f,g\right)}\right)d\ln\left(\frac{w_{g}}{p}\right)=\sum_{g=1}^{F_{\ell}}I_{\left(f,g\right)}d\ln\ell_{g}\\ +\sum_{g=1}^{F_{k}}\Psi_{UR\left(f,g\right)}^{w}d\ln q_{g}-\sum_{n=1}^{N}\Phi_{\left(f,n\right)}^{w}d\ln A_{n}-d\ln y\quad\text{for }f=1,\ldots,F_{\ell}.$ (A19)
Subtracting the equation for $f=F_{\ell}$ from that for any $f\in\{1,\ldots,F_{\ell}-1$} yields
$\sum_{g=1}^{F_{\ell}}\left(\Psi_{UL\left(f,g\right)}^{w}-\Psi_{UL\left(F_{\ell},g\right)}^{w}-I_{\left(f,g\right)}+I_{\left(F_{\ell},g\right)}\right)d\ln\left(\frac{w_{g}}{p}\right)=\sum_{g=1}^{F_{\ell}}\left(I_{\left(f,g\right)}-I_{\left(F_{\ell},g\right)}\right)d\ln\ell_{g}\\ +\sum_{g=1}^{F_{k}}\left(\Psi_{UR\left(f,g\right)}^{w}-\Psi_{UR\left(F_{\ell},g\right)}^{w}\right)d\ln q_{g}-\sum_{n=1}^{N}\left(\Phi_{\left(f,n\right)}^{w}-\Phi_{\left(F_{\ell},n\right)}^{w}\right)d\ln A_{n}.$ (A20)
Solving the system of equations (A17) and (A20) for $d\ln(w_{f}/p)$ and setting $d\ln\ell_{f}=0$ yields
$d\ln\left(\frac{w_{f}}{p}\right)=\sum_{g=1}^{F_{k}}\Upsilon_{\left(f,g\right)}^{q}d\ln q_{g}+\sum_{n=1}^{N}\Upsilon_{\left(f,n\right)}^{A}d\ln A_{n}\quad\text{for }f=1,\ldots,F_{\ell},$ (A21)
where $\Upsilon_{(f,g)}^{q}$ is the $fg$-th element of the $F_{\ell}\times F_{k}$ matrix
$\Upsilon^{q}=$ $\left[\begin{array}[]{ccccc}-\epsilon_{\ell_{F_{\ell}}\ell_{1}}^{\mathcal{C}}&\epsilon_{\ell_{1}\ell_{2}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}\ell_{2}}^{\mathcal{C}}&\cdots&\epsilon_{\ell_{1}\ell_{F_{\ell}-1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}\ell_{F_{\ell}-1}}^{\mathcal{C}}&\epsilon_{\ell_{1}\ell_{F_{\ell}}}^{\mathcal{C}}\\ \epsilon_{\ell_{2}\ell_{1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}\ell_{1}}^{\mathcal{C}}&-\epsilon_{\ell_{F_{\ell}}\ell_{2}}^{\mathcal{C}}&\cdots&\epsilon_{\ell_{2}\ell_{F_{\ell}-1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}\ell_{F_{\ell}-1}}^{\mathcal{C}}&\epsilon_{\ell_{2}\ell_{F_{\ell}}}^{\mathcal{C}}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ \epsilon_{\ell_{F_{\ell}-1}\ell_{1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}\ell_{1}}^{\mathcal{C}}&\epsilon_{\ell_{F_{\ell}-1}\ell_{2}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}\ell_{2}}^{\mathcal{C}}&\cdots&-\epsilon_{\ell_{F_{\ell}}\ell_{F_{\ell}-1}}^{\mathcal{C}}&\epsilon_{\ell_{F_{\ell}-1}\ell_{F_{\ell}}}^{\mathcal{C}}\\ \Lambda_{\ell_{1}}&\Lambda_{\ell_{2}}&\cdots&\Lambda_{\ell_{F_{\ell}-1}}&\Lambda_{\ell_{F_{\ell}}}\end{array}\right]^{-1}$
$\times\left[\begin{array}[]{ccc}\epsilon_{\ell_{1}k_{1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}k_{1}}^{\mathcal{C}}&\cdots&\epsilon_{\ell_{1}k_{F_{k}}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}k_{F_{k}}}^{\mathcal{C}}\\ \vdots&\vdots&\vdots\\ \epsilon_{\ell_{F_{\ell}-1}k_{1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}k_{1}}^{\mathcal{C}}&\cdots&\epsilon_{\ell_{F_{\ell}-1}k_{F_{k}}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}k_{F_{k}}}^{\mathcal{C}}\\ \Lambda_{k_{1}}&\cdots&\Lambda_{k_{F_{k}}}\end{array}\right],$
and $\Upsilon_{(f,n)}^{A}$ is the $fn$-th element of the $F_{\ell}\times N$ matrix
$\Upsilon^{A}=$ $-\left[\begin{array}[]{ccccc}-\epsilon_{\ell_{F_{\ell}}\ell_{1}}^{\mathcal{C}}&\epsilon_{\ell_{1}\ell_{2}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}\ell_{2}}^{\mathcal{C}}&\cdots&\epsilon_{\ell_{1}\ell_{F_{\ell}-1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}\ell_{F_{\ell}-1}}^{\mathcal{C}}&\epsilon_{\ell_{1}\ell_{F_{\ell}}}^{\mathcal{C}}\\ \epsilon_{\ell_{2}\ell_{1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}\ell_{1}}^{\mathcal{C}}&-\epsilon_{\ell_{F_{\ell}}\ell_{2}}^{\mathcal{C}}&\cdots&\epsilon_{\ell_{2}\ell_{F_{\ell}-1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}\ell_{F_{\ell}-1}}^{\mathcal{C}}&\epsilon_{\ell_{2}\ell_{F_{\ell}}}^{\mathcal{C}}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ \epsilon_{\ell_{F_{\ell}-1}\ell_{1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}\ell_{1}}^{\mathcal{C}}&\epsilon_{\ell_{F_{\ell}-1}\ell_{2}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}\ell_{2}}^{\mathcal{C}}&\cdots&-\epsilon_{\ell_{F_{\ell}}\ell_{F_{\ell}-1}}^{\mathcal{C}}&\epsilon_{\ell_{F_{\ell}-1}\ell_{F_{\ell}}}^{\mathcal{C}}\\ \Lambda_{\ell_{1}}&\Lambda_{\ell_{2}}&\cdots&\Lambda_{\ell_{F_{\ell}-1}}&\Lambda_{\ell_{F_{\ell}}}\end{array}\right]^{-1}$
$\times\left[\begin{array}[]{ccc}\Phi_{\left(1,1\right)}^{w}-\Phi_{\left(F_{\ell},1\right)}^{w}&\cdots&\Phi_{\left(1,N\right)}^{w}-\Phi_{\left(F_{\ell},N\right)}^{w}\\ \vdots&\vdots&\vdots\\ \Phi_{\left(F_{\ell}-1,1\right)}^{w}-\Phi_{\left(F_{\ell},1\right)}^{w}&\cdots&\Phi_{\left(F_{\ell}-1,N\right)}^{w}-\Phi_{\left(F_{\ell},N\right)}^{w}\\ \zeta_{1}&\cdots&\zeta_{N}\end{array}\right].$
Equation (A21) immediately implies equation (21). Substituting equation (A21) into equation (A18) yields equation (22).
Proof of Proposition 4
The labor supply function can be derived from the household’s problem as $l_{f}=(\partial\mathcal{U}(c)/\partial c)^{1/\gamma}\chi_{f}^{-1/\gamma}(w_{f}/p)^{1/\gamma}$. Substituting the labor market clearing condition, $\ell_{f}=\mu_{f}l_{f}$, into this yields $\ell_{f}=\mu_{f}(\partial\mathcal{U}(c)/\partial c)^{1/\gamma}\chi_{f}^{-1/\gamma}(w_{f}/p)^{1/\gamma}$. The log-linearization yields
$d\ln\ell_{f}=\frac{c}{\gamma}\frac{\partial^{2}\mathcal{U}\left(c\right)\left/\partial c^{2}\right.}{\partial\mathcal{U}\left(c\right)\left/\partial c\right.}d\ln c+\frac{1}{\gamma}\sum_{g=1}^{F_{\ell}}I_{\left(f,g\right)}d\ln\left(\frac{w_{g}}{p}\right).$
Substituting this into equation (A19) yields
$\sum_{g=1}^{F_{\ell}}\left[\Psi_{UL\left(f,g\right)}^{w}-\left(\frac{1+\gamma}{\gamma}\right)I_{\left(f,g\right)}\right]d\ln\left(\frac{w_{g}}{p}\right)=-d\ln y+\frac{c}{\gamma}\frac{\partial^{2}\mathcal{U}\left(c\right)\left/\partial c^{2}\right.}{\partial\mathcal{U}\left(c\right)\left/\partial c\right.}d\ln c\\ +\sum_{g=1}^{F_{k}}\Psi_{UR\left(f,g\right)}^{w}d\ln q_{g}-\sum_{n=1}^{N}\Phi_{\left(f,n\right)}^{w}d\ln A_{n}\quad\text{for }f=1,\ldots,F_{\ell}.$
Subtracting the equation for $f=F_{\ell}$ from that for any $f\in\{1,\ldots,F_{\ell}-1$} yields
$\sum_{g=1}^{F_{\ell}}\left[\left(\Psi_{UL\left(f,g\right)}^{w}-\Psi_{UL\left(F_{\ell},g\right)}^{w}\right)-\left(I_{\left(f,g\right)}-I_{\left(F_{\ell},g\right)}\right)\left(\frac{1+\gamma}{\gamma}\right)\right]d\ln\left(\frac{w_{g}}{p}\right)\\ =\sum_{g=1}^{F_{k}}\left(\Psi_{UR\left(f,g\right)}^{w}-\Psi_{UR\left(F_{\ell},g\right)}^{w}\right)d\ln q_{g}-\sum_{n=1}^{N}\left(\Phi_{\left(f,n\right)}^{w}-\Phi_{\left(F_{\ell},n\right)}^{w}\right)d\ln A_{n}.$ (A22)
Solving the system of equations (A17) and (A22) for $d\ln(w_{f}/p)$ yields
$d\ln\left(\frac{w_{f}}{p}\right)=\sum_{g=1}^{F_{k}}\Xi_{\left(f,g\right)}^{q}d\ln q_{g}+\sum_{n=1}^{N}\Xi_{\left(f,n\right)}^{A}d\ln A_{n}\quad\text{for }f=1,\ldots,F_{\ell},$ (A23)
where $\Xi_{(f,g)}^{q}$ is the $fg$-th element of the $F_{\ell}\times F_{k}$ matrix
$\Xi^{q}=$ $\left[\begin{array}[]{ccccc}-\epsilon_{\ell_{F_{\ell}}\ell_{1}}^{\mathcal{C}}-\frac{1}{\gamma}&\epsilon_{\ell_{1}\ell_{2}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}\ell_{2}}^{\mathcal{C}}&\cdots&\epsilon_{\ell_{1}\ell_{F_{\ell}-1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}\ell_{F_{\ell}-1}}^{\mathcal{C}}&\epsilon_{\ell_{1}\ell_{F_{\ell}}}^{\mathcal{C}}+\frac{1}{\gamma}\\ \epsilon_{\ell_{2}\ell_{1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}\ell_{1}}^{\mathcal{C}}&-\epsilon_{\ell_{F_{\ell}}\ell_{2}}^{\mathcal{C}}-\frac{1}{\gamma}&\cdots&\epsilon_{\ell_{2}\ell_{F_{\ell}-1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}\ell_{F_{\ell}-1}}^{\mathcal{C}}&\epsilon_{\ell_{2}\ell_{F_{\ell}}}^{\mathcal{C}}+\frac{1}{\gamma}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ \epsilon_{\ell_{F_{\ell}-1}\ell_{1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}\ell_{1}}^{\mathcal{C}}&\epsilon_{\ell_{F_{\ell}-1}\ell_{2}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}\ell_{2}}^{\mathcal{C}}&\cdots&-\epsilon_{\ell_{F_{\ell}}\ell_{F_{\ell}-1}}^{\mathcal{C}}-\frac{1}{\gamma}&\epsilon_{\ell_{F_{\ell}-1}\ell_{F_{\ell}}}^{\mathcal{C}}+\frac{1}{\gamma}\\ \Lambda_{\ell_{1}}&\Lambda_{\ell_{2}}&\cdots&\Lambda_{\ell_{F_{\ell}-1}}&\Lambda_{\ell_{F_{\ell}}}\end{array}\right]^{-1}$
$\times\left[\begin{array}[]{ccc}\epsilon_{\ell_{1}k_{1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}k_{1}}^{\mathcal{C}}&\cdots&\epsilon_{\ell_{1}k_{F_{k}}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}k_{F_{k}}}^{\mathcal{C}}\\ \vdots&\vdots&\vdots\\ \epsilon_{\ell_{F_{\ell}-1}k_{1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}k_{1}}^{\mathcal{C}}&\cdots&\epsilon_{\ell_{F_{\ell}-1}k_{F_{k}}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}k_{F_{k}}}^{\mathcal{C}}\\ \Lambda_{k_{1}}&\cdots&\Lambda_{k_{F_{k}}}\end{array}\right],$
and $\Xi_{(f,n)}^{A}$ is the $fn$-th element of the $F_{\ell}\times N$ matrix
$\Xi^{A}=$ $-\left[\begin{array}[]{ccccc}-\epsilon_{\ell_{F_{\ell}}\ell_{1}}^{\mathcal{C}}-\frac{1}{\gamma}&\epsilon_{\ell_{1}\ell_{2}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}\ell_{2}}^{\mathcal{C}}&\cdots&\epsilon_{\ell_{1}\ell_{F_{\ell}-1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}\ell_{F_{\ell}-1}}^{\mathcal{C}}&\epsilon_{\ell_{1}\ell_{F_{\ell}}}^{\mathcal{C}}+\frac{1}{\gamma}\\ \epsilon_{\ell_{2}\ell_{1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}\ell_{1}}^{\mathcal{C}}&-\epsilon_{\ell_{F_{\ell}}\ell_{2}}^{\mathcal{C}}-\frac{1}{\gamma}&\cdots&\epsilon_{\ell_{2}\ell_{F_{\ell}-1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}\ell_{F_{\ell}-1}}^{\mathcal{C}}&\epsilon_{\ell_{2}\ell_{F_{\ell}}}^{\mathcal{C}}+\frac{1}{\gamma}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ \epsilon_{\ell_{F_{\ell}-1}\ell_{1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}\ell_{1}}^{\mathcal{C}}&\epsilon_{\ell_{F_{\ell}-1}\ell_{2}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}\ell_{2}}^{\mathcal{C}}&\cdots&-\epsilon_{\ell_{F_{\ell}}\ell_{F_{\ell}-1}}^{\mathcal{C}}-\frac{1}{\gamma}&\epsilon_{\ell_{F_{\ell}-1}\ell_{F_{\ell}}}^{\mathcal{C}}+\frac{1}{\gamma}\\ \Lambda_{\ell_{1}}&\Lambda_{\ell_{2}}&\cdots&\Lambda_{\ell_{F_{\ell}-1}}&\Lambda_{\ell_{F_{\ell}}}\end{array}\right]^{-1}$
$\times\left[\begin{array}[]{ccc}\Phi_{\left(1,1\right)}^{w}-\Phi_{\left(F_{\ell},1\right)}^{w}&\cdots&\Phi_{\left(1,N\right)}^{w}-\Phi_{\left(F_{\ell},N\right)}^{w}\\ \vdots&\vdots&\vdots\\ \Phi_{\left(F_{\ell}-1,1\right)}^{w}-\Phi_{\left(F_{\ell},1\right)}^{w}&\cdots&\Phi_{\left(F_{\ell}-1,N\right)}^{w}-\Phi_{\left(F_{\ell},N\right)}^{w}\\ \zeta_{1}&\cdots&\zeta_{N}\end{array}\right].$
Equation (A23) immediately implies equation (A1). Substituting equation (A23) into equation (A18) yields equation (A2).
Proof of Proposition 5
The log-linearization of the profit maximizing condition, $p_{n}=\mathcal{\mathcal{\widetilde{C}}}_{n}(w_{1n},\ldots,w_{Fn})$ $A_{n}$, yields
$d\ln p_{n}=\sum_{f=1}^{F}\lambda_{\ell_{f}n}d\ln w_{fn}-d\ln A_{n}.$ (A24)
Subtracting $d\ln p$ from both sides of equation (A24) and substituting equation (A15) into this yields
$d\ln\left(\frac{p_{n}}{p}\right)=\sum_{f=1}^{F_{\ell}}\lambda_{\ell_{f}n}d\ln\left(\frac{w_{fn}}{p}\right)-\sum_{f=1}^{F_{k}}\lambda_{k_{f}n}d\ln q_{f}-d\ln A_{n}.$ (A25)
The log-linearization of the aggregate price, $p=(\sum_{n=1}^{N}\theta_{cn}^{1/(1-\eta)}p_{n}^{-\eta/(1-\eta)})^{-(1-\eta)/\eta}$, yields $\sum_{n=1}^{N}\zeta_{n}d\ln(p_{n}/p)$ =0=0. Substituting equation (A25) into this yields
$\sum_{f=1}^{F_{\ell}}\sum_{n=1}^{N}\zeta_{n}\lambda_{\ell_{f}n}d\ln\left(\frac{w_{fn}}{p}\right)=\sum_{f=1}^{F_{k}}\Lambda_{k_{f}}d\ln q_{f}+\sum_{n=1}^{N}\zeta_{n}d\ln A_{n}.$ (A26)
The log-linearization of sector $n$’s factor share, $\Lambda_{\ell_{f}n}=\zeta_{n}\lambda_{\ell_{f}n}$, yields
$d\ln\Lambda_{\ell_{f}n}=d\ln\zeta_{n}+d\ln\lambda_{\ell_{f}n}.$ (A27)
The expenditure share takes the same form as equation (6). The log-linearization of equation (6) yields $d\ln\zeta_{n}=-[\eta/(1-\eta)]d\ln p_{n}+[\eta/(1-\eta)]\sum_{m=1}^{N}\zeta_{m}d\ln p_{m}$. Substituting equation (A24) into this yields
$d\ln\zeta_{n}=\left(1-\frac{1}{1-\eta}\right)\left[\sum_{f=1}^{F}\left(\lambda_{\ell_{f}n}d\ln w_{fn}-\sum_{m=1}^{N}\zeta_{m}\lambda_{\ell_{f}m}d\ln w_{fm}\right)-\left(d\ln A_{n}-\sum_{m=1}^{N}\zeta_{m}d\ln A_{m}\right)\right].$ (A28)
The log-linearization of the factor share in sector $n$, $\lambda_{\ell_{f}n}=w_{fn}\ell_{fn}/p_{n}y_{n}$, yields
$d\ln\lambda_{\ell_{f}n}=\sum_{g=1}^{F}\frac{\partial\ln\lambda_{\ell_{f}n}}{\partial\ln w_{gn}}d\ln w_{gn}.$ (A29)
Substituting equations (A28) and (A29) into equation (A27) yields
$d\ln\Lambda_{\ell_{f}n}=\sum_{g=1}^{F}\sum_{m=1}^{N}\Pi_{fg\left(n,m\right)}^{w}d\ln\left(\frac{w_{gm}}{p}\right)+\sum_{m=1}^{N}\Theta_{f\left(n,m\right)}^{w}d\ln A_{m},$ (A30)
where $\Pi_{fg(n,m)}^{w}=I_{(n,m)}(\partial\ln\lambda_{\ell_{f}m}/\partial\ln w_{gm}+[1-1/(1-\eta)]\lambda_{\ell_{g}m})-[1-1/(1-\eta)]\zeta_{m}\lambda_{\ell_{g}m}$ and $\Theta_{f(n,m)}^{w}=-(1-1/(1-\eta))(I_{(n,m)}-\zeta_{m})$. Define the $FN\times FN$ matrix
$\Pi^{w}=\left[\begin{array}[]{cccccc}\Pi_{11}^{w}&\cdots&\Pi_{1F_{\ell}}^{w}&\Pi_{1,F_{\ell}+1}^{w}&\cdots&\Pi_{1,F_{\ell}+F_{k}}^{w}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ \Pi_{F_{\ell}1}^{w}&\cdots&\Pi_{F_{\ell}F_{\ell}}^{w}&\Pi_{F_{\ell},F_{\ell}+1}^{w}&\cdots&\Pi_{F_{\ell},F_{\ell}+F_{k}}^{w}\\ \Pi_{F_{\ell}+1,1}^{w}&\cdots&\Pi_{F_{\ell}+1,F_{\ell}}^{w}&\Pi_{F_{\ell}+1,F_{\ell}+1}^{w}&\cdots&\Pi_{F_{\ell}+1,F_{\ell}+F_{k}}^{w}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ \Pi_{F_{\ell}+F_{k},1}^{w}&\cdots&\Pi_{F_{\ell}+F_{k},F_{\ell}}^{w}&\Pi_{F_{\ell}+F_{k},F_{\ell}+1}^{w}&\cdots&\Pi_{F_{\ell}+F_{k},F_{\ell}+F_{k}}^{w}\end{array}\right],$
whose $fg$-th block matrix is
$\Pi_{fg}^{w}=\left[\begin{array}[]{ccc}\Pi_{fg(1,1)}^{w}&\cdots&\Pi_{fg(1,N)}^{w}\\ \vdots&\ddots&\vdots\\ \Pi_{fg(N,1)}^{w}&\cdots&\Pi_{fg(N,N)}^{w}\end{array}\right],$
and partition it as
$\Pi^{w}=\left[\begin{array}[]{cc}\Pi_{UL}^{w}&\Pi_{UR}^{w}\\ \Pi_{LL}^{w}&\Pi_{LR}^{w}\end{array}\right],$
where $\Pi_{UL}^{w}$, $\Pi_{UR}^{w}$, $\Pi_{LL}^{w}$, and $\Pi_{LR}^{w}$ are $F_{\ell}N\times F_{\ell}N$, $F_{\ell}N\times F_{k}N$, $F_{k}N\times F_{\ell}N$, and $F_{k}N\times F_{k}N$ matrices, respectively. The upper-left and the upper-right submatrices are
$\Pi_{UL}^{w}=$ $\left[\begin{array}[]{ccc}\Pi_{UL,11}^{w}&\cdots&\Pi_{UL,1F_{\ell}}^{w}\\ \vdots&\ddots&\vdots\\ \Pi_{UL,F_{\ell}1}^{w}&\cdots&\Pi_{UL,F_{\ell}F_{\ell}}^{w}\end{array}\right]=\left[\begin{array}[]{ccc}\Pi_{11}^{w}&\cdots&\Pi_{1F_{\ell}}^{w}\\ \vdots&\ddots&\vdots\\ \Pi_{F_{\ell}1}^{w}&\cdots&\Pi_{F_{\ell}F_{\ell}}^{w}\end{array}\right],$
$\Pi_{UR}^{w}=$ $\left[\begin{array}[]{ccc}\Pi_{UR,11}^{w}&\cdots&\Pi_{UR,1F_{k}}^{w}\\ \vdots&\ddots&\vdots\\ \Pi_{UR,F_{\ell}1}^{w}&\cdots&\Pi_{UR,F_{\ell}F_{k}}^{w}\end{array}\right]=\left[\begin{array}[]{ccc}\Pi_{1,F_{\ell}+1}^{w}&\cdots&\Pi_{1,F_{\ell}+F_{k}}^{w}\\ \vdots&\ddots&\vdots\\ \Pi_{F_{\ell},F_{\ell}+1}^{w}&\cdots&\Pi_{F_{\ell},F_{\ell}+F_{k}}^{w}\end{array}\right],$
where both $\Pi_{UL,fg}^{w}$ and $\Pi_{UR,fg}^{w}$ are $N\times N$ matrices. Equation (A30) can then be rewritten as
$d\ln\Lambda_{\ell_{f}n}=\sum_{g=1}^{F_{\ell}}\sum_{m=1}^{N}\Pi_{UL,fg\left(n,m\right)}^{w}d\ln\left(\frac{w_{gm}}{p}\right)+\sum_{g=1}^{F_{k}}\sum_{m=1}^{N}\Pi_{UR,fg\left(n,m\right)}^{w}d\ln\left(\frac{r_{g}}{p}\right)+\sum_{m=1}^{N}\Theta_{f\left(n,m\right)}^{w}d\ln A_{m}\\ \text{for }f=1,\ldots,F_{\ell}.$
Substituting equation (A15) into this yields
$d\ln\Lambda_{\ell_{f}n}=\sum_{g=1}^{F_{\ell}}\sum_{m=1}^{N}\Pi_{UL,fg\left(n,m\right)}^{w}d\ln\left(\frac{w_{gm}}{p}\right)-\sum_{g=1}^{F_{k}}\sum_{m=1}^{N}\Pi_{UR,fg\left(n,m\right)}^{w}d\ln q_{g}+\sum_{m=1}^{N}\Theta_{f\left(n,m\right)}^{w}d\ln A_{m}\\ \text{for }f=1,\ldots,F_{\ell}.$ (A31)
Substituting $\Lambda_{\ell_{f}n}=w_{fn}\ell_{fn}/py$ into this yields
$\sum_{g=1}^{F_{\ell}}\sum_{m=1}^{N}\left(\Pi_{UL,fg\left(n,m\right)}^{w}-I_{\left(f,g\right)}I_{\left(n,m\right)}\right)d\ln\left(\frac{w_{gm}}{p}\right)=\sum_{g=1}^{F_{\ell}}\sum_{m=1}^{N}I_{\left(f,g\right)}I_{\left(n,m\right)}d\ln\ell_{gm}\\ +\sum_{g=1}^{F_{k}}\sum_{m=1}^{N}\Pi_{UR,fg\left(n,m\right)}^{w}d\ln q_{g}-\sum_{m=1}^{N}\Theta_{f\left(n,m\right)}^{w}d\ln A_{m}-d\ln y\quad\text{for }f=1,\ldots,F_{\ell}.$
Subtracting the equation for $(f,n)=(F_{\ell},N)$ from that for any $f\in\{1,\ldots,F_{\ell}-1\}$ and $n\in\{1,\ldots,N\}$ or for any $f=F_{\ell}$ and $n\in\{1,\ldots,N-1\}$ yields
$\sum_{g=1}^{F_{\ell}}\sum_{m=1}^{N}\left[\Pi_{UL,fg\left(n,m\right)}^{w}-\Pi_{UL,F_{\ell}g\left(N,m\right)}^{w}-I_{\left(f,g\right)}I_{\left(n,m\right)}+I_{\left(F_{\ell},g\right)}I_{\left(N,m\right)}\right]d\ln\left(\frac{w_{gm}}{p}\right)\\ =\sum_{g=1}^{F_{\ell}}\sum_{m=1}^{N}\left(I_{\left(f,g\right)}I_{\left(n,m\right)}-I_{\left(F_{\ell},g\right)}I_{\left(N,m\right)}\right)d\ln\ell_{gm}+\sum_{g=1}^{F_{k}}\sum_{m=1}^{N}\left(\Pi_{UR,fg\left(n,m\right)}^{w}-\Pi_{UR,F_{\ell}g\left(N,m\right)}^{w}\right)d\ln q_{g}\\ -\sum_{m=1}^{N}\left(\Theta_{f\left(n,m\right)}^{w}-\Theta_{F_{\ell}\left(N,m\right)}^{w}\right)d\ln A_{m}.$ (A32)
Solving the system of equations (A26) and (A32) for $d\ln(w_{fn}/p)$ and setting $d\ln\ell_{fn}=0$ yields
$d\ln\left(\frac{w_{fn}}{p}\right)=\sum_{g=1}^{F_{k}}\Omega_{f\left(n,g\right)}^{q}d\ln q_{g}+\sum_{m=1}^{N}\Omega_{f\left(n,m\right)}^{A}d\ln A_{m},$ (A33)
where $\Omega_{f(n,g)}^{q}$ is the $ng$-th element of the $N\times F_{k}$ submatrix $\Omega_{f}^{q}$ of the $F_{\ell}N\times F_{k}$ matrix
$\Omega^{q}=$ $\left[\begin{array}[]{ccc}\left(\Omega_{1}^{q}\right)^{\prime}&\cdots&\left(\Omega_{F_{\ell}}^{q}\right)^{\prime}\end{array}\right]^{\prime}$
$=$ $\left[\begin{array}[]{ccccc}-\epsilon_{\ell_{F_{\ell}N}\ell_{11}}^{\mathcal{C}}&\epsilon_{\ell_{11}\ell_{12}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}N}\ell_{12}}^{\mathcal{C}}&\cdots&\epsilon_{\ell_{11}\ell_{F_{\ell},N-1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}N}\ell_{F_{\ell},N-1}}^{\mathcal{C}}&\epsilon_{\ell_{11}\ell_{F_{\ell}N}}^{\mathcal{C}}\\ \epsilon_{\ell_{12}\ell_{11}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}N}\ell_{11}}^{\mathcal{C}}&-\epsilon_{\ell_{F_{\ell}N}\ell_{12}}^{\mathcal{C}}&\cdots&\epsilon_{\ell_{12}\ell_{F_{\ell},N-1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}N}\ell_{F_{\ell},N-1}}^{\mathcal{C}}&\epsilon_{\ell_{12}\ell_{F_{\ell}N}}^{\mathcal{C}}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ \epsilon_{\ell_{F_{\ell},N-1}\ell_{11}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}N}\ell_{11}}^{\mathcal{C}}&\epsilon_{\ell_{F_{\ell},N-1}\ell_{12}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}N}\ell_{12}}^{\mathcal{C}}&\cdots&-\epsilon_{\ell_{F_{\ell}N}\ell_{F_{\ell},N-1}}^{\mathcal{C}}&\epsilon_{\ell_{F_{\ell},N-1}\ell_{F_{\ell}N}}^{\mathcal{C}}\\ \Lambda_{\ell_{1}1}&\Lambda_{\ell_{1}2}&\cdots&\Lambda_{\ell_{F_{\ell}}N-1}&\Lambda_{\ell_{F_{\ell}}N}\end{array}\right]^{-1}$
$\times\left[\begin{array}[]{ccc}\epsilon_{\ell_{11}k_{1N}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}N}k_{1N}}^{\mathcal{C}}&\cdots&\epsilon_{\ell_{11}k_{F_{k}N}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}N}k_{F_{k}N}}^{\mathcal{C}}\\ \vdots&\vdots&\vdots\\ \epsilon_{\ell_{F_{\ell},N-1}k_{1N}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}N}k_{1N}}^{\mathcal{C}}&\cdots&\epsilon_{\ell_{F_{\ell},N-1}k_{F_{k}N}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}N}k_{F_{k}N}}^{\mathcal{C}}\\ \Lambda_{k_{1}}&\cdots&\Lambda_{k_{F_{k}}}\end{array}\right],$
and $\Omega_{f(n,m)}^{A}$ is the $nm$-th element of the $N\times N$ submatrix $\Omega_{f}^{A}$ of the $F_{\ell}N\times N$ matrix
$\Omega^{A}=$ $\left[\begin{array}[]{ccc}\left(\Omega_{1}^{A}\right)^{\prime}&\cdots&\left(\Omega_{F_{\ell}}^{A}\right)^{\prime}\end{array}\right]^{\prime}$
$=$ $-\left[\begin{array}[]{ccccc}-\epsilon_{\ell_{F_{\ell}N}\ell_{11}}^{\mathcal{C}}&\epsilon_{\ell_{11}\ell_{12}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}N}\ell_{12}}^{\mathcal{C}}&\cdots&\epsilon_{\ell_{11}\ell_{F_{\ell},N-1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}N}\ell_{F_{\ell},N-1}}^{\mathcal{C}}&\epsilon_{\ell_{11}\ell_{F_{\ell}N}}^{\mathcal{C}}\\ \epsilon_{\ell_{12}\ell_{11}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}N}\ell_{11}}^{\mathcal{C}}&-\epsilon_{\ell_{F_{\ell}N}\ell_{12}}^{\mathcal{C}}&\cdots&\epsilon_{\ell_{12}\ell_{F_{\ell},N-1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}N}\ell_{F_{\ell},N-1}}^{\mathcal{C}}&\epsilon_{\ell_{12}\ell_{F_{\ell}N}}^{\mathcal{C}}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ \epsilon_{\ell_{F_{\ell},N-1}\ell_{11}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}N}\ell_{11}}^{\mathcal{C}}&\epsilon_{\ell_{F_{\ell},N-1}\ell_{12}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}N}\ell_{12}}^{\mathcal{C}}&\cdots&-\epsilon_{\ell_{F_{\ell}N}\ell_{F_{\ell},N-1}}^{\mathcal{C}}&\epsilon_{\ell_{F_{\ell},N-1}\ell_{F_{\ell}N}}^{\mathcal{C}}\\ \Lambda_{\ell_{1}1}&\Lambda_{\ell_{1}2}&\cdots&\Lambda_{\ell_{F_{\ell}}N-1}&\Lambda_{\ell_{F_{\ell}}N}\end{array}\right]^{-1}$
$\times\left[\begin{array}[]{ccc}\Theta_{1\left(1,1\right)}^{w}-\Theta_{F_{\ell}\left(N,1\right)}^{w}&\cdots&\Theta_{1\left(1,N\right)}^{w}-\Theta_{F_{\ell}\left(N,N\right)}^{w}\\ \vdots&\vdots&\vdots\\ \Theta_{F_{\ell}\left(N-1,1\right)}^{w}-\Theta_{F_{\ell}\left(N,1\right)}^{w}&\cdots&\Theta_{F_{\ell}\left(N-1,N\right)}^{w}-\Theta_{F_{\ell}\left(N,N\right)}^{w}\\ \zeta_{1}&\cdots&\zeta_{N}\end{array}\right].$
The log-linearization of the average wages of labor $f$, $\overline{w}_{f}=\sum_{n=1}^{N}w_{fn}\ell_{fn}/\sum_{n=1}^{N}\ell_{fn}$, yields
$d\ln\overline{w}_{f}=\sum_{n=1}^{N}\left(\frac{\Lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)d\ln w_{fn}+\sum_{n=1}^{N}\left(\frac{\Lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}-\frac{\ell_{fn}}{\ell_{f}}\right)d\ln\ell_{fn}.$
The log change of the average wage of labor $f$ relative to labor $g$ can be written as
$d\ln\left(\frac{\overline{w}_{f}}{\overline{w}_{g}}\right)=\sum_{n=1}^{N}\left(\frac{\Lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)d\ln w_{fn}-\sum_{n=1}^{N}\left(\frac{\Lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)d\ln w_{gn}+\sum_{n=1}^{N}\left(\frac{\Lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}-\frac{\ell_{fn}}{\ell_{f}}\right)d\ln\ell_{fn}-\sum_{n=1}^{N}\left(\frac{\Lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\frac{\ell_{gn}}{\ell_{g}}\right)d\ln\ell_{gn}.$
Substituting equation (A33) into this and setting $d\ln\ell_{fn}=0$ yields equation (A3).
The log-linearization of the factor share, $\Lambda_{\ell_{f}}=\sum_{n=1}^{N}\Lambda_{\ell_{f}n}$, yields
$d\ln\Lambda_{\ell_{f}}=\sum_{n=1}^{N}\frac{\Lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}d\ln\Lambda_{\ell_{f}n}.$
Substituting equation (A31) into this yields
$d\ln\Lambda_{\ell_{f}}=\sum_{g=1}^{F_{\ell}}\sum_{m=1}^{N}\left(\sum_{n=1}^{N}\frac{\Lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\Pi_{UL,fg\left(n,m\right)}^{w}\right)d\ln\left(\frac{w_{gm}}{p}\right)\\ -\sum_{g=1}^{F_{k}}\sum_{m=1}^{N}\left(\sum_{n=1}^{N}\frac{\Lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\Pi_{UR,fg\left(n,m\right)}^{w}\right)d\ln q_{g}+\sum_{m=1}^{N}\left(\sum_{n=1}^{N}\frac{\Lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\Theta_{f\left(n,m\right)}^{w}\right)d\ln A_{m}.$
Substituting equation (A33) into this yields equation (A4), where
$\Pi_{UL,ff\left(n,n\right)}^{w}=$ $\sum_{l\neq n}^{N}\Lambda_{\ell_{f}l}\left(1-\epsilon_{\ell_{fl}\ell_{fn}}^{\mathcal{C}}\right)+\sum_{l=1}^{N}\sum_{h\neq f}^{F}\Lambda_{\ell_{h}l}\left(1-\epsilon_{\ell_{hl}\ell_{fn}}^{\mathcal{C}}\right),$
$\Pi_{UL,ff\left(n,m\right)}^{w}=$ $\Lambda_{\ell_{f}m}\left(\epsilon_{\ell_{fn}\ell_{fm}}^{\mathcal{C}}-1\right)+\sum_{l\neq n,m}^{N}\Lambda_{\ell_{f}l}\left(\epsilon_{\ell_{fn}\ell_{fm}}^{\mathcal{C}}-\epsilon_{\ell_{fl}\ell_{fm}}^{\mathcal{C}}\right)+\sum_{l=1}^{N}\sum_{h\neq f}^{F}\Lambda_{\ell_{h}l}\left(\epsilon_{\ell_{fn}\ell_{fm}}^{\mathcal{C}}-\epsilon_{\ell_{hl}\ell_{fm}}^{\mathcal{C}}\right),$
$\Pi_{UL,fg\left(n,m\right)}^{w}=$ $\Lambda_{\ell_{g}m}\left(\epsilon_{\ell_{fn}\ell_{gm}}^{\mathcal{C}}-1\right)+\sum_{l\neq n}^{N}\Lambda_{\ell_{f}l}\left(\epsilon_{\ell_{fn}\ell_{gm}}^{\mathcal{C}}-\epsilon_{\ell_{fl}\ell_{gm}}^{\mathcal{C}}\right)+\sum_{l\neq m}^{N}\Lambda_{\ell_{g}l}\left(\epsilon_{\ell_{fn}\ell_{gm}}^{\mathcal{C}}-\epsilon_{g_{fl}\ell_{gm}}^{\mathcal{C}}\right)$
$+\sum_{l=1}^{N}\sum_{h\neq f,g}^{F}\Lambda_{\ell_{h}l}\left(\epsilon_{\ell_{fn}\ell_{gm}}^{\mathcal{C}}-\epsilon_{\ell_{hl}\ell_{gm}}^{\mathcal{C}}\right),$
$\sum_{m=1}^{N}\Pi_{UR,fg\left(n,m\right)}^{w}=\sum_{l=1}^{N}\Lambda_{k_{g}l}\left(\epsilon_{\ell_{fn}k_{gl}}^{\mathcal{C}}-1\right)+\sum_{l\neq n}^{N}\Lambda_{\ell_{f}l}\left(\epsilon_{\ell_{fn}k_{gl}}^{\mathcal{C}}-\epsilon_{\ell_{fl}k_{gl}}^{\mathcal{C}}\right)+\sum_{l=1}^{N}\sum_{h\neq f,g}^{F}\Lambda_{\ell_{h}l}\left(\epsilon_{\ell_{fn}k_{gl}}^{\mathcal{C}}-\epsilon_{\ell_{hl}k_{gl}}^{\mathcal{C}}\right).$
Appendix C Additional corollaries
Corollary 5.
Suppose that production requires only two factors. The aggregate elasticity of substitution (12) is identical to the aggregate elasticity of substitution (13), which can be expressed as a weighted average of the elasticities of substitution in production and consumption:
$\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{F}}=\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}=\sum_{n=1}^{N}\zeta_{n}\frac{\lambda_{\ell_{f}n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{f}}\Lambda_{\ell_{g}}}\epsilon_{\ell_{f}\ell_{g}n}^{\mathcal{C}}+\left(1-\sum_{n=1}^{N}\zeta_{n}\frac{\lambda_{\ell_{f}n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{f}}\Lambda_{\ell_{g}}}\right)\frac{1}{1-\eta}.$
The first term represents the substitution between two factors within sectors, and the second term represents the reallocation between sectors with different factor intensities; the relative importance of the two effects depends on the degree of factor intensity (Oberfield and Raval, 2021).
Corollary 6.
Suppose that production requires only one type of capital, $k$, and one type of labor, $\ell$. Let $r$ and $w$ denote their respective prices. Changes in the labor share are given by
$\Delta\ln\Lambda_{\ell}=\frac{1-\epsilon}{\epsilon}\left(1-\Lambda_{\ell}\right)\Delta\ln\left(\frac{k}{\ell}\right)+\psi_{\left(\ell\right)}^{\ell}$
or
$\Delta\ln\Lambda_{\ell}=\left(1-\epsilon\right)\left(1-\Lambda_{\ell}\right)\Delta\ln\left(\frac{w}{r}\right)+\psi_{\left(\ell\right)}^{w},$
where $\epsilon$ is the aggregate elasticity of substitution between capital and labor.
In this case, the labor share can decline with a rise in the capital–labor ratio or the wage–rental price ratio if $\epsilon>1$. Karabarbounis and Neiman (2014) attribute the decline in the labor share to capital-embodied technological change on the basis of findings that the estimated elasticity of substitution exceeds one and that the rental price of capital fell in many countries. Elsby, Hobijn and Şahin (2013) attribute it to other factors on the basis of the observation that the timing of decline in the labor share did not necessarily align with the timing of capital deepening in the United States.
Corollary 7.
Suppose that production requires one type of capital ($k$) and two types of labor (skilled labor, $\ell_{h}$, and unskilled labor, $\ell_{u}$), and the aggregate elasticities of substitution satisfy $\epsilon_{\ell_{h}k}^{\mathcal{F}}=\epsilon_{\ell_{u}k}^{\mathcal{F}}=1$ and $\epsilon_{\ell_{h}\ell_{u}}^{\mathcal{F}}=\epsilon_{\ell_{u}\ell_{h}}^{\mathcal{F}}=\epsilon^{\mathcal{F}}$. Equation (18) becomes the Katz and Murphy (1992) model:
$\Delta\ln\left(\frac{w_{h}}{w_{u}}\right)=-\frac{1}{\epsilon^{\mathcal{F}}}\Delta\ln\left(\frac{\ell_{h}}{\ell_{u}}\right)+\left(\psi_{\left(h\right)}^{\ell}-\psi_{\left(u\right)}^{\ell}\right).$
This equation implies that the relative wages of skilled labor to unskilled labor can increase only with disembodied technological change biased towards skilled labor when the relative supply of skilled labor to unskilled labor increases over time.
Corollary 8.
Suppose that production requires two types of capital (capital equipment, $k_{e}$, and capital structures, $k_{s}$) and two types of labor (skilled labor, $\ell_{h}$, and unskilled labor, $\ell_{u}$), and the aggregate elasticities of substitution satisfy $\epsilon_{\ell_{h}k_{s}}^{\mathcal{F}}=\epsilon_{\ell_{u}k_{s}}^{\mathcal{F}}=1$ and $1/\epsilon_{\ell_{u}\ell_{h}}^{\mathcal{F}}-1/\epsilon_{\ell_{h}\ell_{u}}^{\mathcal{F}}=1/\epsilon_{\ell_{h}k_{e}}^{\mathcal{F}}-1/\epsilon_{\ell_{u}k_{e}}^{\mathcal{F}}$. Equation (18) becomes the Krusell, Ohanian, Ríos-Rull and Violante (2000) model:
$\Delta\ln\left(\frac{w_{h}}{w_{u}}\right)=\left(\frac{\epsilon_{\ell_{u}k_{e}}^{\mathcal{F}}-\epsilon_{\ell_{h}k_{e}}^{\mathcal{F}}}{\epsilon_{\ell_{u}k_{e}}^{\mathcal{F}}\epsilon_{\ell_{h}k_{e}}^{\mathcal{F}}}\right)\Delta\ln\left(\frac{k_{e}}{\ell_{h}}\right)-\frac{1}{\epsilon_{\ell_{h}\ell_{u}}^{\mathcal{F}}}\Delta\ln\left(\frac{\ell_{h}}{\ell_{u}}\right)+\left(\psi_{\left(h\right)}^{\ell}-\psi_{\left(u\right)}^{\ell}\right).$
This equation implies that the skill wage gap can increase with a rise in capital equipment if capital equipment is more complementary to skilled labor than to unskilled labor (i.e., $\epsilon_{\ell_{u}k_{e}}^{\mathcal{F}}>\epsilon_{\ell_{h}k_{e}}^{\mathcal{F}}$). This effect is sufficiently large to account for a rise in the skill premium in OECD countries when ICCT capital is distinguished from non-ICCT capital (Taniguchi and Yamada, 2022).
Appendix D Data Description
D.1 Wages and hours worked
We adjust for changes in the age and education composition of the labor force over time when we construct the data on wages and hours worked. Each type of labor, $\ell_{f}$ for $f\in\left\{mh,fh,mu,fu\right\}$, can be divided into three age groups: young (aged between 15 and 29 years), middle (aged between 30 and 49 years), and old (aged 50 years and older). In addition, unskilled labor consists of two education groups: medium-skilled (entered college or completed high-school education) and low-skilled (dropped out of high school or attended compulsory education only), while skilled labor consists of one education group: high-skilled (completed college). We assume that workers are perfect substitutes within each type of labor.
Suppose that there is no need to make an adjustment to wages and hours worked. Let checks denote unadjusted values. The wages for labor of type $f$ in sector $n$, country $j$, and year $t$ could be calculated as $\check{w}_{f,njt}=\sum_{a_{f}}\rho_{f,njt}^{a_{f}}\check{w}_{f,njt}^{a_{f}}$, where $\rho_{f,njt}^{a_{f}}$ is the share of total hours worked by group $a_{f}$ (i.e., $\rho_{f,njt}^{a_{f}}=\check{\ell}_{f,njt}^{a_{f}}/\sum_{a_{f}}\check{\ell}_{f,njt}^{a_{f}}$). The hours worked by labor of type $f$ in sector $n$, country $j$, and year $t$ could be calculated as $\check{\ell}_{f,njt}=\sum_{a_{f}}\check{\ell}_{f,njt}^{a_{f}}$. We adjust for changes in the age and education composition of the labor force by holding the share of each group constant when we calculate wages and by using the time-invariant efficiency units as weights when we calculate hours worked. Let $T_{j}$ denote the number of years observed for country $j$. The composition-adjusted wages for labor of type $f$ in sector $n$, country $j$, and year $t$ can be calculated as $w_{f,njt}=\sum_{a_{f}}\overline{\rho}_{f,nj}^{a_{f}}\check{w}_{f,njt}^{a_{f}}$, where $\overline{\rho}_{f,nj}^{a_{f}}$ is the sector- and country-specific mean of $\rho_{f,njt}^{a_{f}}$ (i.e., $\overline{\rho}_{f,nj}^{a_{f}}=\sum_{t=1}^{T_{j}}\rho_{f,njt}^{a_{f}}/T_{j}$). The composition-adjusted hours worked by labor of type $f$ in sector $n$, country $j$, and year $t$ can be calculated as $\ell_{f,njt}=\sum_{a_{f}}(\overline{w}_{f,nj}^{a_{f}}/\overline{w}_{f,nj}^{a_{f^{\prime}}})\check{\ell}_{f,njt}^{a_{f}}$, where the efficiency unit is measured by the sector- and country-specific mean of $\check{w}_{f,njt}^{a_{f}}$ (i.e., $\overline{w}_{f,nj}^{a_{f}}=\sum_{t=1}^{T_{j}}\check{w}_{f,njt}^{a_{f}}/T_{j}$) and normalized by $\overline{w}_{f,nj}^{a_{f^{\prime}}}$. We choose middle-aged high-skilled labor as the base group for skilled labor and middle-aged medium-skilled labor as the base group for unskilled labor. Our results do not depend on the choice of the base group.
D.2 Rental price of capital
The rental price of capital ($r_{ft}$) relative to the output price ($p_{t}$) is determined by the relative investment price ($p_{ft}/p_{t}$), depreciation rate ($\delta_{f}$), and interest rate ($\iota_{t}$) for $f\in\left\{i,o\right\}$. The investment price is calculated by dividing the nominal value by the real value of investment. The depreciation rate is the time average of those obtained from the capital law of motion.
We calculate the rental price of capital in two ways. First, we adopt the internal rate of return approach of O’Mahony and Timmer (2009), who calculate the rental price of capital as
$\frac{r_{ft}}{p_{t}}=\delta_{f}\left(\frac{p_{ft}}{p_{t}}\right)+\iota_{t}\left(\frac{p_{f,t-1}}{p_{t-1}}\right)-\left(\frac{p_{ft}}{p_{t}}-\frac{p_{f,t-1}}{p_{t-1}}\right),$ (A34)
where the interest rate is the internal rate of return:
$\iota_{t}=\frac{\sum_{f}\left(\left.r_{ft}\right/p_{t}\right)k_{ft}-\sum_{f}\delta_{f}\left(\left.p_{ft}\right/p_{t}\right)k_{ft}+\sum_{f}\left(\left.p_{ft}\right/p_{t}-\left.p_{f,t-1}\right/p_{t-1}\right)k_{ft}}{\sum_{f}\left(\left.p_{f,t-1}\right/p_{t-1}\right)k_{ft}}.$
Second, we adopt the external rate of return approach of Niebel and Saam (2016), who calculate the rental price of capital as
$\frac{r_{ft}}{p_{t}}=\delta_{f}\left(\frac{p_{ft}}{p_{t}}\right)+\iota_{t}\left(\frac{p_{f,t-1}}{p_{t-1}}\right)-\frac{1}{2}\left[\ln\left(\frac{p_{ft}}{p_{t}}\right)-\ln\left(\frac{p_{f,t-2}}{p_{t-2}}\right)\right]\left(\frac{p_{f,t-1}}{p_{t-1}}\right),$
and the interest rate as
$\iota_{t}=0.04+\frac{1}{5}\sum_{\tau=-2}^{2}\frac{cpi_{t-\tau}-cpi_{t-\tau-1}}{cpi_{t-\tau-1}},$
where cpicpi is the consumer price index. The data can be obtained from OECD.Stat.
D.3 Non-ICCT equipment
The trends in the rental prices of capital differ significantly between ICCT equipment and non-ICCT structures but do not differ significantly between non-ICCT equipment and non-ICCT structures (Figure A1a). The rental price of ICCT equipment fell dramatically, but that of non-ICCT equipment and structures remained almost unchanged. Meanwhile, the quantities of ICCT equipment, non-ICCT equipment, and non-ICCT structures increased. However, the rate of increase in ICCT equipment is far greater than that in non-ICCT equipment and structures (Figure A1b). The rate of increase in non-ICCT equipment is almost the same as that in non-ICCT structures.
Figure A1: ICCT equipment, non-ICCT equipment, and non-ICCT structures
(a) Investment prices
(b) Quantities
Notes: The investment prices of ICCT equipment, non-ICCT equipment, and non-ICCT structures are denoted by $p_{i}$, $p_{oe}$, and $p_{os}$, respectively, and the output price is denoted by $p$. The quantities of ICCT equipment, non-ICCT equipment, and non-ICCT structures are denoted by $k_{i}$, $k_{oe}$, and $k_{os}$, respectively. All the series are logarithmically transformed and normalized to zero in the year 1980.
Appendix E Specification Issues
This section is divided into four parts. First, we examine the consistency of the relative magnitude of substitution parameters across different specifications and discuss the empirical relevance of alternative specifications. Second, we discuss the error terms in the estimating equations. Third, we describe the weights on the substitution parameters in the aggregate elasticities of substitution. Finally, we consider the magnitude of the effects of technological change.
E.1 Substitution parameters
As a starting point, we consider the use of a one-level CES function:
$y=Ak_{o}^{\vartheta_{o}}\left[\left(1-\vartheta_{fh}-\vartheta_{mh}-\vartheta_{fu}-\vartheta_{mu}\right)k_{i}^{\varsigma}+\vartheta_{fh}\ell_{fh}^{\varsigma}+\vartheta_{mh}\ell_{mh}^{\varsigma}+\vartheta_{fu}\ell_{fu}^{\varsigma}+\vartheta_{mu}\ell_{mu}^{\varsigma}\right]^{\frac{1-\vartheta_{o}}{\varsigma}},$ (A35)
where the degree of substitution between ICCT capital and the four types of labor is governed by a single parameter $\varsigma$. We suppress the sector subscript $n$ for notational simplicity but allow all parameters to vary across sectors. Cost minimization implies that the marginal rate of technical substitution of labor for capital equals the relative price of labor to capital for all types of labor. This condition leads to the following first-difference equations:
$\Delta\ln\left(\frac{w_{fh}}{r_{i}}\right)$ $=-\left(1-\varsigma_{fh}\right)\Delta\ln\left(\frac{\ell_{fh}}{k_{i}}\right)+\Delta\nu_{fh},$ (A36)
$\Delta\ln\left(\frac{w_{mh}}{r_{i}}\right)$ $=-\left(1-\varsigma_{mh}\right)\Delta\ln\left(\frac{\ell_{mh}}{k_{i}}\right)+\Delta\nu_{mh},$ (A37)
$\Delta\ln\left(\frac{w_{fu}}{r_{i}}\right)$ $=-\left(1-\varsigma_{fu}\right)\Delta\ln\left(\frac{\ell_{fu}}{k_{i}}\right)+\Delta\nu_{fu},$ (A38)
$\Delta\ln\left(\frac{w_{mu}}{r_{i}}\right)$ $=-\left(1-\varsigma_{mu}\right)\Delta\ln\left(\frac{\ell_{mu}}{k_{i}}\right)+\Delta\nu_{mu}.$ (A39)
The substitution parameter $\varsigma$ can be estimated using any one of equations (A36)–(A39). If the production function (A35) is correctly specified, the coefficients should be the same for all equations. Table A1 presents the results of parameter estimates and hypothesis testing in different specifications. The substitution parameters are estimated jointly by GMM using instruments described in Section 5.2. Panel A of Table A1 indicates $\varsigma_{fh}<\varsigma_{mh}<\varsigma_{fu}<\varsigma_{mu}$ for each sector. The null hypothesis of $\varsigma_{fh}=\varsigma_{mh}=\varsigma_{fu}=\varsigma_{mu}$ is strongly rejected for manufacturing, utilities, and services sectors. The specification (A35) is not valid for the three sectors, though it is for construction sector.
Given the results above, we extend the production function (A35) by incorporating a two-level CES function for the three sectors:
$y=Ak_{o}^{\alpha}\left[\left(1-\vartheta_{mh}-\vartheta_{fu}-\vartheta_{mu}\right)D^{\varsigma^{\prime}}+\vartheta_{mh}\ell_{mh}^{\varsigma^{\prime}}+\vartheta_{fu}\ell_{fu}^{\varsigma^{\prime}}+\vartheta_{mu}\ell_{mu}^{\varsigma^{\prime}}\right]^{\frac{1-\alpha}{\varsigma^{\prime}}},$ (A40)
where the degree of substitution between $k_{i}$ and the three types of labor ($\ell_{mh}$, $\ell_{fu}$, and $\ell_{mu}$) is governed by a parameter $\varsigma^{\prime}$. The degree of substitution between $k_{i}$ and $\ell_{fh}$ is governed by the parameter $\sigma_{fh}$, as in equation (27). Equations (A36)–(A39) are modified as follows:
$\Delta\ln\left(\frac{w_{fh}}{r_{i}}\right)$ $=-\left(1-\sigma_{fh}\right)\Delta\ln\left(\frac{\ell_{fh}}{k_{i}}\right)+\Delta\nu_{fh},$ (A41)
$\Delta\ln\left(\frac{w_{mh}}{w_{fh}}\right)$ $=-\left(1-\varsigma_{mh}^{\prime}\right)\Delta\ln\left(\frac{\ell_{mh}}{D}\right)+\left(1-\sigma_{fh}\right)\Delta\ln\left(\frac{\ell_{fh}}{D}\right)+\Delta\nu_{mh}^{\prime},$ (A42)
$\Delta\ln\left(\frac{w_{fu}}{w_{fh}}\right)$ $=-\left(1-\varsigma_{fu}^{\prime}\right)\Delta\ln\left(\frac{\ell_{fu}}{D}\right)+\left(1-\sigma_{fh}\right)\Delta\ln\left(\frac{\ell_{fh}}{D}\right)+\Delta\nu_{fu}^{\prime},$ (A43)
$\Delta\ln\left(\frac{w_{mu}}{w_{fh}}\right)$ $=-\left(1-\varsigma_{mu}^{\prime}\right)\Delta\ln\left(\frac{\ell_{mu}}{D}\right)+\left(1-\sigma_{fh}\right)\Delta\ln\left(\frac{\ell_{fh}}{D}\right)+\Delta\nu_{mu}^{\prime}.$ (A44)
The substitution parameter $\varsigma^{\prime}$ can be estimated using any one of equations (A42)–(A44). If the production function (A40) is correctly specified, the coefficients on the first terms of the last three equations should be the same. Panel B also indicates $\sigma_{fh}<\varsigma_{mh}^{\prime}<\varsigma_{fu}^{\prime}<\varsigma_{mu}^{\prime}$ for each sector. The null hypothesis of $\varsigma_{mh}^{\prime}=\varsigma_{fu}^{\prime}=\varsigma_{mu}^{\prime}$ is strongly rejected for manufacturing, utilities, and services sectors. Specification (A40) is not valid for the three sectors.
Given the results above, we further extend the production function (A40) by incorporating a three-level CES function for the three sectors:
$y=Ak_{o}^{\alpha}\left[\left(1-\vartheta_{fu}-\vartheta_{mu}\right)C^{\varsigma^{\prime\prime}}+\vartheta_{fu}\ell_{fu}^{\varsigma^{\prime\prime}}+\vartheta_{mu}\ell_{mu}^{\varsigma^{\prime\prime}}\right]^{\frac{1-\alpha}{\varsigma^{\prime\prime}}},$ (A45)
where the degree of substitution between $k_{i}$ and the two types of labor ($\ell_{fu}$ and $\ell_{mu}$) is governed by a parameter $\varsigma^{\prime\prime}$. The degree of substitution between $k_{i}$ and $\ell_{mh}$ is governed by the parameter $\sigma_{mh}$ , as in equation (26). Equations (A41)–(A44) are modified as follows:
$\Delta\ln\left(\frac{w_{fh}}{r_{i}}\right)$ $=-\left(1-\sigma_{fh}\right)\Delta\ln\left(\frac{\ell_{fh}}{k_{i}}\right)+\Delta\nu_{fh},$ (A46)
$\Delta\ln\left(\frac{w_{mh}}{w_{fh}}\right)$ $=-\left(1-\sigma_{mh}\right)\Delta\ln\left(\frac{\ell_{mh}}{D}\right)+\left(1-\sigma_{fh}\right)\Delta\ln\left(\frac{\ell_{fh}}{D}\right)+\Delta\nu_{mh}^{\prime},$ (A47)
$\Delta\ln\left(\frac{w_{fu}}{w_{fh}}\right)$ $=-\left(1-\varsigma_{fu}^{\prime\prime}\right)\Delta\ln\left(\frac{\ell_{fu}}{C}\right)+\left(1-\sigma_{fh}\right)\Delta\ln\left(\frac{\ell_{fh}}{D}\right)-\left(1-\sigma_{mh}\right)\Delta\ln\left(\frac{C}{D}\right)+\Delta\nu_{fu}^{\prime\prime},$ (A48)
$\Delta\ln\left(\frac{w_{mu}}{w_{fh}}\right)$ $=-\left(1-\varsigma_{mu}^{\prime\prime}\right)\Delta\ln\left(\frac{\ell_{mu}}{C}\right)+\left(1-\sigma_{fh}\right)\Delta\ln\left(\frac{\ell_{fh}}{D}\right)-\left(1-\sigma_{mh}\right)\Delta\ln\left(\frac{C}{D}\right)+\Delta\nu_{mu}^{\prime\prime}.$ (A49)
The substitution parameter $\varsigma^{\prime\prime}$ can be estimated using either equation (A48) or (A49). If the production function (A45) is correctly specified, the coefficients on the first terms of the last two equations should be the same. Panel C again indicates $\sigma_{fh}<\sigma_{mh}<\varsigma_{fu}^{\prime\prime}<\varsigma_{mu}^{\prime\prime}$ for each sector. The null hypothesis of $\varsigma_{fu}^{\prime\prime}=\varsigma_{mu}^{\prime\prime}$ is rejected for manufacturing, utilities, and services sectors. Specification (A45) is not valid for the three sectors.
Finally, we extend the production function (A45) by incorporating a four-level CES function, as in equation (24). Panel D indicates $\sigma_{fh}<\sigma_{mh}<\sigma_{fu}<\sigma_{mu}$ for each sector. The null hypotheses of $\sigma_{fh}=\sigma_{mh}$, $\sigma_{mh}=\sigma_{fu}$, and $\sigma_{fu}=\sigma_{mu}$ are strongly rejected for manufacturing, utilities, and services sectors. The results are consistent with specification (24). The relative magnitude of the estimated substitution parameters is consistent across specifications.
Table A1: Production function estimates in alternative specifications
Parameter estimates Hypothesis testing
A. One-level CES A. One-level CES
$\varsigma_{fh}$ $\varsigma_{mh}$ $\varsigma_{fu}$ $\varsigma_{mu}$ $\varsigma_{fh}=\varsigma_{mh}=\varsigma_{fu}=\varsigma_{mu}$
Manufacturing –0.689 0.154 0.349 0.380 70.031
(0.319) (0.057) (0.055) (0.048) [0.000]
Utilities –1.120 0.081 0.221 0.339 31.199
(0.391) (0.130) (0.091) (0.078) [0.000]
Construction 0.096 0.453 0.450 0.507 3.229
(0.320) (0.126) (0.117) (0.097) [0.358]
Services –0.402 0.098 0.259 0.356 66.105
(0.128) (0.063) (0.051) (0.054) [0.000]
B. Two-level CES B. Two-level CES
$\sigma_{fh}$ $\varsigma_{mh}^{\prime}$ $\varsigma_{fu}^{\prime}$ $\varsigma_{mu}^{\prime}$ $\varsigma_{mh}^{\prime}=\varsigma_{fu}^{\prime}=\varsigma_{mu}^{\prime}$
Manufacturing –0.689 0.287 0.467 0.506 105.444
(0.319) (0.046) (0.047) (0.046) [0.000]
Utilities –1.120 0.189 0.303 0.419 34.979
(0.391) (0.105) (0.071) (0.064) [0.000]
Services –0.402 0.343 0.499 0.607 56.943
(0.128) (0.070) (0.047) (0.053) [0.000]
C. Three-level CES C. Three-level CES
$\sigma_{fh}$ $\sigma_{mh}$ $\varsigma_{fu}^{\prime\prime}$ $\varsigma_{mu}^{\prime\prime}$ $\varsigma_{fu}^{\prime\prime}=\varsigma_{mu}^{\prime\prime}$
Manufacturing –0.689 0.287 0.597 0.673 8.765
(0.319) (0.046) (0.047) (0.051) [0.003]
Utilities –1.120 0.189 0.384 0.560 38.617
(0.391) (0.105) (0.067) (0.062) [0.000]
Services –0.402 0.343 0.601 0.755 40.649
(0.128) (0.070) (0.038) (0.049) [0.000]
D. Four-level CES D. Four-level CES
$\sigma_{fh}$ $\sigma_{mh}$ $\sigma_{fu}$ $\sigma_{mu}$ $\sigma_{fh}=\sigma_{mh}$ $\sigma_{mh}=\sigma_{fu}$ $\sigma_{fu}=\sigma_{mu}$
Manufacturing –0.689 0.288 0.600 0.776 10.540 108.465 14.488
(0.319) (0.046) (0.043) (0.054) [0.001] [0.000] [0.000]
Utilities –0.962 0.184 0.404 0.631 15.300 6.140 23.543
(0.362) (0.104) (0.063) (0.058) [0.000] [0.013] [0.000]
Services –0.413 0.340 0.598 0.861 25.515 22.783 61.220
(0.134) (0.067) (0.034) (0.048) [0.000] [0.000] [0.000]
Notes: The first to fourth columns report the estimates of the substitution parameters with standard errors clustered at the country level in parentheses. The fifth to seventh columns report Wald statistics under the stated hypotheses with p-values in square brackets.
E.2 Nesting structure
Our specification (24) is a natural extension of the one used in Fallon and Layard (1975) and Krusell et al. (2000). They estimate the production function of the form:
$y=Ak_{o}^{\vartheta_{o}}\left[\left(1-\vartheta_{u}\right)\left[\left(1-\vartheta_{h}\right)k_{i}^{\varsigma_{h}}+\vartheta_{h}\ell_{h}^{\varsigma_{h}}\right]^{\frac{\varsigma_{u}}{\varsigma_{h}}}+\vartheta_{u}\ell_{u}^{\varsigma_{u}}\right]^{\frac{1-\vartheta_{o}}{\varsigma_{u}}},$ (A50)
where the two substitution parameters ($\varsigma_{h}$, $\varsigma_{u}$) are less than one, and the three share parameters ($\vartheta_{h}$, $\vartheta_{u}$, and $\vartheta_{0}$) lie between zero and one. There are three remarks regarding the specification of the production function (A50). First, new equipment is located inside the CES function to allow for the possibility that capital-embodied technological change may influence the skill wage gap. Second, new equipment is placed in the lowest nest to allow for the possibility that it may be more complementary to skilled labor than to unskilled labor. Finally, the production function is confirmed to be concave (i.e., the estimates of the substitution parameters are less than one); however, not so if unskilled labor is placed in a nest lower than that of skilled labor. This condition ensures that the relative price of skills decreases with a rise in the relative supply of skills.
The production function (A50) proves to be useful in accounting for changes in the skill premium in the United States and other advanced countries (Krusell et al., 2000; Taniguchi and Yamada, 2022). However, it is limited in that male and female labor are assumed to be perfect substitutes. One simple extension is to incorporate the following CES function into equation (A50):
$\ell_{h}$ $=\left[\left(1-\vartheta_{fh}\right)\ell_{mh}^{\varsigma_{fh}}+\vartheta_{fh}\ell_{fh}^{\varsigma_{fh}}\right]^{\frac{1}{\varsigma_{fh}}},$
$\ell_{u}$ $=\left[\left(1-\vartheta_{fu}\right)\ell_{mu}^{\varsigma_{fu}}+\vartheta_{fu}\ell_{fu}^{\varsigma_{fu}}\right]^{\frac{1}{\varsigma_{fu}}},$
where the two substitution parameters ($\varsigma_{fh}$, $\varsigma_{fu}$) are less than one, and the two share parameters ($\vartheta_{fh}$, $\vartheta_{fu}$) lie between zero and one. This specification relaxes the assumption that male and female labor are perfect substitutes yet still imposes the assumption that male and female labor are equally substitutable with ICCT capital for each skill type. This assumption implies that technological change embodied in new equipment has no effect on the gender wage gap. Both the assumption and the implication are inconsistent with the data, as shown above. Therefore, there are only two alternatives. One is our preferred specification (24). The other is the following specification:
$y=Ak_{o}^{\vartheta_{o}}\left[\left(1-\vartheta_{fu}\right)B^{\prime\varsigma_{fu}}+\vartheta_{fu}\ell_{fu}^{\varsigma_{fu}}\right]^{\frac{1-\vartheta_{o}}{\varsigma_{fu}}},$
where
$B^{\prime}$ $=\left[\left(1-\vartheta_{mu}\right)C_{n}^{\prime\varsigma_{mu}}+\vartheta_{mu}\ell_{mu}^{\varsigma_{mu}}\right]^{\frac{1}{\varsigma_{mu}}},$
$C^{\prime}$ $=\left[\left(1-\vartheta_{fh}\right)D^{\prime\varsigma_{fh}}+\vartheta_{fh}\ell_{fh}^{\varsigma_{fh}}\right]^{\frac{1}{\varsigma_{fh}}},$
$D^{\prime}$ $=\left[\left(1-\vartheta_{mh}\right)k_{i}^{\varsigma_{mh}}+\vartheta_{mh}\ell_{mh}^{\varsigma_{mh}}\right]^{\frac{1}{\varsigma_{mh}}}.$
In this specification, male labor is located in a nest lower than that of female labor for each skill type. The four substitution parameters ($\varsigma_{mh}$, $\varsigma_{fh}$, $\varsigma_{mu}$, $\varsigma_{fu}$) are less than one, and the five share parameters ($\vartheta_{mh}$, $\vartheta_{fh}$, $\vartheta_{mu}$, $\vartheta_{fu}$, $\alpha$) lie between zero and one. The problem with this specification is that the production function is not concave, as in the specification where unskilled labor is located in a nest lower than that of skilled labor. Specifically, the estimate of $\varsigma_{fh}$ exceeds one in all sectors regardless of the use of instruments. This result contradicts the economic principle that the relative wages of female skilled labor should decrease with a rise in the relative supply of female skilled labor. To summarize, there are three ways to extend the specification of the production function (A50). Female labor may be located in a nest lower or higher than or in the same nest as that of male labor. Our specification is consistent with the data, but the other two are not.
E.3 Error terms
Error terms are added to equations (28)–(32) for a couple of reasons. First, the gender wage gap and the skill wage gap may be measured with errors. Second, the CES aggregates may be approximated with errors. Finally, the share parameters may change over time, possibly due to disembodied factor-biased technological change. If the share parameters change over time, the first-order Taylor expansion of equations (25)–(27) yields
$\Delta\ln D_{n}$ $\simeq\Delta\ln\widetilde{D}_{n}+\Delta v_{D,n},$
$\Delta\ln C_{n}$ $\simeq\Delta\ln\widetilde{C}_{n}+\Delta v_{C,n},$
$\Delta\ln B_{n}$ $\simeq\Delta\ln\widetilde{B}_{n}+\Delta v_{B,n},$
where
$\Delta v_{D,n}=$ $\frac{1}{\sigma_{fh,n}}\left[\left(1-\varphi_{fh,n}\right)\Delta\ln\left(1-\theta_{fh,n}\right)+\varphi_{fh,n}\Delta\ln\theta_{fh,n}\right],$
$\Delta v_{C,n}=$ $\left(1-\varphi_{mh,n}\right)\Delta v_{D,n}+\frac{1}{\sigma_{mh,n}}\left[\left(1-\varphi_{mh,n}\right)\Delta\ln\left(1-\theta_{mh,n}\right)+\varphi_{mh,n}\Delta\ln\theta_{mh,n}\right],$
$\Delta v_{B,n}=$ $\left(1-\varphi_{fu,n}\right)\Delta v_{C,n}+\frac{1}{\sigma_{fu,n}}\left[\left(1-\varphi_{fu,n}\right)\Delta\ln\left(1-\theta_{fu,n}\right)+\varphi_{fu,n}\Delta\ln\theta_{fu,n}\right].$
Substituting these equations into the marginal rate of technical substitution equations yields a system of equations (28)–(32), where the error terms are
$\Delta v_{fh,n}=$ $\Delta\ln\theta_{fh,n}-\Delta\ln\left(1-\theta_{fh,n}\right).$
$\Delta v_{mh,n}=$ $-\left(\sigma_{mh,n}-\sigma_{fh,n}\right)\Delta v_{D,n}+\Delta\ln\theta_{mh,n}-\Delta\ln\left(1-\theta_{mh,n}\right)-\Delta\ln\theta_{fh,n},$
$\Delta v_{fu,n}=$ $\left(\sigma_{fu,n}-\sigma_{mh,n}\right)\Delta v_{C,n}+\left(\sigma_{mh,n}-\sigma_{fh,n}\right)\Delta v_{D,n}$
$+\Delta\ln\left(1-\theta_{fu,n}\right)-\Delta\ln\theta_{fu,n}+\Delta\ln\left(1-\theta_{mh,n}\right)+\Delta\ln\theta_{fh,n},$
$\Delta v_{mu1,n}=$ $-\left(\sigma_{mu,n}-\sigma_{fu,n}\right)\Delta v_{B,n}+\Delta\ln\theta_{mu,n}-\Delta\ln\left(1-\theta_{mu,n}\right)-\Delta\ln\theta_{fu,n},$
$\Delta v_{mu2,n}=$ $\left(\sigma_{mu,n}-\sigma_{fu,n}\right)\Delta v_{B,n}+\left(\sigma_{fu,n}-\sigma_{mh,n}\right)\Delta v_{C,n}$
$+\Delta\ln\left(1-\theta_{mu,n}\right)-\Delta\ln\theta_{mu,n}+\Delta\ln\left(1-\theta_{fu,n}\right)+\Delta\ln\theta_{mh,n},$
E.4 Weights in the aggregate elasticities
When the sectoral production function is of the nested CES form (24), the aggregate elasticity of substitution (13) can be rewritten as equation (33), where each substitution parameter is summed across sectors with the following weights:
$\pi_{\ell_{f}\ell_{g},n}^{fh}$ $=\begin{cases}\left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(1-\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{1}}\lambda_{\ell_{h}n}}\right)+\left(\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{1}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f},\ell_{g}\in\mathcal{L}_{1}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(1-\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{1}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\notin\mathcal{L}_{1},\ell_{g}\in\mathcal{L}_{1}\\ 0&\text{if }\ell_{g}\notin\mathcal{L}_{1},\end{cases}$
$\pi_{\ell_{f}\ell_{g},n}^{mh}$ $=\begin{cases}\left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{1}}\lambda_{\ell_{h}n}}\right)\left(\frac{\lambda_{\ell_{mh}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)+\left(\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{1}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}=\ell_{mh},\ell_{g}\in\mathcal{L}_{1}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{1}}\lambda_{\ell_{h}n}}\right)\left(\frac{\lambda_{\ell_{mh}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f},\ell_{g}\in\mathcal{L}_{1}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{1}}\lambda_{\ell_{h}n}}\right)\left(\frac{\lambda_{\ell_{mh}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\notin\mathcal{L}_{2},\ell_{g}\in\mathcal{L}_{1}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(1-\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)+\left(\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\in\mathcal{L}_{1},\ell_{g}=\ell_{mh}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(1-\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\notin\mathcal{L}_{1},\ell_{g}=\ell_{mh}\\ 0&\text{if }\ell_{g}\notin\mathcal{L}_{2},\end{cases}$
$\pi_{\ell_{f}\ell_{g},n}^{fu}$ $=\begin{cases}\left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)\left(\frac{\lambda_{\ell_{fu}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)+\left(\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}=\ell_{fu},\ell_{g}\in\mathcal{L}_{2}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)\left(\frac{\lambda_{\ell_{fu}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f},\ell_{g}\in\mathcal{L}_{2}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)\left(\frac{\lambda_{\ell_{fu}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\notin\mathcal{L}_{3},\ell_{g}\in\mathcal{L}_{2}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(1-\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)+\left(\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\in\mathcal{L}_{2},\ell_{g}=\ell_{fu}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(1-\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\notin\mathcal{L}_{2},\ell_{g}=\ell_{fu}\\ 0&\text{if }\ell_{g}\notin\mathcal{L}_{3},\end{cases}$
$\pi_{\ell_{f}\ell_{g},n}^{mu}$ $=\begin{cases}\left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)\left(\frac{\lambda_{\ell_{mu}n}}{\sum_{\ell_{h}\in\mathcal{L}_{4}}\lambda_{\ell_{h}n}}\right)+\left(\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}=\ell_{mu},\ell_{g}\in\mathcal{L}_{3}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)\left(\frac{\lambda_{\ell_{mu}n}}{\sum_{\ell_{h}\in\mathcal{L}_{4}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f},\ell_{g}\in\mathcal{L}_{3}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)\left(\frac{\lambda_{\ell_{mu}n}}{\sum_{\ell_{h}\in\mathcal{L}_{4}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\notin\mathcal{L}_{4},\ell_{g}\in\mathcal{L}_{3}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(1-\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{4}}\lambda_{\ell_{h}n}}\right)+\left(\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{4}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\in\mathcal{L}_{3},\ell_{g}=\ell_{mu}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(1-\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{4}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\notin\mathcal{L}_{3},\ell_{g}=\ell_{mu}\\ 0&\text{if }\ell_{g}\notin\mathcal{L}_{4},\end{cases}$
$\pi_{\ell_{f}\ell_{g},n}^{o}$ $=\begin{cases}\left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{4}}\lambda_{\ell_{h}n}}\right)\lambda_{k_{o}n}&\text{if }\ell_{f},\ell_{g}\in\mathcal{L}_{4}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{4}}\lambda_{\ell_{h}n}}\right)-\pi_{\ell_{f}\ell_{g},n}^{c}&\text{if }\ell_{f}=k_{o}\\ \frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\pi_{\ell_{f}\ell_{g},n}^{c}&\text{if }\ell_{g}=k_{o},\end{cases}$
and
$\pi_{\ell_{f}\ell_{g},n}^{c}$ $=\begin{cases}\left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{4}}\lambda_{\ell_{h}n}}\right)\left(1-\lambda_{k_{o}n}\right)&\text{if }\ell_{g}\in\mathcal{L}_{4}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\lambda_{\ell_{g}n}&\text{if }\ell_{g}=k_{o},\end{cases}$
where the sets of factors are defined as $\mathcal{L}_{1}=\{k_{i},\ell_{fh}\}$, $\mathcal{L}_{2}=\{k_{i},\ell_{fh},\ell_{mh}\}$, $\mathcal{L}_{3}=\{k_{i},\ell_{fh},\ell_{mh},\ell_{fu}\}$, and $\mathcal{L}_{4}=\{k_{i},\ell_{fh},\ell_{mh},\ell_{fu},\ell_{mu}\}$.
E.5 Magnitude of the effects of technological change
We show that the effects of technological progress embodied in ICCT capital tend to increase as the differences in the substitution parameters in the production function increase and the income share of ICCT capital relative to the income shares of other factors increases. For this purpose, we assume here that the aggregate production function is of the nested CES form (24) (i.e., $N=1$) and that the capital–skill–gender complementarity hypothesis holds (i.e., $\sigma_{fh}<\sigma_{mh}<\sigma_{fu}<\sigma_{mu}$). The effects of technological change embodied in ICCT capital on the relative wages and income shares of the four types of labor can be obtained by setting $d\ln q_{o}=d\ln A_{n}=0$ in equations (A1) and (A2), respectively.
We first consider the effects on the four income shares. The effect on the income share of female skilled labor is
$\frac{\partial\ln\Lambda_{\ell_{fh}}}{\partial\ln q_{i}}=\left(1-\epsilon_{k_{i}\ell_{fh}}^{\mathcal{C}}\right)\left(\frac{\Lambda_{k_{i}}}{\Lambda_{\ell_{fh}}}\right)\left(\frac{H_{fh}}{H}\right)+\left(\epsilon_{\ell_{mh}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{fh}k_{i}}^{\mathcal{C}}\right)\left(\frac{H_{mh}}{H}\right)\\ +\left(\epsilon_{\ell_{fu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{fh}k_{i}}^{\mathcal{C}}\right)\left(\frac{H_{fu}}{H}\right)+\left(\epsilon_{\ell_{mu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{fh}k_{i}}^{\mathcal{C}}\right)\left(\frac{H_{mu}}{H}\right),$ (A51)
where
$H=H_{fh}+H_{mh}+H_{fu}+H_{mu},$
$H_{fh}=$ $\left(\epsilon_{k_{i}\ell_{mu}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{k_{i}\ell_{fu}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{k_{i}\ell_{mh}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\Lambda_{\ell_{fh}}+\left(\epsilon_{\ell_{mu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{fu}k_{i}}^{\mathcal{C}}\right)\left(\epsilon_{\ell_{fu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{mh}k_{i}}^{\mathcal{C}}\right)\left(\epsilon_{\ell_{mh}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{fh}k_{i}}^{\mathcal{C}}\right)\left(\frac{\Lambda_{\ell_{fh}}\Lambda_{\ell_{mh}}\Lambda_{\ell_{fu}}\Lambda_{\ell_{mu}}}{\Lambda_{k_{i}}\Lambda_{k_{i}}\Lambda_{k_{i}}}\right)$
$+\left(\epsilon_{k_{i}\ell_{fu}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{k_{i}\ell_{mu}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{\ell_{mh}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{fh}k_{i}}^{\mathcal{C}}\right)\left(\frac{\Lambda_{\ell_{fh}}\Lambda_{\ell_{mh}}}{\Lambda_{k_{i}}}\right)+\left(\epsilon_{k_{i}\ell_{mu}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{\ell_{fu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{mh}k_{i}}^{\mathcal{C}}\right)\left(\epsilon_{\ell_{mh}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{fh}k_{i}}^{\mathcal{C}}\right)\left(\frac{\Lambda_{\ell_{fh}}\Lambda_{\ell_{mh}}\Lambda_{\ell_{fu}}}{\Lambda_{k_{i}}\Lambda_{k_{i}}}\right)$
$+\left(\epsilon_{k_{i}\ell_{mh}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{k_{i}\ell_{mu}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{\ell_{fu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{fh}k_{i}}^{\mathcal{C}}\right)\left(\frac{\Lambda_{\ell_{fh}}\Lambda_{\ell_{fu}}}{\Lambda_{k_{i}}}\right)+\left(\epsilon_{k_{i}\ell_{fu}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{\ell_{mu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{mh}k_{i}}^{\mathcal{C}}\right)\left(\epsilon_{\ell_{mh}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{fh}k_{i}}^{\mathcal{C}}\right)\left(\frac{\Lambda_{\ell_{fh}}\Lambda_{\ell_{mh}}\Lambda_{\ell_{mu}}}{\Lambda_{k_{i}}\Lambda_{k_{i}}}\right)$
$+\left(\epsilon_{k_{i}\ell_{mh}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{k_{i}\ell_{fu}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{\ell_{mu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{fh}k_{i}}^{\mathcal{C}}\right)\left(\frac{\Lambda_{\ell_{fh}}\Lambda_{\ell_{mu}}}{\Lambda_{k_{i}}}\right)+\left(\epsilon_{k_{i}\ell_{mh}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{\ell_{mu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{fu}k_{i}}^{\mathcal{C}}\right)\left(\epsilon_{\ell_{fu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{fh}k_{i}}^{\mathcal{C}}\right)\left(\frac{\Lambda_{\ell_{fh}}\Lambda_{\ell_{fu}}\Lambda_{\ell_{mu}}}{\Lambda_{k_{i}}\Lambda_{k_{i}}}\right),$
$H_{mh}=$ $\left(\epsilon_{k_{i}\ell_{fh}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{k_{i}\ell_{fu}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{k_{i}\ell_{mu}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\Lambda_{\ell_{mh}}+\left(\epsilon_{k_{i}\ell_{fh}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{\ell_{mu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{fu}k_{i}}^{\mathcal{C}}\right)\left(\epsilon_{\ell_{fu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{mh}k_{i}}^{\mathcal{C}}\right)\left(\frac{\Lambda_{\ell_{mh}}\Lambda_{\ell_{fu}}\Lambda_{\ell_{mu}}}{\Lambda_{k_{i}}\Lambda_{k_{i}}}\right)$
$+\left(\epsilon_{k_{i}\ell_{fh}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{k_{i}\ell_{mu}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{\ell_{fu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{mh}k_{i}}^{\mathcal{C}}\right)\left(\frac{\Lambda_{\ell_{mh}}\Lambda_{\ell_{fu}}}{\Lambda_{k_{i}}}\right)+\left(\epsilon_{k_{i}\ell_{fh}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{k_{i}\ell_{fu}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{\ell_{mu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{mh}k_{i}}^{\mathcal{C}}\right)\left(\frac{\Lambda_{\ell_{mh}}\Lambda_{\ell_{mu}}}{\Lambda_{k_{i}}}\right),$
$H_{fu}=\left(\epsilon_{k_{i}\ell_{mu}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{k_{i}\ell_{mh}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{k_{i}\ell_{fh}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\Lambda_{\ell_{fu}}+\left(\epsilon_{k_{i}\ell_{mh}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{k_{i}\ell_{fh}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{\ell_{mu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{fu}k_{i}}^{\mathcal{C}}\right)\left(\frac{\Lambda_{\ell_{fu}}\Lambda_{\ell_{mu}}}{\Lambda_{k_{i}}}\right),$
$H_{mu}=\left(\epsilon_{k_{i}\ell_{fu}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{k_{i}\ell_{mh}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{k_{i}\ell_{fh}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\Lambda_{\ell_{mu}}.$
The effect on the income share of male skilled labor is
$\frac{\partial\ln\Lambda_{\ell_{mh}}}{\partial\ln q_{i}}=\left(1-\epsilon_{k_{i}\ell_{mh}}^{\mathcal{C}}\right)\left(\frac{\Lambda_{k_{i}}}{\Lambda_{\ell_{mh}}}\right)\left(\frac{H_{mh}}{H}\right)+\left(\epsilon_{\ell_{fu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{mh}k_{i}}^{\mathcal{C}}\right)\left(\frac{H_{fu}}{H}\right)+\left(\epsilon_{\ell_{mu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{mh}k_{i}}^{\mathcal{C}}\right)\left(\frac{H_{mu}}{H}\right).$ (A52)
The effect on the income share of female unskilled labor is
$\frac{\partial\ln\Lambda_{\ell_{fu}}}{\partial\ln q_{i}}=\left(1-\epsilon_{k_{i}\ell_{fu}}^{\mathcal{C}}\right)\left(\frac{\Lambda_{k_{i}}}{\Lambda_{\ell_{fu}}}\right)\left(\frac{H_{fu}}{H}\right)+\left(\epsilon_{\ell_{mu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{fu}k_{i}}^{\mathcal{C}}\right)\left(\frac{H_{mu}}{H}\right).$ (A53)
The effect on the income share of male unskilled labor is
$\frac{\partial\ln\Lambda_{\ell_{mu}}}{\partial\ln q_{i}}=\left(1-\epsilon_{k_{i}\ell_{mu}}^{\mathcal{C}}\right)\left(\frac{\Lambda_{k_{i}}}{\Lambda_{\ell_{mu}}}\right)\left(\frac{H_{mu}}{H}\right).$ (A54)
The magnitude of the first terms in equations (A51)–(A54) depends on the substitution parameters in the production function and the income share of ICCT capital relative to the income shares of other factors.
$\left(1-\epsilon_{k_{i}\ell_{f}}^{\mathcal{C}}\right)\left(\frac{\Lambda_{k_{i}}}{\Lambda_{\ell_{f}}}\right)=\left(1-\frac{1}{1-\sigma_{f}}\right)\left(\frac{\Lambda_{k_{i}}}{\Lambda_{\ell_{f}}}\right)\qquad\text{for }f\in\left\{mh,fh,mu,fu\right\}.$
The magnitude of the second to the last terms in equations (A51)–(A54) depends on the differences between the aggregate elasticities, which in turn depend on the differences between the substitution parameters in the production function and the income share of ICCT capital relative to the income shares of other factors.
$\epsilon_{\ell_{mh}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{fh}k_{i}}^{\mathcal{C}}=\left(\frac{1}{1-\sigma_{mh}}-\frac{1}{1-\sigma_{fh}}\right)\left(\frac{\Lambda_{k_{i}}}{\Lambda_{k_{i}}+\Lambda_{\ell_{fh}}}\right)>0,$ (A55)
$\epsilon_{\ell_{fu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{mh}k_{i}}^{\mathcal{C}}=\left(\frac{1}{1-\sigma_{fu}}-\frac{1}{1-\sigma_{mh}}\right)\left(\frac{\Lambda_{k_{i}}}{\Lambda_{k_{i}}+\Lambda_{\ell_{fh}}+\Lambda_{\ell_{mh}}}\right)>0,$ (A56)
$\epsilon_{\ell_{mu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{fu}k_{i}}^{\mathcal{C}}=\left(\frac{1}{1-\sigma_{mu}}-\frac{1}{1-\sigma_{fu}}\right)\left(\frac{\Lambda_{k_{i}}}{\Lambda_{k_{i}}+\Lambda_{\ell_{fh}}+\Lambda_{\ell_{mh}}+\Lambda_{\ell_{fu}}}\right)>0,$ (A57)
$\epsilon_{\ell_{mu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{mh}k_{i}}^{\mathcal{C}}=\left(\epsilon_{\ell_{mu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{fu}k_{i}}^{\mathcal{C}}\right)+\left(\epsilon_{\ell_{fu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{mh}k_{i}}^{\mathcal{C}}\right)>0,$ (A58)
$\epsilon_{\ell_{fu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{fh}k_{i}}^{\mathcal{C}}=\left(\epsilon_{\ell_{fu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{mh}k_{i}}^{\mathcal{C}}\right)+\left(\epsilon_{\ell_{mh}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{fh}k_{i}}^{\mathcal{C}}\right)>0,$ (A59)
and
$\epsilon_{\ell_{mu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{fh}k_{i}}^{\mathcal{C}}=\left(\epsilon_{\ell_{mu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{fu}k_{i}}^{\mathcal{C}}\right)+\left(\epsilon_{\ell_{fu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{mh}k_{i}}^{\mathcal{C}}\right)+\left(\epsilon_{\ell_{mh}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{fh}k_{i}}^{\mathcal{C}}\right)>0.$ (A60)
The effects of technological progress embodied in ICCT capital on the income shares of male and female skilled labor are positive, while those on the income shares of male and female unskilled labor are negative, if $\sigma_{fh}<0$, $\sigma_{mh}\simeq 0$, and $\sigma_{fu},\sigma_{mu}>0$. Thus, the magnitude of the effects are all proportional to the differences in the substitution parameters in the production function and the income share of ICCT capital relative to the income shares of other factors.
We next consider the effects on the four wage gaps. The effect on the skilled gender wage gap is
$\frac{d\ln\left(\left.w_{mh}\right/w_{fh}\right)}{d\ln q_{i}}=-\frac{1}{H}\left(\epsilon_{k_{i}\ell_{mu}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{k_{i}\ell_{fu}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{\ell_{mh}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{fh}k_{i}}^{\mathcal{C}}\right)\left(1-\Lambda_{k_{o}}\right)<0.$
The magnitude of the effect is proportional to the difference between the aggregate elasticities (A55). The effect on the unskilled gender wage gap is
$\frac{d\ln\left(\left.w_{mu}\right/w_{fu}\right)}{d\ln q_{i}}=-\frac{1}{H}\left(\epsilon_{k_{i}\ell_{mh}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{k_{i}\ell_{fh}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{\ell_{mu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{fu}k_{i}}^{\mathcal{C}}\right)\left(1-\Lambda_{k_{o}}\right)<0.$
The magnitude of the effect is proportional to the difference between the aggregate elasticities (A57). The effect on the male skill wage gap is
$\frac{d\ln\left(\left.w_{mh}\right/w_{mu}\right)}{d\ln q_{i}}=\frac{1}{H}\left(\epsilon_{k_{i}\ell_{fh}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{k_{i}\ell_{fu}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{\ell_{mu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{mh}k_{i}}^{\mathcal{C}}\right)\left(1-\Lambda_{k_{o}}\right)\\ +\frac{1}{H}\left(\epsilon_{k_{i}\ell_{fh}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{\ell_{mu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{fu}k_{i}}^{\mathcal{C}}\right)\left(\epsilon_{\ell_{fu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{mh}k_{i}}^{\mathcal{C}}\right)\left(1-\Lambda_{k_{o}}\right)\left(\frac{\Lambda_{\ell_{fu}}}{\Lambda_{k_{i}}}\right)>0.$
The magnitude of the effect is proportional to the differences between the aggregate elasticities (A56), (A57), and (A58). The effect on the female skill wage gap is
$\frac{d\ln\left(\left.w_{fh}\right/w_{fu}\right)}{d\ln q_{i}}=\frac{1}{H}\left(\epsilon_{k_{i}\ell_{mu}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{k_{i}\ell_{mh}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{\ell_{fu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{fh}k_{i}}^{\mathcal{C}}\right)\left(1-\Lambda_{k_{o}}\right)\\ +\frac{1}{H}\left(\epsilon_{k_{i}\ell_{mu}}^{\mathcal{C}}+\frac{1}{\gamma}\right)\left(\epsilon_{\ell_{fu}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{mh}k_{i}}^{\mathcal{C}}\right)\left(\epsilon_{\ell_{mh}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{fh}k_{i}}^{\mathcal{C}}\right)\left(1-\Lambda_{k_{o}}\right)\left(\frac{\Lambda_{\ell_{mh}}}{\Lambda_{k_{i}}}\right)>0.$
The magnitude of the effect is proportional to the differences between the aggregate elasticities (A55), (A56), and (A59). Thus, the magnitude of the effects on the four wage gaps is proportional to the differences in the substitution parameters in the production function and the income share of ICCT capital relative to the income shares of other factors.
Appendix F Additional Results
Table A2 report the changes attributable to the three types of technological change when labor is not perfectly mobile across sectors. Table A3 reports the effects of technological change embodied in ICCT capital on the relative wages and income shares of the four types of labor by country for the cases in which the supply of labor is perfectly inelastic and unit elastic. Table A4 compares the magnitude of the changes in relative wages and income shares attributable to ICCT equipment obtained by estimating and aggregating sectoral production functions with that obtained by assuming and estimating the aggregate production function.
Table A2: Effects of technological change under imperfect labor mobility, 1980–2005
$\Delta\ln q_{i}$ $\Delta\ln q_{o}$ $\Delta\ln A_{n}$
Relative wages
$\Delta\left(\frac{w_{mh}}{w_{fh}}\right)$ –0. 936 –0. 016 0. 072
(0. 056) (0. 001) (0. 005)
$\Delta\left(\frac{w_{mu}}{w_{fu}}\right)$ –0. 062 –0. 001 –0. 004
(0. 006) (0. 000) (0. 001)
$\Delta\left(\frac{w_{mh}}{w_{mu}}\right)$ 0. 240 0. 004 –0. 010
(0. 014) (0. 000) (0. 001)
$\Delta\left(\frac{w_{fh}}{w_{fu}}\right)$ 1. 337 0. 022 –0. 104
(0. 058) (0. 001) (0. 005)
$\Delta\left(\frac{\overline{w}_{M}}{\overline{w}_{F}}\right)$ –0. 251 –0. 004 0. 014
(0. 015) (0. 000) (0. 001)
$\Delta\left(\frac{\overline{w}_{H}}{\overline{w}_{U}}\right)$ 0. 564 0. 009 –0. 040
(0. 013) (0. 000) (0. 001)
Labor shares
$\Delta\left(\frac{w_{mh}\ell_{mh}}{py}\right)$ 0. 005 0. 000 0. 000
(0. 001) (0. 000) (0. 000)
$\Delta\left(\frac{w_{fh}\ell_{fh}}{py}\right)$ 0. 017 0. 000 –0. 001
(0. 001) (0. 000) (0. 000)
$\Delta\left(\frac{w_{mu}\ell_{mu}}{py}\right)$ –0. 036 –0. 001 0. 002
(0. 002) (0. 000) (0. 000)
$\Delta\left(\frac{w_{fu}\ell_{fu}}{py}\right)$ –0. 008 0. 000 0. 001
(0. 001) (0. 000) (0. 000)
$\Delta\left(\frac{\sum_{f}w_{f}\ell_{f}}{py}\right)$ 0. 001 0. 000 0. 000
(0. 004) (0. 000) (0. 000)
Notes: The changes in the levels of relative wages and labor shares attributable to technological change are computed using equations (A3) and (A4) of Proposition 5. Standard errors in parentheses are computed using bootstrap with 500 replications.
Table A3: Effects of technological change embodied in ICCT capital for each country
All AUS AUT CZE DEU DNK FIN GBR ITA JPN NLD USA
Relative wages, $1/\gamma=0$
$\Delta\left(\frac{w_{mh}}{w_{fh}}\right)$ –0. 886 –1. 351 –0. 565 –0. 770 –0. 879 –3. 325 –0. 264 –0. 938 –0. 552 –1. 318 –1. 301 –0. 770
(0. 052) (0. 082) (0. 033) (0. 048) (0. 053) (0. 239) (0. 013) (0. 057) (0. 036) (0. 099) (0. 092) (0. 043)
$\Delta\left(\frac{w_{mu}}{w_{fu}}\right)$ –0. 074 –0. 140 –0. 038 –0. 070 –0. 058 –0. 205 –0. 025 –0. 085 –0. 043 –0. 089 –0. 081 –0. 081
(0. 006) (0. 010) (0. 003) (0. 006) (0. 005) (0. 017) (0. 003) (0. 007) (0. 004) (0. 009) (0. 007) (0. 007)
$\Delta\left(\frac{w_{mh}}{w_{mu}}\right)$ 0. 262 0. 412 0. 226 0. 306 0. 216 1. 105 0. 102 0. 225 0. 223 0. 188 0. 419 0. 176
(0. 014) (0. 022) (0. 014) (0. 016) (0. 011) (0. 075) (0. 005) (0. 012) (0. 015) (0. 008) (0. 022) (0. 008)
$\Delta\left(\frac{w_{fh}}{w_{fu}}\right)$ 1. 283 1. 822 1. 037 1. 359 1. 239 5. 308 0. 303 1. 446 0. 829 1. 665 2. 240 0. 934
(0. 053) (0. 076) (0. 041) (0. 062) (0. 056) (0. 274) (0. 009) (0. 064) (0. 030) (0. 108) (0. 121) (0. 037)
Relative wages, $1/\gamma=1$
$\Delta\left(\frac{w_{mh}}{w_{fh}}\right)$ –0. 486 –0. 737 –0. 315 –0. 420 –0. 470 –1. 599 –0. 155 –0. 503 –0. 299 –0. 660 –0. 655 –0. 431
(0. 036) (0. 045) (0. 019) (0. 027) (0. 028) (0. 110) (0. 006) (0. 030) (0. 017) (0. 047) (0. 044) (0. 024)
$\Delta\left(\frac{w_{mu}}{w_{fu}}\right)$ –0. 096 –0. 177 –0. 050 –0. 085 –0. 074 –0. 238 –0. 033 –0. 109 –0. 053 –0. 102 –0. 100 –0. 108
(0. 010) (0. 021) (0. 007) (0. 011) (0. 010) (0. 032) (0. 006) (0. 015) (0. 008) (0. 014) (0. 013) (0. 014)
$\Delta\left(\frac{w_{mh}}{w_{mu}}\right)$ 0. 264 0. 422 0. 230 0. 303 0. 212 0. 992 0. 105 0. 229 0. 206 0. 177 0. 400 0. 189
(0. 013) (0. 020) (0. 012) (0. 017) (0. 011) (0. 058) (0. 009) (0. 012) (0. 011) (0. 012) (0. 022) (0. 011)
$\Delta\left(\frac{w_{fh}}{w_{fu}}\right)$ 0. 767 1. 092 0. 634 0. 807 0. 717 2. 853 0. 196 0. 844 0. 507 0. 876 1. 226 0. 567
(0. 039) (0. 044) (0. 025) (0. 038) (0. 032) (0. 131) (0. 006) (0. 037) (0. 017) (0. 054) (0. 059) (0. 023)
Labor shares, $1/\gamma=0$
$\Delta\left(\frac{w_{mh}\ell_{mh}}{py}\right)$ 0. 006 0. 005 0. 004 0. 004 0. 005 0. 015 0. 002 0. 004 0. 004 0. 005 0. 009 0. 002
(0. 001) (0. 001) (0. 001) (0. 001) (0. 001) (0. 002) (0. 000) (0. 001) (0. 000) (0. 001) (0. 001) (0. 001)
$\Delta\left(\frac{w_{fh}\ell_{fh}}{py}\right)$ 0. 017 0. 016 0. 016 0. 022 0. 015 0. 022 0. 013 0. 008 0. 010 0. 006 0. 007 0. 021
(0. 001) (0. 001) (0. 001) (0. 001) (0. 001) (0. 001) (0. 000) (0. 000) (0. 000) (0. 000) (0. 000) (0. 001)
$\Delta\left(\frac{w_{mu}\ell_{mu}}{py}\right)$ –0. 039 –0. 074 –0. 025 –0. 027 –0. 024 –0. 107 –0. 013 –0. 037 –0. 032 –0. 032 –0. 050 –0. 036
(0. 002) (0. 004) (0. 002) (0. 002) (0. 002) (0. 007) (0. 001) (0. 002) (0. 002) (0. 002) (0. 003) (0. 002)
$\Delta\left(\frac{w_{fu}\ell_{fu}}{py}\right)$ –0. 008 –0. 011 –0. 006 –0. 007 –0. 005 –0. 027 –0. 005 –0. 006 –0. 005 –0. 004 –0. 004 –0. 008
(0. 001) (0. 001) (0. 001) (0. 001) (0. 001) (0. 003) (0. 000) (0. 001) (0. 001) (0. 000) (0. 001) (0. 001)
Labor shares, $1/\gamma=1$
$\Delta\left(\frac{w_{mh}\ell_{mh}}{py}\right)$ 0. 014 0. 014 0. 010 0. 010 0. 011 0. 028 0. 004 0. 009 0. 008 0. 012 0. 019 0. 010
(0. 001) (0. 001) (0. 001) (0. 001) (0. 001) (0. 002) (0. 000) (0. 000) (0. 000) (0. 001) (0. 001) (0. 001)
$\Delta\left(\frac{w_{fh}\ell_{fh}}{py}\right)$ 0. 020 0. 020 0. 020 0. 026 0. 017 0. 024 0. 016 0. 009 0. 012 0. 006 0. 008 0. 026
(0. 001) (0. 001) (0. 001) (0. 001) (0. 001) (0. 001) (0. 001) (0. 000) (0. 000) (0. 000) (0. 000) (0. 001)
$\Delta\left(\frac{w_{mu}\ell_{mu}}{py}\right)$ –0. 070 –0. 128 –0. 046 –0. 046 –0. 042 –0. 173 –0. 024 –0. 066 –0. 052 –0. 051 –0. 083 –0. 067
(0. 005) (0. 009) (0. 004) (0. 004) (0. 004) (0. 016) (0. 003) (0. 006) (0. 005) (0. 005) (0. 008) (0. 006)
$\Delta\left(\frac{w_{fu}\ell_{fu}}{py}\right)$ –0. 007 –0. 009 –0. 005 –0. 008 –0. 004 –0. 019 –0. 007 –0. 004 –0. 003 –0. 003 0. 001 –0. 010
(0. 001) (0. 002) (0. 001) (0. 001) (0. 001) (0. 005) (0. 001) (0. 001) (0. 001) (0. 001) (0. 001) (0. 001)
Technological change embodied in ICCT capital
$\Delta\ln q_{i}$ 2. 001 2. 552 2. 005 1. 333 1. 379 3. 365 1. 332 1. 599 1. 754 1. 381 1. 821 1. 954
Notes: Country names are abbreviated as follows: AUS, Australia; AUT, Austria; CZE, the Czech Republic; DEU, Germany; DNK, Denmark; FIN, Finland; GBR, the United Kingdom; ITA, Italy; JPN, Japan; NLD, the Netherlands; and USA, the United States. The changes in the levels of relative wages and labor shares attributable to technological change are computed using equations (A1) and (A2) of Proposition 4. The first column reproduces the first and third columns of Table 7. The second to last columns report the effects evaluated at the sample means over time for each country. Standard errors in parentheses are computed using bootstrap with 500 replications.
Table A4: Results from sectoral and aggregate production functions
$\Delta\ln k_{i}$ $\Delta\ln q_{i}$
Sectoral Aggregate Sectoral Aggregate
Relative wages Relative wages
$\Delta\ln\left(\frac{w_{mh}}{w_{fh}}\right)$ –0. 884 –1. 109 $\Delta\left(\frac{w_{mh}}{w_{fh}}\right)$ –0. 886 –1. 015
(0. 070) (0. 097) (0. 052) (0. 061)
$\Delta\ln\left(\frac{w_{mu}}{w_{fu}}\right)$ –0. 070 –0. 075 $\Delta\left(\frac{w_{mu}}{w_{fu}}\right)$ –0. 074 –0. 073
(0. 005) (0. 006) (0. 006) (0. 005)
$\Delta\ln\left(\frac{w_{mh}}{w_{mu}}\right)$ 0. 222 0. 236 $\Delta\left(\frac{w_{mh}}{w_{mu}}\right)$ 0. 262 0. 254
(0. 014) (0. 017) (0. 014) (0. 015)
$\Delta\ln\left(\frac{w_{fh}}{w_{fu}}\right)$ 1. 036 1. 270 $\Delta\left(\frac{w_{fh}}{w_{fu}}\right)$ 1. 283 1. 436
(0. 069) (0. 096) (0. 053) (0. 063)
$\Delta\ln\left(\frac{\overline{w}_{M}}{\overline{w}_{F}}\right)$ –0. 228 –0. 276 $\Delta\left(\frac{\overline{w}_{M}}{\overline{w}_{F}}\right)$ –0. 249 –0. 276
(0. 017) (0. 022) (0. 014) (0. 015)
$\Delta\ln\left(\frac{\overline{w}_{H}}{\overline{w}_{U}}\right)$ 0. 454 0. 531 $\Delta\left(\frac{\overline{w}_{H}}{\overline{w}_{U}}\right)$ 0. 564 0. 602
(0. 024) (0. 032) (0. 013) (0. 014)
Labor shares Labor shares
$\Delta\ln\left(\frac{w_{mh}\ell_{mh}}{py}\right)$ 0. 092 0. 101 $\Delta\left(\frac{w_{mh}\ell_{mh}}{py}\right)$ 0. 006 0. 006
(0. 015) (0. 018) (0. 001) (0. 001)
$\Delta\ln\left(\frac{w_{fh}\ell_{fh}}{py}\right)$ 0. 976 1. 211 $\Delta\left(\frac{w_{fh}\ell_{fh}}{py}\right)$ 0. 017 0. 019
(0. 071) (0. 098) (0. 001) (0. 001)
$\Delta\ln\left(\frac{w_{mu}\ell_{mu}}{py}\right)$ –0. 130 –0. 135 $\Delta\left(\frac{w_{mu}\ell_{mu}}{py}\right)$ –0. 039 –0. 037
(0. 004) (0. 003) (0. 002) (0. 002)
$\Delta\ln\left(\frac{w_{fu}\ell_{fu}}{py}\right)$ –0. 060 –0. 059 $\Delta\left(\frac{w_{fu}\ell_{fu}}{py}\right)$ –0. 008 –0. 007
(0. 005) (0. 005) (0. 001) (0. 001)
$\Delta\ln\left(\frac{\sum_{f}w_{f}\ell_{f}}{py}\right)$ –0. 003 0. 012 $\Delta\left(\frac{\sum_{f}w_{f}\ell_{f}}{py}\right)$ –0. 001 0. 006
(0. 007) (0. 009) (0. 004) (0. 004)
Notes: The figures in the upper and lower halves of the first two columns are computed using equation (18) of Proposition 2 and equation (15) of Proposition 1, while the figures in the upper and lower halves of the last two columns are computed using equations (21) and (22) of Proposition 3. The figures in the first column of the left panel are reproduced from the fifth columns of Tables 5 and 6, while those in the first column of the right panel are reproduced from the first column of Table 7. The figures in the second columns of the left and right panels are computed using the estimates of the aggregate production function parameters reported in the last row of Table 2.